A Complete Bibliography of The American Statistician: 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

07 October 2024 Version 1.06

Title word cross-reference

(X_{1} < X_{2} ⋯ < X_{k}) [HL96]. 1: m [Lev97]. 2 × 2 [MW92, MH93]. 2 × k [GK99]. ^{T M} [Pie90]. _{100p} [Stu90]. X̄ [Zeh91]. F [LD98, ML92b, Pei93, PS98b]. L [Hos92]. L1 [All97]. P [Don99, SSC99, Sch96a, Ber96a, MW92]. p^{k} [Bec96]. p^{n−k} [Bec96]. R [BK92]. R^{2} [And94, Mag90, SW91]. S^{2} [Zeh91]. t [CA97, DG97, Gro90, KG90, LaM94b, LD98, MC91]. U [Sch91a]. x_{2} [Sch93a]. x_{2}|x_{1} [Sch93a].

-Distance [All97].

1 [Kah91]. 1-2-3 [Kah91]. 14th [RvE97]. 1919 [Ben93]. 1990 [TW91]. 1991 [Wan91]. 1994 [CR95].

2} [Won98]. 2.0 [Maa99]. 21st [McD99]. 240 [DeL94, Fle94, Kut94, Lam94a, LDL^{+}94, Rog94]. 2nd [Hol99, KK99b, Mac99].

3 [Kah91]. 3.1 [Spi90]. 3rd [Hol99, Joh99].

43 [Ang90, GY90]. 46 [DeL94, Fle94, Kut94, Lam94a, LDL^{+}94, Rog94]. 47 [WS94]. 48 [DN94a]. 4th [Kaf99b].

5} [Gol99, Tan99]. 5.0 [Mac93]. 52 [Ano98n]. 5V [KDV94].

6th [Pil98].

7th [Obr98].

9th [Pil98].

A. [Wan91]. Abandon [Dal90a, Dal90c]. Absolute [Bas97, SGC90]. Abstract [Goo90]. Abuses [Pau92]. Academia [Bic95]. Academic [CRV95]. Accent [Ano91a, Ano91b, Ano92a, Ano93a, Ano95a, Ano95b, Ano95c, Ano96a, Ano96b]. Accepted [GKR94, PG97]. accessible [DK95]. Account [MBB90]. Accuracy [TM99]. Acknowledgement [DN94a]. Across [MW92]. Active [Maa99]. Activities [S^{+}90]. Activity [Die93]. ActivStats [Maa99]. Additive [Hil93, S.N91]. Address [CJ92]. Adjusted [GCGG95, SM91]. Adjustment [LK99]. Advanced [RSW99]. Advantage [HS94]. Advertisements [DWBQ98]. Advertising [Dan92]. Advisory [Eld90]. Against [Hel95, Jam99, Sil97, Gab91]. Agencies [Eld90, Gon94]. Aggregation [Lic90]. Ahead [Gna90]. Aided [GPA97]. Alai [CG95a]. Algebra [Bag95, Nas95]. Algebraically [Sha91b]. Algorithm [BS95, CG95b, Gan95, Mil96, Nav97b, Pos93]. Algorithms [MS98]. Alignment [Faw90]. Alternative [BH98, Coh97]. Alternatives [MKV92, Suk92]. Amer [Ang90, DN94a, GY90]. Amer. [WS94]. America [NOMLT95]. American [Ano91h, DeL94, Fle94, Kut94, Lam94a, LDL^{+}94, Rog94, Ano98c, HBF90, Lan98, Lan99a, MMR90, MMS90, Wio90, ZH90]. Among [SR97, Wil91a]. Amstat [DWBQ98]. Analogues [VBB94]. Analyses [HV95, PS94]. Analysis [AR91, Ank98, Atk95, CT91, Cle99c, Coh97, CW99, DeL94, Edl91, FM92, Fle94, FN92, GW94, HG97, HHB91, Hol99, Hsi95, Joh99, JK99, KT94, KDV94, Kut94, Lam94a, Lar92, LDL^{+}94, MV94, Mol91, NK96, NL95, OW91, OCE93, Pei93, Pie90, PR98, Pro91, Rog94, Rum99, Sap91, Sea94, She98, SC99, SW90, Smi98, Tho95, Van96b, Vis91, Wat95, WC91, Alb92, Hol99, Nor95]. Analyzing [Cot98, Faw90]. Animal [ES93]. annotated [Wio90]. Anomalies [God91, HBC^{+}91]. ANOVA [BK96, Hin96, Saw92, Sch93d, Ver99]. Any [O’C90]. APL [RM91]. Appear [Wri90]. Apple [SMP97]. Application [DuM99a, DuM99b, HBC^{+}91, LS99b, Mad99, Nas94b, OS99, RS93, Rie90]. Applications [Ber97c, BTV90, BHO98, DB97, Epp93, GKT98, Kah91, NS99, WTH95, Knü94]. Applied [Bon98, GH91, KK99b, SM97, SW90, Hol99, Hol99, Joh99]. Approach [Alb97c, Bri96, Chu99, EGCD91, FS93, FN92, Gro91, Hil96b, Lor90, MSS91, Oks91, Van96b, MKV92]. Approximate [AC98a, AW91a, BfJTH93, LT96, LC99, ME95]. Approximation [BS98, CGO91, GC99, PA91, Rea98, McC90]. Approximations [Huz99]. areas [McC90]. Arguments [GM96]. ARIMA [McL99]. Arising [Jon99]. Arithmetic [Ste96]. Arithmetic-Geometric-Harmonic [Ste96]. ARMA [McL99]. Arrival [BGG98]. Articles [Pet99]. Artifacts [JSM95]. Artificial [DG97]. ASA [Ano90a, Ano90b, Ano90k, CR90, Haa90, Inm90, KLM90, MBB90, Nor90]. Ask [Sim98]. Aspect [BWM98]. Aspects [CS94, JK90]. Assess [Loo95]. Assessing [GKT91, McC98a, McC98b, McC99]. Assessment [Rus98]. Assessors [TM99]. Association [Ano91h, Lee93, MMS90, HBF90, MMR90, Wio90, ZH90]. Assumptions [HM93]. Asymptotic [KB93, Sch91a]. Audio [Pos90]. Augmenting [CH96]. Autocovariance [Per93]. Automatic [OL96]. Average [RS91b]. Averages [JK90]. Awareness [CF91, Gon94]. AWAY [BH98]. Axum [Edl91].

Babies [CP99]. Back [Ano90c, Ano90d, Ano90e, Ano90f, Ano91c, Ano91d, Ano91e, Ano91f, Ano92b, Ano92c, Ano92d, Ano92e, Ano93b, Ano93c, Ano93d, Ano93e, Ano94a, Ano94b, Ano94c, Ano94d, Ano95d, Ano95e, Ano95f, Ano95g, Ano96c, Ano96d, Ano96e, Ano96f, Ano97b, Ano97c, Ano97d, Ano97e, Ano98d, Ano98e, Ano98f, Ano98g, Ano99b, Ano99c, Ano99d, Ano99e, SL92, SW90]. Back-of-the-Envelope [SL92]. Bad [Ver99]. Badly [Hog91]. Bagplot [RRT99]. Balanced [Sch93d]. Ball [Gun93]. Baseball [Bor99, HV95, Lac90, RMY96, WC97]. Based [Ano98b, Ano98t, BR97, BN98, Bas97, Bis98, Czi98, DS98, Ede90a, Haw91, Kot91, LS97b, Mag90, MV94, ME95, PS98a, Pet99, SO94, Sho95, SS98, Spu98, Var98, Wan97, Wan99, Coh97]. Basic [Low98, Pet99, Smi98, Wel98]. Basketball [Car96, SMH91, SSH96, SS99, War95b, Wes90]. Basu [BHO98]. Bayes [Alb97a, Alb97b, Alb97c, Ber97a, Ber97b, Fre97, LS99a, Lin97, Moo97a, Moo97b, Moo97c, Sch97, Sho97, Wit97a, Wit97c]. Bayesian [Sho98, Alb92, Alb93, BCJ97, Ber96c, Ber97c, CU92, CB95b, DuM99a, DuM99b, Gel98, LS97b, LS99b, Mad99, Mat99, NS97, OS99, SG92]. Bayesians [Ros96a]. Be [Kon95, Olk92, SS99]. Bearing [Gun93]. Beginners [Alb97a, Alb97b, Ber97a, Ber97b, Fre97, Lin97, Moo97a, Moo97b, Moo97c, Sch97, Sho97, Wit97a, Wit97c]. Behavior [Mag98]. Bernoulli [Ber96c, LT96, VBB94, War97]. Best [BR97, BH98, DB97, Sea97, SM91]. Best-of-Seven [BH98]. Bet [CW93]. Beta [Bai92, Cha95, Pos93]. Better [AC98a]. Between [JK99, LL97, Nav97a, OW91, Sap91, Sch93a, Sch96c, Sha91b, SF93]. Beyond [MH93]. Bias [Sal99, Sto98]. Biased [Ano90g, Ano91g, SP94]. bibliography [Wio90]. Bin [Wan97, Wan99]. Binary [CW92, EP91, Jor94, Lee93, Mat99, PPS96, Stu90]. Binomial [AC98a, BCJ97, DN94b, DN94c, LC99, Lin92, Pul90, SL92, Tie94, VBB94, DN94a, Knü94]. Biological [GPA97]. Biomedical [LD98]. Biopharmaceutical [Fre90]. Biostatistical [AW91b, NOMLT95]. Biostatisticians [DAF^{+}94]. Biostatistics [Der93, Her90, Kle95]. Birth [Nun92]. Birthday [Nun92]. Bivariate [Beh90, Cai94, GM91, Gre96, JH96, RRT99]. Blind [Sto98]. BLiP [LT97]. Block [FM92, Gat95]. Blocking [Faw90, Old95]. Blocks [MG97]. BLUE [BF91]. BLUEs [VAD90]. BLUP [Sea97]. BMDP [KDV94]. BMDP-5V [KDV94]. Boards [Gib90]. Bonaparte [Gun98]. Bonferroni [KG90]. Book [Agr96, AB94, Ame93, And96, And99, Ank98, Bag95, Bar97, Bar98, Bar99, Bis95, Bol91, Bon98, Boo97, Boy94, Can98, Car93, Che94b, Che94a, Cho97, Cla93, Cle99a, Cle99b, Cle99c, CW94, Coo91, Cot98, Cry91, Cry97, D’A96, Del97, Del95, Die97, Dob93a, Dob93b, Dob98, Don93, Don96, Dou97, DS91, DN92, Elt96, Eme96, Fra97, Fre94, Fri96, GKM^{+}91, Gar97, Gas97, Gol98, Gol99, Gor96, Gun97, Gun98, Gup94, Hal92, Hal94a, Hal94b, Hal95a, Hal95b, Hal95c, Han95, Har97, Har95, Hay94, Hay96, Hay99, Hei91, Hel99, Her95, Hil93, Hil94, Hil96a, Hoe98, Hoe97, Hof91, Hol99, Hol95, Jam99, Joh96, Joh99, Jos93, Kaf97, Kaf98, KK99a, Kaf99a, KK99b, Kaf99b, Kas94, Kas98, KT94, Kin95, Koe97a, Koe97b, Koo90, Koo97]. Book [KDV94, Len98, Les95, Lev92, Loc93, Low98, Lur95, Lye96, LLL91, MW95, Maa99, Mac99, Mal94, Mal97, Man91, Mar94, Mcc97, Mid96, Mon97a, Moo96, Mor91, Nab97, Nag90, Nag98, Nas95, NO98, Nei96, Nei93, Nic94a, Nic95, Nor95, Obr98, OL96, Par98a, Par98b, Pec97, Pet99, PRSE94, Pic97, Pil98, PD93, Pos94, Pos96, RS95, Ree94, Rei94, RR92, RR93, Ros93, Ros94, Ros95b, Ros96b, Rum98, Rum99, Rus98, SP98, Sap95, Sax94, Sch96b, Sch99, Sch93c, Sch93e, Sch93f, Sha94, She98, She93, Sho96a, Sho96b, Sho98, Sim96, Smi97, Smi98, Sne98, Sol95, Son90, Spi90, Tan92, Tan95, Tan99, The90, Tho95, Tom93, War95a, War93, Was93, Wat96, Wat95, Wea97, Wel98, Win95, Wit93, Wit94, Wit97b, Won98, Woo93, Wue98]. Books [Ano93f, Ano93g, Ano94e, Ano97m, Ano97n, Ano97o, Ano97p, Ano98i, Ano98p, Sch93d]. Boosters [MZ92]. Bootstrap [All97, BS98, Pol99, Wil94b]. Both [Atk94]. Bound [Gre96, Lev97, PN95]. Bounded [JH96]. Bounds [Ass96, CT91, Nav97a, Pol99, TM95, WB91, Wri90]. Bowling [CS94]. Box [HN98]. Boxplot [RRT99]. Breakdown [Bas91]. Breeze [Bru91]. Breeze/STAT [Bru91]. Bribery [CG95a]. Bridge [Dun94, Per98]. Brief [Ano91h, Rot90, MMR90]. Bringing [WC97]. Bronchial [HHB91]. Buffon [Rie90]. Building [MG97]. Built [Sea97]. Built-In [Sea97]. Bulbs [Daw96, MD96]. Business [Bon98, BO92, Hil96b, PS98a, Smi98, Rum98].

Cables [Per98]. Calculating [Edw96, Tie94, Tie98, TM99]. Calculation [HL96, Lor90]. Calculations [DP91]. Calculator} [Bar98, NO98]. Calibration [FS93]. Can [Olk92, SS99, Wat91]. Canonical [Ren92]. Carcinogenesis [HHB91]. Career [BK93, But90]. Carlo [BS98, KCGN98, Rie90, RM91]. Case [FT93, PS98a, Wie99, DK95].

Case-Based [PS98a]. Casebook} [Smi98]. Casino [CW93]. Catastrophic [MZ92]. Categorical [Fri95, Gan95, God91, MGLK99, Sim98, vS96]. Cauchy [AB92]. Causal [Old95]. Cautionary [HL98, HS92, Hsi95]. CD [Sne98]. CD-ROM [Sne98]. Celebrates [KLM90]. Celebration [Ano90a, Ano90b, MBB90]. Cement [PR98]. Censored [BS95, WT92]. Census [Oko99]. Censusing [RS93]. Center [AW91b, BCJ97, Ede90a, WKB^{+}95, KM94]. Central [DR99, Jol95]. Century [McD99]. Certification [Boe94, Fie94, Gro94, Gun94, Hod94, HSF^{+}94, Imr94b, Imr94c, Kad94, Kaf94, Loh94, Sti94, Ver94]. Chain [KCGN98]. Challenge [Rus98]. Challenges [Cha99, CRV95, Kir91]. Chances [Der93]. Change [vS96]. changed [Boa94]. Chapters [MMS90]. Characteristic [Epp93, Luc97, Wal95a, WTH95]. Characterization [Ano90h, BF91, BK99, GY89, GY90, Kel94]. Characterize [McC94]. Characterizing [SBL97]. Charts [CG95a]. Cheating [Kva96]. Chebyshev [SS97]. Chemical [KM92, RKW98]. Chernoff [Nav97a, PN95]. Chi [Bai92, Dri99, Lev97]. Chi-Square [Dri99]. Chi-Squared [Bai92, Lev97]. Choice [Wan97, Wan99]. Choosing [NP97]. Churchill [CR95]. Citizenry [Sch90]. Class [Ahm96, FK97, Gel98, Gro98]. Class-Participation [Gel98]. Classes [Mag96]. Classification [DS98, SS98]. Classroom [Kva96]. Clinical [AW91b, Bli90, Hsi91, WKB^{+}95]. Clinicians} [Bar99]. Closeness [FK97]. Cluster [WC91]. Cocaine [Shu91]. Coefficient [AH95, CW92, RvE97, Tur97, VM90, Van96a, Wri92, ZM94, vM93]. Coefficients [Bri94, KG90, Sha91b, SR97, Wil91a, Wri90]. Coin [Kon95]. Coincidences [Han92b]. Cointegration [Mur94]. Collaborators [Ano90l, Ano91j, Ano92h, Ano93j, Ano94g, Ano95i, Ano96h, Ano97a, Ano98a, Ano99a]. Collection [FT93]. Collections [Hay96, JR92]. Collector [Daw91, Rea98]. Colleges [LW98]. Column [Dal90a, Ehr90, Nor90, Sch90, Sch91c]. Combining [Peñ97, SC94]. Come [SATV94]. Comment [Boe94, CW92, Fie94, Gro94, Gun94, Hod94, Hog92b, HSF^{+}94, Kad94, Kaf94, Loh94, Min92, MD97, PG97, Sey91, Sol92, Sti94, Ver94, Wel92]. Comments [PR94]. Committees [Eld90]. Common [Dav95, Dav98a, MW92, OCE93, Tie98]. Community [AS99]. COMPACT [Bli90]. Companion [Sne98]. Comparative [Nor95, OCE93]. Compare [Bri96, LC99]. Comparing [Ano94f, GCGG95, Han92a, Hos92, Kha92, She97, WS93, WS94]. Comparison [Git94, LT96, Loc93, Mar94, Mat99, MV94, MW92, Mon97b, Nav97a, Nay94, Sav90]. Comparisons [And94, Day98, PF95, Wil91a, ZPH99]. Competitive [ST95, SB98]. Complementary [Pie92]. completely [HB96]. Complex [Coh97, KG90]. Complications [DR99]. Component [AB92, CM92, NK96, RS91a]. Components [BK99, HL98, JK99, Ren92]. Computations [Kha91, RM91]. Computer [Ano90g, Ano91g, Bli90, DeL94, DP91, Fle94, GPA97, Kut94, Lam94a, LDL^{+}94, NL95, Rog94, Sch93d, Sul96, McC90, Bag95, Nas95]. Computer-Aided [GPA97]. Computers [Ano90j, Obe91, Mar94].

Computing [Ano90r, Ano90s, Ano91o, Ano91p, Ano91q, Ano91r, Ano93o, Ano93p, Ano95q, Ano95r, Ano95s, Ano96q, Ano96r, Ano96s, Ano96t, Ano97q, Ano98r, Ano98s, CR90, Cha99, Dal90a, Dal90c, GKT98, Hyn96, MHT90, Nag98, New92a, New92b, Sea94, TM95]. Concentration [GH91]. Concept [Sul96]. Concepts [Bar99, BO92, Cha99, DS98, Dun99, Fra92, Kaf99b, Kha91, MMCsS96, Pau92, SS98]. Conceptual [Fri95]. Concerning [Goo90]. Concordance [RS91b]. Concrete [Loo97]. Condition [McL99]. Conditional [BL94, Gre91, GAR92, MH93, PP98]. Conditionality [Hel95]. Conditionally [GM91]. Conditionals [MS94]. Conditions [MM90]. Conference [Wat95]. Confessions [Kon95]. Confidence [Ano90i, AS98, Ber96a, Dou93, Ede90a, Fer90, FS93, Juo93, LD94, LK99, ME95, MW92, Nic94b, Pes90, Pol99, SL92, Sch96c, TM95, TL93, Van96a, Ver99, WL99, Wri90, Ang89, Ang90, Khu93]. Confounding [Old95]. Connection [Sap91, Sch96c, Wri92]. Connections [LL97]. Conservative [Nic94b]. Considerations [NK96]. Consistency [HS91]. Consonance [NP97]. Constructed [LaM94b]. Constructing [BK96, God91]. Construction [JT96, Nic94b]. Consultant [Dal90b, ES93, WKB^{+}95]. Consulting [Eng93, Hog92b, KM94, Kir91, Loh95, Min92, NOMLT95, Sol92, Wel92, Wil92]. Contingency [GK99, Gre91, MW92, MH93]. Continuity [LC99]. Contribution [Rie90, Sne90]. Control [BO92, CG95a, Git94, Kas98, PF95, Sto98, Str90]. Controversies [Kaf99b]. Controversy [Vos99]. Convergence [Mil96]. Conversion [Hil96a]. Cook [HH95, OH95]. Cooperative [Die93]. Coordination [WKB^{+}95]. COPY [Hal92, Hal92]. Core [AW91b, Hog94]. Cores [Der93]. Corner [Ano98t, Bis98, Czi98, Spu98, Var98]. Corrected [ML92a, ML92b]. Correction [Ang90, Ano90g, Ano91g, Ano92f, Ano92g, Ano93h, Ano93i, Ano94f, Ano95h, Ano96g, Ano97f, Ano98h, Ano98i, Ano99f, Bro94, DN94b, GY90, KB93, LC99, Mur94, Wan99, WS94, Yu94a]. Corrections [Ano90k, Ano90h, Ano90i, Ano90j, Ano91h, Ano91i, Dav98a]. Correlated [PPS96, VBB94]. Correlation [DGC97, GK99, Gre96, KAB^{+}97, MC96, Nel98b, RM94, RvE97, SR97, TO93, Tur97, VM90, VAD90, Wil91a, Wri92, ZM94]. Correspondence [Tho95]. Corroborate [CG95a]. Counterexamples [Hel95, JT96]. Counterparts [Epp93]. Counting [Edw96]. Coupon [Daw91]. Course [BO92, DNG98, Gab91, Hay96, Mac99, Pil98, SO94, Ste90]. Courses [AC98b, Bir95, Bra96a, CRV95, Eng93, Kle95, Led95, LW98, PS98a, SM97, SR96, WC97]. Court [GKR94, HS94, PG97]. Covariance [FM92, MS98, Van97]. Covariate [AFV96]. Coverage [Pol99]. Cox [Hsi95]. CQI [Bry99b, Cob99b, Gro99b, Hig99a, Hog99a, Hog99b, New99]. Craig [DK95]. Criteria [Day98]. Criterion [NP97]. Critical [HV95]. Cross [Gna90]. Cross-Disciplinary [Gna90]. Crossing [He97]. Crossover [PS94]. Cumulants [BL94, Smi95]. Cumulative [Tie94, Tie98]. Curious [Per93, Sil97]. Curriculum [Ano98b, Ano98t, BN98, Bis98, Czi98, MHT90, SO94, Spu98, Var98]. curve [McC90]. Curves [FT93, GKT91, He97, JR92, Mon95]. Cuthbert [Hun98]. Cutter [Sto98].

Dallal [DeL94, Fle94, Kut94, Lam94a, LDL^{+}94, Rog94]. Daniel [Hun98]. Dare [Dal90a, Dal90c]. Data [Alb97c, AR91, Bar98, Bec96, BM91, Can98, Cha99, Cle99c, Coh97, Daw96, DGC97, DuM99a, DuM99b, Dun94, Edl91, Fri95, Gol92a, Gon94, GR96, Han98, HBC^{+}91, Hay96, Hsi95, KK91, KG90, KG98, Kot91, LW90, LS99b, MD96, Mad99, McN92, MGLK99, OS99, PB90, PR98, Rum99, Sch93d, Sea94, SC99, SW90, Str90, Stu90, Vis91, WT92, Wal98, Wan97, Wan99, Wat95, WC91, Cle99c, Mon97b, Rum98, Sch99]. Data-Based [Wan97, Wan99]. Data-Oriented [Alb97c]. DBMS [Hal92, Hal92]. DBMS/COPY [Hal92, Hal92]. Deal [MCDD91a, MCDD91b, Sey91]. Dealing [Kir91]. Deception [Gun98]. Decision [Gel98, HS91, Rum98, Sch99, SRW95]. Decisions [SRW95]. Defectives [Wri90]. Deficiencies [GM96]. Defying [Han92b]. Degree [Lev97]. Degrees [AW91a, Lee93, LW98, Sul96]. Delta [Oeh92]. Deming [Boa94, Git94, Hah95]. Demonstrating [Old95]. Demonstration [Lee95, Lev93]. Demonstrations [Gel98]. Density [Ano90q, BfJTH93, HN98, Hyn96]. Department [CRV95]. Dependence [WSS93]. Dependent [Ano93h, Beh90, Dav93, SK93]. Depends [Sim98]. Depicting [Wai96]. Derivation [BL94, VM90]. Derivations [Sam92]. Derive [Ano96g, LC99, Wal95b]. Derived [BSY92, MS95]. DERIVE(R) [Nas95]. Deriving [BC92]. Describing [Gro91]. Description [Kah91, Sch93a]. Design [AC98b, Ank98, GM96, Gun93, Law96, LE93, Mat99, MKV92, Pie90, PS94, RKW98, Par98b]. Designed [BM91, Czi99]. Designing [AC98b]. Designs [CH96, Faw90, FM92, Gat95, LS97b]. Desk [BM91]. Detecting [Bor97, DGC97]. Determination [Ano95h, Bro93, Bro94, CW92, Sha91b]. Determinations [Dal90b]. Developed [Coh97, Vis91]. Developing [PS98a]. Developments [Ano90r, Ano91o, Ano91p, Ano93o, New92a, New92b, Zur91]. Deviant [Olk92]. Deviates [Lee93]. Deviation [MN97, O’C90]. Deviations [SGC90]. Device [Old95]. Diagrams [LL97, Loh95]. Diamond [Unw91]. Dice [KK91]. Did [Ben93]. Difference [SF93]. Differences [Sho95]. Different [Wie99]. Differential [BSY92]. Differently [PF95]. Differing [KG95]. Difficult [Wat91]. Digesting [Bra96b]. Digits [Ley96]. Dilemma [MCDD91a, MCDD91b, Sey91]. Dimensional [EP91, LT97, ZCM95]. Dirac [AT99]. Direction [Hig99c]. Disciplinary [Gna90]. Discipline [Hig99c]. Discovery [Bar98, Cle99c]. Discrete [BSY92, JSM95, Lin92, Nav97a, Pes90, RM91, TL93, WW97]. Discriminant [Mol91, OW91, Ren92, Sap91]. Discriminate [OW91]. Discussion [Bis98, Bry99a, Bry99b, Cob99a, Cob99b, Czi98, Fre97, Gro99a, Gro99b, KCGN98, Lin97, LS99b, Mad99, New99, OS99, Sch97, Sho97, Spu98, Var98, Wit97a, Wit97c]. Dispersion [Sho95]. Display [MC96, SC99]. Displaying [Ber96b, Faw90, JR92, KAB^{+}97]. Distance [All97]. Distinguish [Wal95a]. Distinguishing [HB96]. Distribution [AS98, BTV90, Cha95, CA97, Dun99, Ede90a, GKT91, Gho96, Gro91, Jol95, LT97, Ley96, Mad98, McC94, Mon95, Pos93, RMY96, SS91, Ste92, WTH95, Kel94, Knü94]. Distribution-Fitting [Mad98]. Distributional [Bai92, Hos92, Jon99]. Distributions [AB92, BS99, BM94, BSY92, Dan92, Dri99, Fer90, GM91, JH96, Lee93, LC99, Lin92, PA91, Pes90, SS91, SBL97, SS97, SC97, Tie98, Wal95a, WW97, Wel98]. Divisions [Gro90]. Do [PF95, Wat91]. Does [Bar90, Daw96, HS94, MD96, McC94, Wal95a]. Dog [Mur94]. Dose [AR91, Stu90]. Dose- [Stu90]. Dose-Response [AR91]. Dot [Wil99]. Double [Edw96]. Double-Elimination [Edw96]. Drug [Bor97]. Drunk [Mur94]. dual [Knü94]. Dummy [Coh91]. Dynamic [Atk95, HBC^{+}91]. Dynamics [GM96].

Early [Rie90]. Easier [Dav98b, Wat91]. Easy [SR96]. Eckler [Wan91]. Ecological [WC91]. Econometric [KT94]. Economics [Bon98, Chu99]. ed [Pil98, Stu90]. ed. [Obr98]. edition [Hol99, Gol98, Kaf99b, Hol99, Joh99, KK99b, Mac99]. Editor [ATHM^{+}94, Ano97l, Ano98o, Ano99f, Ano99m, BNL^{+}94, BHR^{+}94, BCUF96, BJK^{+}90, BRG^{+}91, BGS^{+}94, BGD^{+}92, CHD^{+}95, CIM^{+}95, DSS^{+}92, FJH^{+}90, Gat90, Hig99b, Hom98, KJM^{+}91, LS97a, Lan98, LOS^{+}96, LFL^{+}93, MJ99, ML99, MOS^{+}98, MDCW99, RNB^{+}96, RON93, RCU^{+}93, SCHI95, SFP^{+}90, SGO^{+}96, SNK^{+}93, SGN^{+}92, SBS^{+}95, TDF^{+}92, WTA^{+}90, vSMC^{+}91, Ano92i, Ano95j, Ano95k, Ano95l, Ano96i, Ano96j, Ano96k, Ano96l, Ano98c, Ano98q, Ano99g, Dal90a, Ehr90, Gol91, Gol92b, Gol92c, Gol92d, GV93, Gol93a, Gol93b, Gol93c, Gol94a, Gol94b, Gol94c, Gol95a, Gol95b, Gol95c, Gol95d, Gol96a, Gol96b, Gol96c, Gol96d, Gol97a, Gol97b, Hil97a, Hil97b, Hil98a, Hil98b, Hil99, Lan97, Lan99a, Lan99b, Lar94, Lar96, McC90, Nas94a, Nor90, Sch90, Sch91c, Sch92, Sch93b]. Editorial [Ano90l, Ano91j, Ano92h, Ano93j, Ano94g, Ano95i, Ano96h, Ano97a, Ano98a, Ano99a, Gib90, Sch91b]. Education [Ano90k, Hog91, Inm90, MCGM95, RS91a, Sne93, Sne95, Rus98]. Edwards [Boa94]. Effect [Dor96, GR96, Mat99, SRW95]. Effective [Pos93]. Effectively [Bir95]. Effectiveness [MM90]. Effects [BH98, BK96, CJ92, MPB97, SG91, Sea94, Sea97, Alb92]. Efficacy [NS97]. Efficiency [GT95]. Efroymson [Mil96]. Eight [Mor98]. Eisenhart [CR95]. Elasticity [Ass96]. Electric [Pal99]. Electronic [BM94, Sne98, SATV94, Wil91b]. Elemental [MG97]. Elementary [Ano90h, Ber97c, Hu92, Pil98, Sam92, Sho98, GY89, GY90]. Elements [SP98]. Elimination [Edw96]. Elusive [Hel99]. Embracing [Wil94a]. Emerging [HHJL93]. Empirical [Epp93, Mon95]. Engineering [Ano98b, Ano98t, BN98, Bis98, Czi98, Hog94, SP98, Spu98, Var98, ZH90]. Engineers [Bis91, Cle99d]. Enumeration [Hog92a]. Envelope [SL92]. Environment [HHJL93, LS90, RKW98]. Environmental [SG91]. Environments [Nag98]. Epidemiologic [Hsi91]. Epidemiology [Kle95, Ost98]. Equating [Ano96g, Wal95b]. Equation [HL99]. Equations [Bar99]. Equivalence [LW90, ML92a, ML92b, St.90]. Equivariant [Bas91]. Error [Ass96, Far90, Gat95, GAR92, Kaz95, LK99, Mur94, Wai96]. Errors [Bas97, CR96, Str90]. Errors-in-Variables [CR96]. Establishing [AW91b]. Estimable [EGCD91]. Estimate [Kva96]. Estimates [Ano93h, BC92, Bly93, EM93, KG95, SK93]. Estimating [AW91a, AH95, Bor97, Hau97, HL99, Hou97, KU96, Per98, Rus91, Sch91a, Sol93, Str90, Stu90]. Estimation [AC98a, Ano90q, Ano95h, BR97, BGG98, Bro93, Bro94, DC90, Far90, Hen99, Kor90, LK99, ME95, MS94, NWS90, PB92, PT91, PR94, Pul90, War97]. Estimator [Ang92, Oak93, SI90]. Estimators [Bas91, Buz97, CHS99, DN94b, DN94c, FK97, Kha92, LT96, Lin93, Nay94, SC94, DN94a]. Ethnicity [Saw98, Wal98]. Evaluating [LC96]. Evaluation [Coh97, Lac90]. Even [Vid94]. EVENTCHARTS [Gol92a]. Events [Gol92a, WSS93]. Evidence [Luc97, She98]. Evolving [CY99]. Exact [AC98a, BK96, Meh91, MW92, MH93, Nic94b, Wil91b, ZCM95]. Exam [Kva96]. Examination [Ost98]. Example [AFV96, Hos92, Sil97, Tur97]. Examples [Beh90, KB93, KG95, NWS90, OS95]. Excel [Cle99c, Bon98, Sch99]. Excerpts [Pet99]. Exchange [Inm94, PFI94, CU92]. Exclusion [Ham97, HR98]. Exercises [Cle99c, KM92]. Expect [PP98]. Expectation [PP98]. Expected [Hom97, SSC99]. Expenditure [CJ92]. Experience [Bra96a, CRV95]. Experiment [AC98b, PR98, PF95, Rie90]. Experimental [Gat95, GM96, Gun93, HHB91, Par98b]. Experiments [Ber96b, Czi99, Dan94, DeL94, Fle94, Kut94, Lam94a, Law96, LDL^{+}94, NL95, Pie90, PF95, Rog94, Sal99, Sul96, Ver99, Ank98, Cot98]. Explained [KS91]. Explaining [CG92, GC99]. Explanation [GKR94, PG97]. Exploration [She97]. Exploratory [Atk95]. Exploring [HBC^{+}91, HH95, OH95, Rum99]. Explosion [GT95]. Exponential [BL94, CM92, MS94, She97]. Exponentiated [Gre96]. Exposing [GM96]. Exposure [Dan92]. Exstatix [BM91]. Extensions [JK90, NWS90]. Extrapolation [But90, Van96b]. Extreme [CT91, Hsi95, SB98].

F [ML92a]. Faces [Str92]. Factor [Ber96b, Czi99, Dan94]. Factorial [Bec96, Dan94, DeL94, Fle94, Kut94, Lam94a, LDL^{+}94, Nav97a, NL95, Rog94]. Factorials [LL97]. Factors [DeL94, Fle94, Kut94, Lam94a, LS99a, LDL^{+}94, Rog94, Sea94, Sti95]. Fail [AS94]. Failure [MZ92]. Fall [LW98]. Fallibility [FBG96]. Families [CM92, Fer90]. Family [MS94, She97]. Fast [Buz97, Unw91]. Fastat [BM91]. Fastest [Ano97f, Tib97]. Fate [Gun98]. Faulty [Ste92]. Faulty-Inspection [Ste92]. FDA [DuM99a, DuM99b, LS99b, Mad99, OS99]. Features [JR92]. Federal [Gon94, Eld90, Wal98]. Feel [PF95]. Fiducial [Sch96c]. Fig [Lam90]. File [Hil96a]. Filter [CT91]. Fin [MCGM95]. Financial [KR99]. Finding [EGCD91, Hel99]. Finite [DC90, Hau97, Hou97]. FIRM [Haw91]. First [AC98b, Dav95, Dav98a, Hay96, Pil98]. Fisher [Inm94, McL99, PFI94].

Fit [GKT91, HL99, KS91, Sim98]. Fitting [Gan95, HD94, Mad98]. Five [CS94, KDV94, Loc93, Lur95, Mar94, Ost98]. Five-Pin [CS94]. Fixed [Lee93, Sea94]. Fixed-Effects [Sea94]. Flipper [Kon95]. Florida [CRV95]. Folded [Mon95]. Follow [LKNP95]. Follow-Up [LKNP95]. Football [Ste91]. Forecasting [OL96]. Foreign [Eng93]. Foreword [Tuc95]. form [Knü94]. Formal [Haw91]. Forms [DCWP93, Dri99, Saw92]. Formula [Ano96g, Hut93, LC99, Wal95b]. Formulas [DCWP93, Jon99, Pei93]. Formulated [Ano91i, Pei90]. Formulation [Olk92, Smi95]. Formulations [LW90]. Foundations [Sti99]. Four [BM94, KB93]. Fourth [DR99]. Fractional [Bec96, McL99]. Framework [HG97]. Frank [NL95]. Freedom [AW91a, Lev97, Sul96]. Frequencies [Nun92]. Frequency [DuM99a, DuM99b, LS99b, Mad99, OS99]. Frequentist [Ber96c]. Front [Ano90m, Ano90n, Ano90o, Ano90p, Ano91k, Ano91l, Ano91m, Ano91n, Ano92j, Ano92k, Ano92l, Ano92m, Ano93k, Ano93l, Ano93m, Ano93n, Ano94h, Ano94i, Ano94j, Ano94k, Ano95m, Ano95n, Ano95o, Ano95p, Ano96m, Ano96n, Ano96o, Ano96p, Ano97g, Ano97h, Ano97i, Ano97j, Ano98j, Ano98k, Ano98l, Ano98m, Ano99h, Ano99i, Ano99j, Ano99k]. FU [LKNP95]. FU-PLOT [LKNP95]. Function [AT99, Cai94, Cha95, Gro91, Koo93, McC94, Mon95, Pos93, Sch91a, Wal95a, vM93]. Functions [Ano96g, DCWP93, EGCD91, Epp93, Jor94, Luc97, Ren92, WTH95, Wal95b]. Funding [Der93]. Funnel [Gun93]. Further [Luc97]. Future [Bar90, Bra90, Ehr90, Goo90, HHHZ99, Ham90, RSW99, ZH90]. Futures [Bax90].

G [DeL94, Fle94, Kut94, Lam94a, LDL^{+}94, Rog94]. GAIM [S.N91]. Gambling [CW93]. Game [Chu99, Ste91, Wes90, Kel94]. Games [Pau92]. Gamma [Wie99]. Gastwirth [MD97, PG97]. Gauss [Ano92g, Jes94, Sha91a]. Gaussian [SS91]. Gendered [Ham97]. General [HH99, Mor98, Pes90, Sal90, Stu90, TM95, DK95]. Generalized [AT99, BC92, Bir90, Boo92, Chu99, HD94, HL99, Oks91, Pie92, S.N91, Tju98, ZPH99, Hil93, Hil94]. Generate [Daw96, MD96]. Generating [BS95, Cai94, EP91, Gan95, Koo93, Lee93, McC94, PPS96]. Generation [DAF^{+}94]. GENMOD [KDV94]. Genstat [Ano92f, Zur91]. Geometric [Bri96, Sam92, Sch93a, Ste96, TO93]. Geometrical [Bai92, Epp93]. Geometry [GM96, SK95]. Gibbs [CG92, MB94]. Give [Wie99]. Given [MS94]. Glimpse [HHHZ99]. Global [HBC^{+}91]. Go [Ver99]. Goals [Saw98]. Gods [Jam99]. Going [MH93]. Good [Ver99]. Goodness [KS91, Sim98]. Goodness-of-Fit [Sim98]. Got [VAD90]. Government [Ros95a]. Graduate [BK93]. Graphical [DGC97, Fra92, Gun98, KAB^{+}97, Lee95, LKNP95, MC96, Nav97b, PS94, Sti95, Wes90]. Graphically [HH95, OH95]. Graphics [Atk95, Edl91, Haa90, HBC^{+}91, MMCsS96, Mar94, Nas94b, Unw91]. Graphing [Bar98, Hyn96, McN92, NO98]. Graphs [CW99]. Graybill [St.90]. Group [DS98, Eme96, SS98]. Groups [Bir95, OW91]. Growth [Lor90]. Guide [Cle99c, Hoe98, Mur97, Par98b, Pet99].

H [TW91]. Haenszel [Lev97, MW92]. Hald [PR98]. Half [LC99]. Hand [War95b]. Hansen [TW91]. Harmonic [Ste96]. Hasse [Loh95]. Hastings [CG95b]. Having [Lee93]. Health [DSJV90]. Her [Mur94]. Here [SATV94]. Heredity [Nel98a]. Hesitate [Alb97a, Alb97b, Ber97a, Ber97b, Fre97, Lin97, Moo97a, Moo97b, Moo97c, Sch97, Sho97, Wit97a, Wit97c]. Heterogeneous [Bir95]. Hierarchical [CS91, SC99]. High [EP91, HK92]. High-Dimensional [EP91]. Higher [ZCM95]. Highest [Hyn96]. Hindsight [Bos90]. Histogram [BfJTH93, Wan97, Wan99]. Historical [Ano90k, Haa90, Inm90]. History [Ano91h, CR90, DSJV90, Fre90, Gol90, HBF90, Hes90, LS90, Rot90, MMR90]. hitters [Alb92]. HLM [KDV94]. Hold [Bar90]. Home [HS94, Alb92, BH98]. HOME-AWAY [BH98]. Home-Court [HS94]. Homogeneity [MS92]. Honeycomb [Vid94]. Hope [Ehr90]. Horvitz [OS95]. Hot [War95b]. Hypotheses [GKR97a, Sal90, Sil97]. Hypothesis [All97, Ano93h, GR96, Loo97, SK93]. Hypothesis-Dependent [Ano93h, SK93]. Hypotheticals [GKR97a].

IDAMS [Vis91]. Identities [Bai92]. Identity [MS95, Wri92]. Ignoring [AFV96]. II [BS95, BM91, McC99]. III [CH96]. Illustrate [Pau92]. Illustrating [Jes94, Tur97]. Illustration [Jol95, Mur94, Nav97b]. Illustrations [Nas92]. Image [Yu94a, Yu94b]. Imbalance [Kaz95]. Impact [Bor99, Hah95, HHHZ99, Ham97, HT90, Kaz95, LC96, SG91]. Implementation [SR96]. Implementing [RKW98]. Importance [GM96]. Important [JR92]. Imposing [AR95]. Impossible [GKR94, GKR97b]. Improved [Car96, Dri99, Lev93]. Improvement [HHHZ99, KM92, Hoe98]. Improvements [Hog91]. Improving [Der93, Far90, Gon94, SW90]. Including [Nas94b]. Inclusion [Sol93]. Incorrect [Pei93]. Increasing [CF91, Mag96]. Independence [CA97, JT96, LaM94b, SK95, WSS93]. Independent [LC99, Zeh91]. Index [Ano97k, Ano98n, Ano99l]. Indexe [Ley96]. Induction [Gho96]. Inductive [SS91]. Industrial [Hah95, HHJL93, Hun98, MHT90, RKW98]. Industrial/Chemical [RKW98]. Industry [Dan92, Ket95, KM92, McD99, ST95]. Inequalities [SS97, Ste96]. Inequality [Cua93, MN97]. Inertia [Nel98b]. Inference [Epp93, HM93, Haw91, HT90, LS97b, Meh91, NS97, RS93, Rum99]. Inference-Based [Haw91]. Inflation [Sti95]. Influence [Gro91]. Information [Ano90t, Ano91s, Ano92n, Ano93q, Ano94l, Ano95t, Ano96u, Ano97r, Ano98u, Ano99o, BK93, Car96, Chu99, Day98, Don99, GT95, Igl98, McL99, Peñ97, PW91, Sne95, TM99]. Informative [DBD90]. Inning [RMY96]. Input [Lar92]. Insight [Fra92, Gun93]. Inspection [Ste92]. Institutional [Gib90]. Instructor [Kon95]. Instructors [LW98]. Instruments [NO98]. Integrated [HG97]. Integrating [BO92, SO94]. Interaction [SG91]. Interactive [Atk95, Gol98, Unw91]. Interactively [Dun99]. Intercept [GAR92]. Internally [GW94]. Internationally [Vis91]. Internet [Sne95]. Interpretation [Bai92, Ber96c, PS98b, Ren92, Sti95, TO93]. Interpretations [Epp93]. Interpreted [Ale91]. Interpreting [Cua93, Pet99]. Interrater [Mac93]. Interstates [LC96]. Interval [AC98a, Ano95h, Ber96a, Bro93, Bro94, Ede90a, LT96, LK99, SL92, VW93, War97]. Intervals [Ano90i, Ano93i, BK96, Fer90, Juo93, LD94, MW92, Pes90, Pol99, TL93, Van96a, Var92, Ver99, WL99, Ang89, Ang90, Khu93]. Introduction [Alt92, Ank98, Hay99, Kut95, Law96, Rum99, Wue98]. Introductory [BO92, Bra96a, DNG98, Gab91, Led95, Mag96, Ste90, SR96, Wat91, WC97]. Invariance [Neu92, PB92]. Inverse [SS91]. Inversion [WTH95]. Invited [Boe94, Dal90a, DeL94, Ehr90, Fie94, Fle94, Gro94, Gun94, Hod94, Hog92b, HSF^{+}94, Kad94, Kaf94, Kut94, Lam94a, LDL^{+}94, Loh94, Min92, Nor90, Rog94, Sch90, Sch91c, Sol92, Sti94, Ver94, Wel92]. Involving [Fer90, Sil97]. Iowa [LC96]. Iterative [Gan95, HD94, Jor94].

Jackknife [Buz97]. Jackson [Ben93]. Jai [CG95a]. Japanese [Git94]. Jeffreys [LS97b]. JMP [BM91]. Job [DWBQ98]. Jobs [DWBQ98]. Joe [Ben93]. Joint [AS98, Mag90]. Journal [Pet99, Sne95]. Journals [Gib90]. Jumping [Han92b]. Just [Lar92, Lin92]. Justification [Gre91, HD94].

Kalman [CT91]. Kaplan [Oak93]. Karl [Inm94, PFI94]. Kendall [VM90]. Kernel [Alt92, Ano90q]. Key [GT95]. Kinetic [NQ96]. Known [AH95, LC99, Tur97]. Krieger [MD97, PG97]. Kronecker [MS98]. Kurtosis [Gro98, SI90].

Labor [SG98]. Laboratory [Ano98b, Ano98t, BN98, Bis98, Cle99c, Czi98, Spu98, Var98, Cry91]. Laboratory-Based [Ano98b, Ano98t, BN98, Bis98, Czi98, Spu98, Var98]. Lack [HS91]. Lagrange [Wri92]. Language [Ale91, Ham97]. Large [DuM99a, DuM99b, FT93, HS94, JR92, Koo93, LS99b, Mad99, Mag96, MC96, OS99]. Larger [Bai95]. Laws [Koo93]. Lead [Vid94]. League [Gro90, HV95, RMY96, WC97]. Learn [Ros96a, Wat91]. Learning [Die93, Moo93]. Least [Bas97, BK92, BC92, Dal91, Dun94, HH99, HS92, Lic90, ML92a, ML92b, Oks91, Rie90, Van96b]. Least-Squares [Lic90]. Legal [Gas92]. Lehmann [Suk92]. Lesson [BK92]. Lessons [Mos97]. Let [Hog99a, MCDD91a]. Let’s [Bry99b, Cob99b, Gro99b, Hig99a, Hog99b, MCDD91b, New99, Sey91]. Letter [McC90]. Letters [ATHM^{+}94, Ano97l, Ano98o, Ano99f, Ano99m, BNL^{+}94, BHR^{+}94, BCUF96, BJK^{+}90, BRG^{+}91, BGS^{+}94, BGD^{+}92, CHD^{+}95, CIM^{+}95, DSS^{+}92, FJH^{+}90, Hig99b, Hom98, KJM^{+}91, LS97a, LOS^{+}96, LFL^{+}93, MJ99, ML99, MOS^{+}98, MDCW99, RNB^{+}96, RON93, RCU^{+}93, SCHI95, SFP^{+}90, SGO^{+}96, SNK^{+}93, SGN^{+}92, SBS^{+}95, TDF^{+}92, WTA^{+}90, vSMC^{+}91]. Level [LL97, Pro91].

Leverage [Sal90]. Library [Kah91]. Life [DB97, SC99]. Life-Testing [DB97]. Life-Time [SC99]. Light [Daw96, MD96]. lightly [Wio90]. Likelihood [Ass96, BC92, Bir90, DN94b, DN94c, Hen99, Kor90, Mag90, ME95, PB92, PT91, Sil97, Ste92, Tju98, DN94a]. Limit [Jol95, LC96]. Limitations [MGLK99]. Limited [TM99]. Limiting [DCWP93]. Limits [Sch96c]. Line [Wri92]. Linear [BR97, CR96, DB97, EGCD91, GAR92, Hil94, HD94, HT90, Ibr97, Kot91, LaM94b, LW90, MN97, MS95, MV94, MSS91, Mol91, MPB97, Nic94b, Pie92, Sal90, Sea97, SM91, TM95, Tju98, Van97, WB91, ZPH99, vS96]. Lines [Nel90, Nel98b]. List [Dav98a]. Literacy [GNM^{+}98]. Literate [Sch90]. Literature [Ham97, Mur97]. LMS [Dal91]. Local [HBC^{+}91]. Locate [Hel99]. Locating [HBC^{+}91]. Log [Mol91, Pie92, Wie99, vS96]. Log-Linear [Mol91, vS96]. Log-Normal [Wie99]. Logical [Gre91]. Logistic [Bra96b, HHB91, OW91, Sim98, Str92, WL99]. Logit [Sap91]. LOGLINDI [Mol91]. Lognormal [Rea98]. Longitudinal [Dun94, LKNP95]. Look [Dan92, HV95, JL97, Kot91, RvE97]. Looking [Gna90]. LORANK [CGO91]. Lottery [Pau92]. Lotus [Kah91]. Lower [CGO91, Gre96, Lev97].

Macintosh [BM91, SMP97]. Magnitudes [Sch93a]. Major [HV95, RMY96, WC97]. Majors [GH91]. Make [Dav98b, MCDD91a, MCDD91b, Sey91, Wat91]. Making [Mon97b, Rum98]. Man [Ano97f, Tib97]. Management [Bli90, Git94, Vis91, Obr98]. Managers [Bon98]. Mann [Ahm96]. Mantel [Lev97, MW92]. Manual [Kha91]. Many [Daw96, MD96, Str92]. Map [SM97]. Maple [Bag95, Tan99]. Margin [SS99]. Marginal [Lee93]. Marginals [JH96]. Market [SG98]. Marketing [Rot90]. Markov [Ano92g, Jes94, KCGN98, RM91, Sha91a]. Mass [DCWP93]. Matched [Lev97]. Matches [RvE97]. Matching [RS91b]. Materials [Ano91a, Ano91b, Ano92a, Ano93a, Ano95a, Ano95b, Ano95c, Ano96a, Ano96b, Ano97m, Ano97n, Ano97o, Ano97p, Ano98i, Ano98p]. Mathematica [CB95b, Bag95]. Mathematical [Dav95, Mac99]. Mathematics [GH91, HK92]. Matrices [CGO91, KAB^{+}97, MS98, MC96, RM94]. Matrix [Ale91, McL99, Olk92, Ale91]. Matter [Ano90c, Ano90d, Ano90e, Ano90f, Ano90m, Ano90n, Ano90o, Ano90p, Ano91c, Ano91d, Ano91e, Ano91f, Ano91k, Ano91l, Ano91m, Ano91n, Ano92b, Ano92c, Ano92d, Ano92e, Ano92j, Ano92k, Ano92l, Ano92m, Ano93b, Ano93c, Ano93d, Ano93e, Ano93k, Ano93l, Ano93m, Ano93n, Ano94a, Ano94b, Ano94c, Ano94d, Ano94h, Ano94i, Ano94j, Ano94k, Ano95d, Ano95e, Ano95f, Ano95g, Ano95m, Ano95n, Ano95o, Ano95p, Ano96c, Ano96d, Ano96e, Ano96f, Ano96m, Ano96n, Ano96o, Ano96p, Ano97b, Ano97c, Ano97d, Ano97e, Ano97g, Ano97h, Ano97i, Ano97j, Ano98d, Ano98e, Ano98f, Ano98g, Ano98j, Ano98k, Ano98l, Ano98m, Ano99b, Ano99c, Ano99d, Ano99e, Ano99h, Ano99i, Ano99j, Ano99k]. Maximal [GK99]. Maximum [BC92, DN94a, DN94b, DN94c, GW94, Hen99, Kor90, LC96, ME95, PB92].

May [CW93, Dav98b]. Mean [AH95, AS98, BS99, GAR92, Lev93, MC91, O’C90, Per93, Sam91, She97, SP94, Ste96]. Meaning [Sen90]. Means [Ano93h, BC92, GR96, MS92, SK93, SF93, vM93]. Measure [MN97]. Measurement [LK99]. Measures [DGC97, GKT98, Gro98, Hos92, Mag90]. Measuring [AG95, Saw98, vS96]. Media [Dan92, GNM^{+}98]. Median [Dal91, HS92, Lee95, O’C90, SGC90, She97, SP94]. Medical [CJ92]. Medium [CW99]. Meet [MHT90]. Meeting [Saw98]. Meier [Oak93]. Memorian [NL95]. Message [CF91, CW99]. Meta [Van96b, Nor95]. Meta-Analysis [Van96b, Nor95]. Method [BK96, EP91, Faw90, GCGG95, HS92, LKNP95, Oeh92, PPS96, Pes90]. Methods [Alb93, Bon98, Cle99d, Die93, Hes90, MV94, Mur97, Joh99, And99, Eme96]. Metropolis [CG95b]. Microcomputer [HG97]. MicrOsiris [VV90]. Microsoft [Bon98, Cle99c]. Mid [MW92]. Mid- [MW92]. Milliken [St.90]. Minimizes [SGC90]. Minimum [Ang92, Cai94]. Mining [DuM99a, DuM99b, Han98, LS99b, Mad99, OS99, PW91]. Minitab [Cle99c, Alb93, Ste90]. Misleading [VW93]. Missing [PT91, Sne93, HB96]. Missingness [MGLK99]. Mississippi [BES98]. Misuse [CR96]. Mixed [MSS91, MPB97, Sch93d, Sea97, Vos99]. Mixed-Effects [MPB97]. Mixed-Model [Sch93d]. MIXSOFT [Pie90]. Mixture [Beh90, FN92, Pie90, PR98]. Mixtures [Hau97, Hou97, Sap91]. ML3 [KDV94, Pro91]. MLE’s [SS91]. Mode [AG95, Sen90]. Model [AR95, And94, Bri96, HT90, Kot91, Sap91, Sch93d, Stu90, TM95, Van97, Alb92]. Model-Based [Kot91]. Modeling [Car96, Dan92, GPA97, Haw91, NWS90, NQ96, Peñ97, RKW98, RMY96, SB98]. Models [Ano91i, BL94, BF91, CR96, CM92, EGCD91, Fri95, GKT91, Hil93, HD94, HL99, Ibr97, LaM94b, MN97, MS95, MV94, MSS91, McL99, MPB97, NS97, Nel98a, Pie92, S.N91, Sch99, SMH91, SSH96, Sea94, Sea97, Sha91b, Tju98, Vos99, Wie99, ZPH99, vS96, Pei90, Hil94]. Modern [KK99b, Pil98, Rum99, SP98]. Modernizing [Leh95, Gar95, Mor95a]. Modification [Lev93]. Moment [Cai94, DR99, Koo93, McC94, Nav97a, PN95]. Moment-Generating [Cai94, McC94]. Moments [BSY92, Hos92, Nel98b, Sam92, Smi95]. Monotone [Ste92]. Monotonic [Bas91]. Monte [BS98, KCGN98, Rie90, RM91]. Morris [TW91]. Most [Bir90, Ede90b]. Motivating [DNG98]. Mountain [Mon95]. mph [LC96]. MS [Mar94, Ano90j, Obe91]. MS-DOS [Mar94, Obe91]. MS/PC [Ano90j]. MS/PC-DOS [Ano90j]. MSOD [Ehr90]. Multi [Bor99]. Multi-Tiered [Bor99]. Multicenter [WKB^{+}95]. Multilevel [KDV94, ZPH99]. Multimedia [VM96]. Multinomial [Ano90i, Ano95h, Bro93, Bro94, KU96, Ang89, Ang90]. Multiple [Cua93, Fra92, LD94, Nic94b, PF95, Sav90, TM99, WW97, Wil91a]. Multivariable [Joh99]. Multivariance [FB91]. Multivariate [Ano90q, CY99, EP91, FN92, Gan95, Loo95, Sap91]. Muncher [Mon97b].

Musical [MBB90]. Mutual [SK95]. Myths [BKM96].

N [Ano90q, Bra90]. N-KERNEL [Ano90q]. Napoleon [Gun98]. National [CJ92]. Natural [OS95]. Nature [BGG98, Inm94, PFI94]. NBA [CY99]. NCAA [Car96, SMH91, SSH96, SS99]. Nearest [Alt92, SM91]. Nearest-Neighbor [Alt92]. Nearly [JH96, Nel90]. Nearness [Kha92, Nay94]. Necessary [McL99]. Necessity [ST95]. Need [HG97, Hel99]. Needed [Hog91]. Needle [Rie90]. Needs [Bic95, Ket95, MHT90, McD99, Ros95a]. Negative [VBB94]. Neighbor [Alt92, SM91]. Nested [DeL94, Fle94, Kut94, Lam94a, LDL^{+}94, LS97b, Rog94, Sea94]. Networks [Pal99, WM96]. Neural [WM96]. News [DWBQ98]. Next [Bra90, DAF^{+}94]. Nigerian [Oko99]. NLEX [NWS90]. No [Lin92, Ang90, DN94a, GY90, WS94]. Nominal [VW93]. Noncalculus [SGC90]. Noncentral [Cha95, Pos93]. Nongaming [Bor97]. Nonidentifiable [NS97]. Nonlinear [Ano90q, NWS90, Tju98]. Nonlocal [Mag98]. Nonmathematical [Hig99c]. Nonparametric [Alt92, DS98, Meh91, SS98]. Nonrandom [MGLK99]. Nonsingular [McL99]. Nonstandard [vS96]. Nonstatisticians [Bra96a]. Nonuniform [Nun92]. Normal [Ano96g, AS98, BS99, BfJTH93, Cai94, CP99, GM91, Gre96, Hut93, JH96, Sap91, SF93, Van96a, Wal95b, Wie99, McC90]. Normality [DBD90, Loo95]. North [NOMLT95]. Note [Ang89, Ang90, Ano90k, Ano90i, BR97, Cha95, CH96, Dav93, Don99, Ede90b, EM93, GM91, HL96, Hen99, HS92, Hsi95, Inm90, Koo93, LaM94b, LK99, Oak93, Oeh92, PT91, Pes90, Pul90, SI90, Ste96, TL93, Van97, Khu93]. Notes [Ano92i, Ano95j, Ano95k, Ano95l, Ano96i, Ano96j, Ano96k, Ano96l, Ano98q, Gol91, Gol92b, Gol92c, Gol92d, GV93, Gol93a, Gol93b, Gol93c, Gol94a, Gol94b, Gol94c, Gol95a, Gol95b, Gol95c, Gol95d, Gol96a, Gol96b, Gol96c, Gol96d, Gol97a, Gol97b, HL98, Hil97a, Hil97b, Hil98a, Hil98b, Hil99, Nas94a]. Number [Ang92]. Numbers [Hel99, Koo93]. Numerical [Luc97, WTH95]. Numerically [Wal95a].

Objection [MD97]. Observation [Olk92]. Obtaining [Pes90, Smi95, TL93, WTH95]. Occurrence [Dav95, Dav98a]. Odd [Ang92]. Odds [Gab91, Han92b, MW92, WL99]. Old [Sen90, Smi95]. OLS [BF91]. Olympic [NK96]. Once [LC99]. One [Ano98b, Ano98t, BN98, BK96, Bis98, Czi98, Czi99, Dan94, LT97, Lev97, LC99, O’C90, Sil97, Spu98, Var98]. One-Dimensional [LT97]. One-Factor-at-a-Time [Czi99, Dan94]. One-Half [LC99]. One-Semester [Ano98b, Ano98t, BN98, Bis98, Czi98, Spu98, Var98]. One-Sided [Sil97]. One-Way [BK96]. Online [Hel99]. Operator [BSY92]. Opportunities [Gna90]. Opposition [Dav98b]. Optimal [CB95a]. Optimality [Lee95]. Order [Ano90h, BR97, Dav93, DN94b, DN94c, SC97, DN94a, GY89, GY90]. Order-Statistics [Ano90h, GY89, GY90]. Orders [KR99]. Ordinal [VW93].

Oriented [Alb97c, Ano98b, Ano98t, BN98, Bis98, Czi98, Spu98, Var98]. Orthogonal [CR96]. Other [Gol92a, Kir91, Lin92, SG98, Sne95, Var92, Ano93i]. Our [Bry99b, Cob99b, Gro99b, Hig99a, Hog99a, Hog99b, New99]. Outcome [GAR92, GM96, Mat99, SRW95]. Output [OCE93, Sea94]. Outreach [KM94]. Overdispersion [Tju98]. Overruled [MD97]. Overview [Hog92a].

P [Lam90]. Package [Bli90, CGO91, DS98, Meh91, S.N91, Sea94, SS98, Spi90, Tho95]. Packages [Coh97, Dal90a, Dal90c, Eme96, Hau97, Hou97, HF96, Lur95, Mor98, OCE93, Ost98, Sch93d, ZPH99, Loc93]. Paired [DG97, Day98, McN92, MC91]. Paired-Comparisons [Day98]. Pairing [DG97]. Pairs [Dou93, EGS96, Gre96]. Pairwise [SK95]. Paper [Sto98]. Paradox [AFV96, CU92, War95b]. Paradoxical [Hom97]. Parallel [Nel90]. Parameter [Dou93, LT96, MKV92]. Parameters [DN94b, DN94c, Fer90, RS93, SS91, TL93, War97, WB91, DN94a]. Part [Eng93, McC98a, McC98b, McC99, Sch93d]. Partial [KR99, TO93]. Participation [Gel98, Mag96]. Partitioning [CS91]. Past [RSW99, ZH90]. Pathological [KB93]. PC [Mar94, Coh97, Dal90b, FB91, VV90]. PC-Based [Coh97]. PC-DOS [Mar94, Ano90j]. PC-SIZE [Dal90b]. Pearson [Inm94, PFI94]. Peculiar [Ley96]. People [Kir91]. Percentage [Wil91b]. Percentile [Pol99]. Percentiles [Dun91]. Perfect [Mad91]. Performance [GAR92, RMY96]. Period [PS94]. Permutation [LD98]. Perot [Gun98]. Personal [MBB90, Obe91]. Perspective [Bai95, OS95, SG92]. Perspectives [RKW98]. Ph.D [Gar95]. Ph.D. [Leh95, Mor95a]. Phenomena [GPA97, NQ96]. Phrases [GT95]. Physical [Old95, ZH90]. Pin [CS94]. Pinch [Dun91]. Pitcher [RMY96]. Pitfalls [VM96]. Pitman [FK97, Kha92, Nay94]. Place [Moo93]. Placement [RSW99]. Player [WC97, MCDD91a, MCDD91b, Sey91]. Players [Lac90]. Playoff [BH98]. Playoffs [Bor99]. Plot [FM92, HN98, LT97, LKNP95]. Plot-Density [HN98]. Plots [GKT91, HN98, KR99, Mon95, Nel90, Sal90, SW94, Wil99]. Plotting [Bec96, WT92]. Plus [Hal92, KK99b]. Pocket [Hoe98, McC90]. Point [Car96, GR96]. Points [FT93, Wil91b]. Poisson [GY90, Alb92, Ano90h, Ano96g, GY89, Jol95, Kel94, Knü94, Neu92, PA91, Tie94, Wal95b, War97]. Policy [Saw98, Kaf98]. Polynomial [Ano91i, Mag98, Pei90]. Pooling [Hin96]. Population [DC90, Kha92, Nay94, RM91, SP94]. Position [BTV90]. Positive [PN95, VAD90]. Possibilities [But90]. Possible [Bly93, PG97]. Post [Hog92a]. Post-Enumeration [Hog92a]. Posterior [Ros96a]. Power [Ano90j, DP91, Kor90, MH93, Ott96, PA91, SG91, She97, TM95, WB91]. POWER.FOR [DP91]. Powerful [Ber96a, Bir90, DBD90, Ede90b]. Powers [Suk92]. Pr [HL96]. Practical [Ano90q, Hoe98, Huz99, NK96, Par98b, Sav90]. Practice [Can98, KCGN98, Maa99, Per98, SB98]. Pragmatic [GM96]. Pragmatics [Hin96]. Precise [Dun91]. Predict [SS99]. Predicting [BCJ97]. Prediction [SM91, Knü94]. Predictive [Ibr97]. Predictor [Ver99]. Predictors [DB97, Sea97, Sim98]. Preliminary [Gol98, MM90]. Preparing [HK92]. Preposterous [Han92b]. Present [RSW99, ZH90]. Presentation [BKM96, Wes90]. Pretesting [GAR92]. Prevalence [Kva96]. Previous [Kor90]. Principal [BK99, FT93, HL98, JK99, NK96, Ren92]. Principle [Hel95, Nel98a]. Principles [Can98]. Prior [LS97b, Mat99, Ros96a]. priority [DN94a]. Priors [Ibr97]. Probabilistic [GPA97, She98]. Probabilities [Ano95h, Bro93, Bro94, KU96, PN95, Pul90, Tie94, Wil91b]. Probability [Ang92, Ano96g, Bar99, BM94, Bor97, BSY92, Dav98a, DCWP93, GKT91, Hut93, Kaz95, Mad91, OS95, SP98, SMH91, SSH96, Sol93, Ste91, WT92, Wal95b, Kel94]. Problem [BES98, Chu99, Daw91, Day98, Hil96b, Hu92, Mal98, Nun92, Pal99, RS91b, Smi95, Knü94, Rea98, Rie90]. Problem-Solving [Hil96b]. Problems [Dan92, Oko99, Sim94, UE90]. PROC [Mon97b]. Procedure [HD94, ML92a, ML92b, St.90]. Procedures [Ber96c, HS91, Sav90, UE90]. Proceedings [Wat95]. Process [Ano90h, Ber96c, KM92, Neu92, RM91, Sto98, GY89, GY90]. Processes [Per93]. productivity [HBF90]. Products [MS98]. Profession [CF91, Igl98]. Professional [Bor99, Lac90]. Profiles [LRG90]. Program [Ano90q, Dal90b, DP91, Hsi91, WC91]. Programs [BM91, Bry99b, Cob99b, Gar95, Gro99b, Hig99a, Hog99a, Hog99b, KDV94, Leh95, Mar94, Mor95a, Nas94b, New99, Tho95]. Progression [Mad91]. Progressive [BS95]. Project [AC98b, Yu94a, Yu94b]. Projecting [Kor90]. Projection [BK99]. Projections [Nic94b]. Projects [AS99, GNM^{+}98, Led95]. Promises [VM96]. Proof [Ano90h, SS91, SGC90, DK95, GY89, GY90]. Properties [Ibr97, Per93, WSS93]. Property [Ano91i, DB97, Lee95, Neu92, SC97, Tur97, Pei90]. Proportion [RvE97]. Proportional [Gan95]. Proportions [AC98a, Ano90i, Ano94f, WS93, Ang89, Ang90, WS94]. Prospects [Oko99]. Prosper [HHJL93]. Proving [Zeh91]. Pseudorandom [Dor96]. Public [CF91, Kaf98, Saw98]. Publication [SRW95]. Publications [SATV94, Wio90]. Publish [SRW95]. Pursuit [BK99]. Putting [SW90]. Puzzling [Sim94].

Q [GCGG95]. Q-TWiST [GCGG95]. Quadratic [Dri99, Saw92, Sch96c]. Quadratic-Type [Sch96c]. Quality [Ano98b, Ano98t, BN98, Bis98, Boe94, BO92, Czi98, Fie94, GCGG95, Git94, Gon94, Gro94, Gun94, Hod94, HHJL93, Hoe98, HSF^{+}94, Imr94b, Imr94c, Kad94, Kaf94, Kas98, Loh94, Sne90, Spu98, Sti94, Str90, Var98, Ver94, HBF90]. Quality-Adjusted [GCGG95]. Quality-Oriented [Ano98b, Ano98t, BN98, Bis98, Czi98, Spu98, Var98]. Quantal [AR91]. Quantifying [BWM98]. Quantile [Gro98, He97, KR99]. Quantiles [HF96, Sho95]. Quantitative [AR91, CS94]. Quantitatively [Sch90].

Quantities [JSM95]. Quasi [Sal99, Tju98]. Quasi-Experiments [Sal99]. Question [AS94]. Quick [Ano90i, Ang89, Ang90]. Quota [BWM98].

R [Ano91t, Ano91u]. R. [Inm94, PFI94]. Race [Saw98, Wal98]. Racing [SB98]. Random [Ano92g, Beh90, BK96, Cai94, HR98, Lee93, Nav97a, SSC99, Sea97, Sha91a, WSS93, Alb92, HB96]. Randomization [Old95]. Randomly [JK90]. Rank [CGO91, LW90, MV94, Suk92, VM90, WB91]. Rank-Based [MV94]. Ranking [Gro90]. Rate [BGG98]. Rates [Lor90, TM99]. Ratings [Bas97]. Ratio [Bir90, DC90, JSM95, Mag90, MW92, RS93, Sil97, Ste92, VW93, vM93]. Ratios [AB92, WL99]. Reaction [Ste90]. Read [Mur97]. Ready [SATV94]. Real [Ber97c, Dor96]. Realism [WC97]. Really [Gat95]. Realm [Han92b]. Reasoning [And99, Bra96a, Gas92, Hay99]. Reasons [Alb97a, Alb97b, Ber97a, Ber97b, Fre97, Lin97, Moo97a, Moo97b, Moo97c, Sch97, Sho97, Wit97a, Wit97c]. Received [Ano93f, Ano93g, Ano94e]. Reconstruction [Yu94a, Yu94b]. Record [BTV90, Sol93]. Records [NK96]. Recreational [SB98]. Recurrence [DCWP93]. Recursive [Haw91, Smi95, Tie94, Tie98]. Reflections [Hah95]. Region [Nic94b]. Regional [SMH91, SSH96]. Regions [Dou93, Hyn96, ME95]. Regression [Alt92, Ano90q, Ano91i, BCJ97, Bra96b, Bri94, Bri96, CR96, CHS99, Coh91, Cua93, Dal91, DC90, Dun94, EM93, Fra92, GW94, GAR92, HL98, HHB91, Hol99, HL99, Hsi95, HT90, Joh99, JK99, Jor94, KG90, Kot91, LaM94b, LD94, Lev93, Lic90, MG97, MC91, Mil96, Nas92, Nel98b, Nic94b, OW91, Obe91, OCE93, Pei93, Pie92, Sam91, Sen90, Sim98, Smi98, Str92, Tju98, UE90, WL99, Pei90]. Regression-Toward-the-Mean [Lev93]. Regressions [BF91, Mag98]. Regressors [Ano92g, Sha91a]. Reinforcing [CF91, Kha91]. Rejoinder [MCDD91b]. Related [Sha91b, SC97]. Relating [Dri99]. Relation [JK99]. Relationship [Old95, Sch93a, Sha91b, SR97]. Relationships [Ber96b, Jon99, RS91b]. Relative [Far90]. Release [Tan99]. Reliability [Mac93, McC98a, McC98b, McC99]. Remaining [Str90]. Remarkable [Jam99, SG93]. Reminder [FBG96]. Removable [Sal99]. Repeated [DGC97]. Replicated [LE93]. Replications [Ott96]. Reply [Alb97a, Alb97b, Ano98b, Ano98t, Ano99n, Ber97a, Ber97b, DuM99b, GKR97a, GKR97b, Hig99a, Hig99d, Hog99b, Hog99c, Imr94a, Imr94c, Moo97b, Moo97c]. Report [Ano98c, Ano99g, Lan97, Lan99b, Lar94, Lar96, Sch92, Sch93b, Mon97b, Lan98, Lan99a]. Reported [Don99]. Reporting [DuM99a, DuM99b, LS99b, Mad99, OS99]. Representation [Gib90]. Representing [FT93]. Requirements [SL92]. Resample [Ros96a]. Resampling [Sim94, SG92, Wil94b]. Research [AW91b, Hes90, Hol99, LRG90, LD98, Nag98]. Residual [KS91, Nel90]. Residuals [GW94]. Resolution [CU92, CH96]. Resolving [Vos99]. Resource [OS95, Won98]. Resources [Sne95]. Respect [AG95]. Respondent [Gar95, Mor95a, Mor95b]. Response [AR91, DeL94, Fle94, Kut94, Lam94a, MKV92, Nel98a, Rog94, Stu90, Van96b].

Response-Surface [Nel98a]. Responses [CW92, LDL^{+}94]. Restrict [Bly93]. Restriction [DN94b, DN94c, DN94a]. Restrictions [Sea97]. Result [Dri99, SC97]. Results [Bra96b, GK99, Wie99]. Reveals [Wri92]. Revelations [Gun98]. Reverse [Shu91]. Reversion [Sam91]. Review [Agr96, AB94, Ame93, And96, And99, Ank98, Bag95, Bar97, Bar98, Bar99, Bis95, Bol91, Boo97, Boy94, Can98, Car93, Che94b, Che94a, Cho97, Cla93, Cle99a, Cle99b, CW94, Coo91, Cot98, Cry91, Cry97, D’A96, Del97, Del95, Die97, Dob93a, Dob93b, Dob98, Don93, Don96, Dou97, DS91, DN92, Elt96, Eme96, Fra97, Fre94, Fri96, GKM^{+}91, Gar97, Gas97, Gol98, Gor96, Gun97, Gun98, Gup94, Hal92, Hal94a, Hal94b, Hal95a, Hal95b, Hal95c, Han95, Har97, Har95, Hau97, Hay94, Hay96, Hay99, Hei91, Hel99, Her95, Hil93, Hil94, Hil96a, Hoe98, Hoe97, Hof91, Hol95, HL99, Hou97, Jam99, Joh96, Joh99, Jos93, Kaf97, Kaf98, KK99a, Kaf99a, KK99b, Kaf99b, Kas94, Kas98, KT94, Kin95, Koe97a, Koe97b, Koo90, Koo97, KDV94]. Review [Len98, Les95, Lev92, Loc93, Low98, Lur95, Lye96, LLL91, MW95, Maa99, Mac99, Mal94, Mal97, Man91, Mar94, Mcc97, Mid96, Mon97a, Moo96, Mor98, Mor91, Nab97, Nag90, Nag98, Nas95, NO98, Nei96, Nei93, Nic94a, Nic95, Nor95, Obr98, OL96, Par98a, Par98b, Pec97, Pet99, PRSE94, Pic97, Pos90, PD93, Pos94, Pos96, RS95, Ree94, Rei94, RR92, RR93, Ros93, Ros94, Ros95b, Ros96b, Rum98, Rum99, Rus98, SP98, Sap95, Sax94, Sch96b, Sch93c, Sch93e, Sch93f, Sha94, She98, She93, Sho96a, Sho96b, Sho98, Sim96, Smi97, Sne98, Sol95, Son90, Spi90, SMP97, Tan92, Tan95, The90, Tho95, Tom93, War95a, War93, Was93, Wat96, Wat95, Wea97, Wel98, Win95, Wit93, Wit94, Wit97b, Won98, Woo93, Wue98, Cle99c, Gol99, Sch99, Tan99, Hay96, KT94]. Review [KDV94, Nor95, Wat95]. Reviews [Ano90s, Ano91q, Ano91r, Ano93p, Ano95q, Ano95r, Ano95s, Ano96q, Ano96r, Ano96s, Ano96t, Ano97m, Ano97n, Ano97o, Ano97p, Ano97q, Ano98i, Ano98p, Ano98r, Ano98s, Bon98, Hol99, Pil98, Smi98]. Revising [Wal98]. Revisited [Daw91, KK91, SRW95, ZM94]. Revisiting [NK96]. Rho [RS91b]. Ridge [Obe91]. Right [Ama99]. Risk [DNG98, KS91, KR99, MZ92, Jam99]. Road [SB98]. Roads [Vid94]. Robust [Bas97, MKV92, NWS90]. Rocket [MZ92]. Role [Bil98, LaM94b, Old95, SK95]. ROM [Sne98]. Rome [Vid94]. Root [Hut93]. Rosenbaum [MD97, PG97]. Ross [Gun98, Wan91]. Roundtable [KCGN98]. Rudolf [Rie90]. Ruin [Pal99]. Rule [Alb97c, JL97, Puk94]. run [Alb92]. Runners [SB98]. Runs [RMY96]. Rural [GKT98, LC96].

S [KK99b]. S-Plus [KK99b]. Sacco [She98]. Saddlepoint [GC99, Huz99]. Salaries [HV95, WC97]. Salary [Lac90]. Same [Ano94f, WS93, WS94]. Sample [Ama99, Ano90j, Ano95h, Bro93, Bro94, CJ92, Dal90b, Die93, DP91, Ede90a, HR98, Hsi91, HF96, Jol95, KG95, LC99, Mat99, MC91, MS92, NP97, Per93, Ros96a, SL92, She97, SP94, Suk92, TM95, War97, Wri90, vM93, Knü94]. Sample-Size [Dal90b]. Sampler [CG92, MB94]. Samples [Ano90g, Ano91g, BS95, Lev97, Pes90]. Sampling [Alb93, CLM92, DC90, Dor96, FKL^{+}96a, FKL^{+}96b, OS95, SS91, SG92]. SAS [Kas98, AR91, DS98, Mon97b, SS98, UE90]. SAS-Based [DS98, SS98]. Say [Lin92]. Scale [WB91]. Scaled [FK97]. Scatterplot [SG93]. Scatterplots [EGS96, KG98]. Scheffé [Khu93]. Schemes [Tie94, Tie98]. Scholarly [GT95]. Scholarship [Hog92b, Min92, Sol92, Wel92, Wil92]. School [HK92]. Schools [Wat95]. Schucany [Ano91t, Ano91u]. Science [Ber97c, CF91]. Sciences [DSJV90, ZH90, Bon98, Wue98]. Scientific [SO94]. Scientists [SG98]. Score [Boo92, St.90]. Scores [Kva96]. Searching [Ama99]. Second [Dav98a]. Section [Ano90k, CR90, Fre90, Gol90, Haa90, Hes90, Hil98b, Inm90, LS90, ZH90, HBF90, Ano98q, Gol97a, Gol97b, Hil97a, Hil97b, Hil98a, Hil99]. Seeding [SS99]. Seemingly [BF91]. Segmented [BWM98]. Seldom [Don99]. Selected [Igl98]. Selecting [Die93]. Selection [Ano90g, Ano91g, BWM98, HT90, Nel98a, OW91, Sal99]. Selection/Quota [BWM98]. Semester [Ano98b, Ano98t, BN98, Bis98, Czi98, Spu98, Var98]. Semi [RM91]. Semi-Markov [RM91]. Semiconductor [ST95]. Separate [PF95]. Sequences [BH98]. Sequential [Eme96]. Series [BH98, Ben93, CHS99, Cry91, Gab91, KT94, PT91, PA91, Unw91]. Service [AS99, Sne95]. Sesquicentennial [Ano90a, Ano90b, KLM90, MBB90, Nor90, S^{+}90]. Set [Daw96, MD96, Str90, WSS93]. Sets [AS98, FS93]. Setting [Eng93, Gas92]. Seven [BH98]. Seventh [Hut93]. Seventh-Root [Hut93]. Several [MW92]. Shape [Hos92, RM94]. Shaping [McD99]. Shareware [Nas92]. Shoeless [Ben93]. Shoes [AR95]. Shoestring [Wil94b]. Short [Bir95, CRV95, Eng93, Kle95]. Shortest [Fer90, Juo93, WL99]. Should [Ros96a]. Show [Chu99]. showcase [CB95a]. showdown [CB95a]. Shuttle [MZ92]. Sided [Ede90b, Sil97]. Sides [Atk94]. Siècle [MCGM95]. Sigma [HHHZ99, Puk94]. Significance [Mag90, ZCM95]. SIMCLUST [WC91]. Similar [JR92]. Simple [BS95, BF91, BK96, EGCD91, Eub94, GAR92, GM96, HR98, Hel95, HD94, Jon99, Oks91, PPS96, SGC90, Tur97, Yu94a, Yu94b]. Simplified [FS93, VM90]. Simplifying [HH99]. Simpson [AFV96, War95b]. Simulated [Dun91]. Simulation [CLM92, Lin93, MMCsS96, RM91, Rus91, Sim94]. Simulational [BS95]. Simultaneous [Ano90i, KG90, LD94, Ang89, Ang90]. Single [PF95]. Siobhan [Daw91]. Situations [Hom97]. Six [HHHZ99]. Size [Ama99, Ano90g, Ano90j, Ano91g, Ano95h, Bro93, Bro94, Dal90b, Dor96, DP91, Ede90a, GW94, Hsi91, LC99, NP97, SL92, SP94, TM95, vM93, Dal90b]. Size-Biased [Ano90g, Ano91g, SP94]. Sizes [Jol95]. Sketch [Haa90]. Skewness [AG95, Gro91]. Skills [HHJL93]. Slowly [Gun93]. Small [War97]. Small-Sample [War97]. SMD [RKW98]. Smoothing [Eub94]. Social [Bon98, Gol90, Nag98, Wue98]. Software [Ano90s, Ano91q, Ano91r, Ano93p, Ano95q, Ano95r, Ano95s, Ano96q, Ano96r, Ano96s, Ano96t, Ano97q, Ano98r, Ano98s, Coh97, DS98, HG97, HL99, KT94, Mad98, McC98a, McC98b, McC99, Mor98, NWS90, NQ96, Nor95, Ost98, Pie90, Pro91, RM91, SS98, Spi90, SMP97, Vis91, ZPH99, Hil93, Hil96a]. Sole [AR95]. Solid [MZ92]. Solution [Hu92, Nun92, Sav90]. Solving [Hil96b]. Some [Alb97a, Alb97b, Ber97a, Ber97b, Dan92, Dun99, Fre90, Fre97, GK99, Gel98, God91, HL98, Hah95, HV95, JK90, Lin97, Mad98, Moo97a, Moo97b, Moo97c, NK96, Nay94, PR94, RS91b, Sch97, Sho97, Sim94, Wit97a, Wit97c]. Sort [Ver99]. Space [MZ92]. Spatial [HBC^{+}91]. Spearman [RS91b]. Specification [Mat99]. Specified [Lee93]. Speculations [Goo90]. Speed [LC96]. Spherical [TO93]. Spline [Eub94]. Split [FM92]. Split-Block [FM92]. Split-Plot [FM92]. Splitting [PB90]. Spontaneous [DuM99a, DuM99b, LS99b, Mad99, OS99]. Sports [Bas97, Mos97]. Spread [Car96]. Spreadsheets [Wil94b]. SPSS [Mac93]. Square [Dri99]. Squared [Bai92, BK92, GAR92, Lev97]. Squares [BK92, BC92, Dal91, Dun94, HH99, HS92, Lic90, ML92a, ML92b, MS98, Oks91, Rie90, Saw92, SBL97, SR97, Van96b]. SSIZE [Hsi91]. SSR [Sch93a]. Stability [Sto98]. Standard [O’C90, Wal98, vS96, McC90]. Standardize [Bri94]. Standardized [GW94]. Stanford [Sti99]. STAT [Bru91]. Stata [Gol99]. States [Gon94]. Stationary [Per93]. Statist [Ang90, DN94a, GY90, WS94]. Statistic [FBG96, HH95, Lev97, OH95, PS98b, SC97]. Statistical [Ano90k, Ano90r, Ano90s, Ano91h, Ano91o, Ano91p, Ano91q, Ano91r, Ano93o, Ano93p, Ano95q, Ano95r, Ano95s, Ano96q, Ano96r, Ano96s, Ano96t, Ano97q, Ano98r, Ano98s, BWM98, Bon98, Bra96a, CR90, Cle99c, Cle99d, CW99, Dal90a, Dal90c, Dan92, Eld90, ES93, Eng93, Epp93, Gas92, GNM^{+}98, Gon94, GKT98, GM96, Haa90, HBF90, HM93, Ham97, HS91, Hay99, Hog91, HF96, Imr94b, Inm90, Inm94, JSM95, Kah91, KM94, Kir91, Loh95, Lur95, MHT90, MMCsS96, MMR90, MMS90, McC98a, McC98b, McC99, Meh91, Mon97b, Mur97, Nag98, Nas92, Nas94b, New92a, New92b, OCE93, Ost98, Pau92, PFI94, Peñ97, Per98, RKW98, Sam91, Sne90, Sne93, SMP97, SRW95, SB98, Tho95, WM96, WOR98, Wil92, Wio90, Yu94a, Yu94b, ZH90, Boe94, Fie94, Gro94, Gun94, Hod94]. Statistical [Hog92b, HSF^{+}94, Imr94c, Kad94, Kaf94, Loh94, Min92, Sol92, Sti94, Ver94, Wel92, And99, Eme96, Kas98, Spi90]. Statistically [SO94]. Statistician [AR95, Ano98c, Bil98, DeL94, Fle94, Hoe98, Hun98, Kut94, Lam94a, Lan98, LDL^{+}94, Pos90, Rog94, Shu91, WKB^{+}95, Boa94, Lan99a]. Statisticians [Bar90, But90, HHJL93, LRG90, SG98]. Statistics [Ahm96, Alb93, AS99, Ano90h, Ano98b, Ano98t, BR97, Bar98, BN98, Bax90, Ber97c, BM91, Bil98, Bis91, Bis98, BK93, Bon98, BO92, Bra96a, Bra90, Bry99b, Cle99c, Cob99b, CF91, CB95b, Czi98, DNG98, Dav93, Dav95, Dav98a, DSJV90, Ehr90, Gab91, Gar95, GH91, Gel98, Gib90, Gna90, Gol90, Gol98, Gol99, Goo90, Gro99b, Hah95, HHHZ99, Ham90, Han98, HK92, Hel99, Hig99a, Hig99c, Hil96b, Hog94, Hog99a, Hog99b, Igl98, KK99b, Kaf99b, KB93, KG90, LaM94b, Lar92, Law96, Led95, Leh95, LS90, Loc93, LW98, Mac99, Mad98, Mag96, McD99, Moo93, MCGM95, Mor98, Mor95a, Mos97, New99, NS99, OCE93, PS98a, Pet99, Pil98, RS91a, RSW99, Rot90, Rum98, Rum99, Rus98, SP98, SO94, SM97, Sch91a, Sch99, SW90, SG92]. Statistics [Smi98, Sne95, Sne98, SATV94, ST95, Spu98, Ste90, Sti99, S^{+}90, SR96, Var98, VM96, Wat91, Wel98, Wil94a, WC97, Wue98, GY89, GY90, Can98, Low98, Sho98, Maa99, Hay96, Kaf98, Obr98, SP98]. Statistix [Spi90]. STATTAB [BGS91]. Status [Eld90]. StatView [BM91]. StatXact [Meh91]. Stepwise [Coh91, Mil96, OW91]. Sting [Shu91]. Stirling [Ano96g, Wal95b]. Stock [Ley96]. stopping [CB95a]. Story [Jam99]. Straight [Wri92]. Straight-Line [Wri92]. Straightforward [BGG98]. Strategies [CW93, GM96]. Strategy [Loo97]. Stratification [Kaz95]. Streets [SP94]. Strength [Car96, Per98]. Strengths [MGLK99]. Strong [Nel98a]. Stronger [Dav98b]. Structure [DGC97, Van97]. Student [GNM^{+}98, Loc93, Mag96, Ste90, Tan99]. Studentized [GW94]. Students [Ano98b, Ano98t, BN98, Bis98, BK93, Czi98, Hog94, Mur97, Spu98, Var98]. Studies [Hsi91, LKNP95, LE93]. Study [DWBQ98, Kor90, Ste90, Wie99]. Styles [Moo93]. Subjects [Cot98]. Subsampling [MB94]. Subsection [DSJV90, Rot90]. Subsets [MG97]. Success [McD99]. Successes [Ang92]. Sufficiency [PR94]. Sufficient [McL99]. Suggestion [DBD90, SW94]. Suggestions [Bir95]. Sum [SGC90]. Summary [Lar92, OCE93]. Sums [MS98, Saw92, SBL97, SR97]. Sunflower [SW94]. SuperANOVA [BM91]. Superior [LD98]. Support [HR98]. Supporting [Luc97]. Suppression [SR97]. Surface [MKV92, Nel98a, Van96b]. Survey [Ano94f, CLM92, Coh97, FKL^{+}96a, FKL^{+}96b, HG97, Hes90, Hog92a, KG90, KG95, KG98, Kot91, NOMLT95, WS93, WS94, CJ92]. Survival [Atk95, BCJ97, GCGG95, Gol92a, HG97, LW90]. Survive [AS94, HHJL93]. Suspension [Per98]. Symbolic [Ham97]. Symmetric [SC97]. Symmetry [vS96]. Symphony [Kah91]. Synergism [HN98]. System [CY99, DuM99a, DuM99b, LS99b, Mad99, OS99, Kas98]. Systems [Bag95, BWM98, Nas95, Obe91].

Table [MH93]. Tables [BM94, DuM99a, DuM99b, GK99, Gre91, Hin96, Lin92, LS99b, Mad99, MW92, Mon97b, OS99, Wil91b, ZCM95]. TABULATE [Mon97b]. Taguchi [MKV92]. Tail [Jor94, PN95]. Take [Daw96, MD96]. Tales [Gun98, Wel98]. Tape [Pos90]. Teach [Gun93, HK92, Koo93, MMCsS96, Sim94]. Teacher [Ano98t, Bis98, Czi98, Spu98, Var98]. Teachers [HK92, Sne95]. Teaching [Alb93, Alb97c, Ano91a, Ano91b, Ano92a, Ano93a, Ano95a, Ano95b, Ano95c, Ano96a, Ano96b, Ano97m, Ano97n, Ano97o, Ano97p, Ano98i, Ano98p, Ber97c, Bir95, Bis91, Bra96a, CLM92, CRV95, Dal90a, Dal90c, DSJV90, FKL^{+}96a, FKL^{+}96b, GPA97, Hil96b, Kle95, LW98, Loh95, Loo97, Mad98, ME95, Moo93, NS99, Oks91, SM97, SW90, Spi90, Sto98, VM96]. Team [Car96, HS94]. Teams [Gro90]. Tears [SG92]. Technical [Edl91, Kah91]. Technique [Bor97, KAB^{+}97]. Techniques [DGC97, Kha91, Nas92, SM97, Str90]. Ten [Bir95]. Terms [Dav95, Dav98a, Nel98a]. Test [DG97, Dor96, MM90, MC91, ML92a, ML92b, MS92, Sil97, St.90]. Testing [All97, Ano93h, DB97, KG90, Lin93, Loo97, SK93]. Tests [Ber96a, Bir90, Boo92, DBD90, Ede90b, Gre91, GR96, Inm94, LW90, Loo95, LD98, Mag90, MH93, PFI94, SG91, Sho95, SRW95, Suk92, WB91, WW97, ZCM95]. Texas [NO98]. Their [Bor99, Dal90a, Dal90c, Epp93, SBL97]. Theme [DNG98]. Theorem [Ano92g, BHO98, Jes94, Jol95, OS95, Sha91a, DK95]. Theory [Dun99, Gel98, Git94, NS99, Oks91, Sim94, SB98]. Thinking [Sne90]. Thompson [OS95]. Three [Dun99, JL97, Per93, Pro91, Puk94]. Three-Level [Pro91]. Throughout [NOMLT95]. Throw [Ben93]. TI [NO98]. TI-83 [NO98]. Tiered [Bor99]. Tighter [PN95]. Time [CHS99, Czi99, Dan94, Hom97, KT94, PT91, SC99, Unw91]. Time-Series [CHS99]. Timed [Gol92a]. Timed-Events [Gol92a]. Times [Cry91, SB98]. TIMESLAB [Cry91]. Timing [LKNP95]. Tolerance [BK96]. Tool [Kha91, Hol99]. Tools [Dun99, Nas94b, WM96, WOR98]. Total [Git94, Sne90]. Tournament [Car96, SS99]. Tournaments [Edw96, SMH91, SSH96]. Trace [HN98]. Track [NK96]. Traffic [Hu92]. Training [DAF^{+}94]. Transformations [MPB97, SG91, SW91]. Transforming [Atk94, AT99]. Trauma [BCJ97]. Treatments [Faw90, GCGG95, Han92a]. Tree [Atk94]. Trials [Bli90, VBB94, WKB^{+}95]. Triangles [TO93]. Trigonometric [Jon99]. Trivariate [Hut93]. True [Pol99]. Truncated [BS99]. Truncation [Fer90]. Truth [Sch93d]. Twice [LC99]. TWiST [GCGG95]. Two [Ano94f, DGC97, DS98, Ede90b, Gre91, Gro90, Han92a, Hay96, Hos92, LW90, LC99, LL97, Mat99, MV94, MS92, OW91, PS94, RS93, SS98, SF93, Suk92, WS93, Knü94, WS94]. Two-by-Two [Gre91]. Two-Group [DS98, SS98]. Two-Level [LL97]. Two-Period [PS94]. Two-Sample [Mat99, MS92, Suk92, Knü94]. Two-Sided [Ede90b]. Type [Ahm96, Kaz95, Pol99, Sch96c, BS95]. Type-II [BS95]. Typologies [VW93].

U.S. [Gib90, Ley96]. Ultimate [Won98]. Unbalanced [Sea94]. Unbiased [Ang92, BR97, DB97, PR94, SC94, Sea97, SM91]. Unbiasedness [Lin93, MS95]. Unconditional [MH93]. Uncorrelated [Beh90]. Uncovering [Old95]. Undergraduate [GH91, Hig99c]. Understanding [Can98, CG95b, CHS99, Cle99c, Low98, Sul96, WM96, Wue98]. Undesirable [HR98]. Unexpected [PP98]. Unified [Lor90, MSS91]. Uniformly [Bir90, Ede90b]. Unifying [OS95]. Unimodal [Ede90a, SS97]. Unit [AW91b]. United [Gon94]. Units [NOMLT95]. Univariate [Loo95, TM95]. Universal [Sam91]. Universities [CF91, LW98]. University [Kir91, CRV95, KM94]. Unknown [Ede90a, Per93]. Unordered [EGS96]. Unquestionably [GKR97b]. Unrelated [BF91]. untitled [Agr96, AB94, Ame93, And96, Bar97, Bis95, Bol91, Boo97, Boy94, Car93, Che94b, Che94a, Cho97, Cla93, Cle99a, Cle99b, CW94, Coo91, Cry97, D’A96, Del97, Del95, Die97, Dob93a, Dob93b, Dob98, Don93, Don96, Dou97, DS91, DN92, Elt96, Fra97, Fre94, Fri96, GKM^{+}91, Gar97, Gas97, Gor96, Gun97, Gup94, Hal94a, Hal94b, Hal95a, Hal95b, Hal95c, Han95, Har97, Har95, Hay94, Hei91, Her95, Hoe97, Hof91, Hol95, Joh96, Jos93, Kaf97, KK99a, Kaf99a, Kas94, Kin95, Koe97a, Koe97b, Koo90, Koo97, Len98, Les95, Lev92, Lye96, LLL91, MW95, Mal94, Mal97, Man91, Mcc97, Mid96, Mon97a, Moo96, Mor91, Nab97, Nag90, Nei96, Nei93, Nic94a, Nic95, Par98a, Pec97, PRSE94, Pic97, PD93, Pos94, Pos96, RS95, Ree94, Rei94, RR92]. untitled [RR93, Ros93, Ros94, Ros95b, Ros96b, Sap95, Sax94, Sch96b, Sch93c, Sch93e, Sch93f, Sha94, She93, Sho96a, Sho96b, Sim96, Smi97, Sol95, Son90, Tan92, Tan95, The90, Tom93, War95a, War93, Was93, Wat96, Wea97, Win95, Wit93, Wit94, Wit97b, Woo93]. Unweighted [KG95]. Updating [EM93]. Upon [GAR92]. Upper [Pol99, Wri90]. Urban [GKT98]. Use [Bry99b, CR96, Cob99b, Gro99b, HL98, Hay96, Hig99a, Hog99a, Hog99b, KG90, Loo95, Mor98, New99, SM97, Sim94, Ste90]. Used [LC99, SS99]. Useful [DGC97, DB97]. Usefulness [Luc97]. User [Bor97]. Uses [Mad98]. Using [Alb93, All97, AR91, Ano93h, AT99, Beh90, Bon98, BSY92, CLM92, CG95a, Cle99c, CB95b, DBD90, Gab91, Gan95, GCGG95, GKT98, Kas98, Kha92, Koo93, Kva96, MMCsS96, MS98, Pau92, SK93, Smi98, SB98, TO93, Wil94b]. Usual [WL99]. Utilizes [SI90].

V [Tan99]. Value [Don99, Rus91, SB98]. Values [Ber96a, Bly93, Boe94, Fie94, Gro94, Gun94, Hod94, HSF^{+}94, Imr94b, Imr94c, Kad94, Kaf94, Loh94, PT91, SSC99, Sch96a, Sti94, Ver94]. Vanzetti [She98]. VARCL [KDV94]. Variable [GR96, OW91]. Variables [AT99, Beh90, Bri96, Cai94, CR96, Coh91, Lic90, Nav97a, SSC99, vS96]. Variance [Ang92, Ano93h, AS98, BS99, CM92, FN92, Kha92, Lar92, Lin93, MM90, MS92, Nay94, Per93, Sch91a, Sea94, SI90, SK93, Sti95, VM90]. Variances [BS98]. Variates [Bai92, Dav93, EP91, Gan95, PPS96, Ren92]. Variation [AH95, KS91, Sto98, Van96a, vM93]. Vary [HS94]. Venn [LL97]. Versa [Smi95, SRW95]. Versatile [LT97]. Version [Hal92, S.N91, Tan99]. Versions [Loc93]. Versus [Czi99, PF95, GKT98]. Via [Don99, Ano90j, Car96, DCWP93, Gho96, Pal99]. Vice [Smi95, SRW95]. Victory [SS99]. Video [Gab91, Moo93]. View [Wil94a]. Views [AR95]. Violin [HN98]. Visibility [GT95]. Visual [Fri95, Gun98]. Visualizing [Dun99, Gol92a, LKNP95]. Volume [Ano90t, Ano91s, Ano92n, Ano93q, Ano94l, Ano95t, Ano96u, Ano97k, Ano97r, Ano98u, Ano99l, Ano99o, Ano98n].

W. [Boa94]. Waiting [Hom97]. Wald [FBG96, Mag90]. Wanna [CW93]. Way [BK96, RvE97]. Weak [Nel98a]. Weak-Heredity [Nel98a]. Web [WOR98]. Weighted [JK90, KG95, ML92a, ML92b]. Weights [Jor94]. Weldon [KK91]. Welfare [ES93]. Well [Ano91i, LC99, Tur97, Pei90]. Well-Formulated [Ano91i, Pei90]. Well-Known [LC99, Tur97]. We’ve [VAD90]. Whale [RS93]. Where [DWBQ98]. Which [BF91]. Whitney [Ahm96]. Who [Ano97f, Sim98, Tib97, Boa94]. Whose [AB92]. Wide [WOR98]. Wider [Wil94a]. Width [Wan97, Wan99]. Wilcoxon [Ahm96]. William [Ano91t, Ano91u]. Williamsburg [Per98]. Win [Dav98b]. Windows [Mac93, Mon97b, Hil96a, Lur95]. Winning [Ste91, Kel94]. Wishart [Gho96]. Within [Cot98, O’C90]. Within-Subjects [Cot98]. Without [Bar99, He97, SG92]. Wolf [Rie90]. Women [S^{+}90]. Words [GT95]. Work [CW93, LRG90, Pos90]. Workbook [Pet99]. Workshop [Bar98, Cle99c, Law96]. World [Ano97f, Ben93, Tib97, Vid94, Boa94, WOR98]. Would [Kon95]. Would-Be [Kon95]. Writing [RS91a, SO94, SR96].

Yates [NL95]. Year [But90]. Years [Bra90].

Zero [Wri90]. Zeroth [Mal98].

References

Arnold:1992:DWC

[AB92] Barry C. Arnold and Patrick L. Brockett. On distributions whose component ratios are Cauchy. The American Statistician, 46(1): 25–26, February 1992. CODEN ASTAAJ. ISSN 0003-1305 (print),

           1537-2731 (electronic). URL http://www.jstor.org/stable/
           2684405.

Albert:1994:BRB

[AB94] Jim Albert and Mark Berliner. Book review: [untitled]. The American Statistician, 48(2):129–131, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

           jstor.org/stable/2684265.

Agresti:1998:ABT

[AC98a] Alan Agresti and Brent A. Coull. Approximate is better than “Exact” for interval estimation of binomial proportions. The American Statistician, 52(2):119–126, May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

           amstat.org/publications/tas/abstracts_98/agresti.html;
            http://www.jstor.org/stable/2685469.

Anderson-Cook:1998:DFE

[AC98b] C. M. Anderson-Cook. Designing a first experiment: a project for design of experiment courses. The American Statistician, 52(4):338–342, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.amstat.org/publications/tas/abstracts_98/anderson-
cook.html;   http://www.jstor.org/stable/2685437.

Appleton:1996:ICE

[AFV96] David R. Appleton, Joyce M. French, and Mark P. J. Vanderpump. Ignoring a covariate: An example of Simpson’s paradox. The American Statistician, 50(4):340–341, November 1996. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684931.

Arnold:1995:MSR

[AG95] Barry C. Arnold and Richard A. Groeneveld. Measuring skewness with respect to the mode. The American Statistician, 49(1):34–38, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684808.

Agresti:1996:BRB

[Agr96] Alan Agresti. Book review: [untitled]. The American Statistician, 50(1):95–96, February 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685052.

Arnholt:1995:EMK

[AH95] Alan T. Arnholt and Jaimie L. Hebert. Estimating the mean with known coefficient of variation. The American Statistician, 49(4):367–369, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684576.

Ahmad:1996:CMW

[Ahm96] Ibrahim A. Ahmad. A class of Mann–Whitney–Wilcoxon type statistics. The American Statistician, 50(4):324–327, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684929.

Albert:1992:BAP

[Alb92] James Albert. A Bayesian analysis of a Poisson random effects model for home run hitters. The American Statistician, 46(4): 246–253, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685306.

Albert:1993:TBS

[Alb93] James H. Albert. Teaching Bayesian statistics using sampling methods and MINITAB. The American Statistician, 47(3):182– 191, August 1993. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2684973.

Albert:1997:BBS

[Alb97a] Jim Albert. [Bayes for beginners? some reasons to hesitate]: Reply. The American Statistician, 51(3):271–272, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684904.

Albert:1997:RBB

[Alb97b] Jim Albert. Reply — Bayes for beginners? some reasons to hesitate. The American Statistician, 51(3):271–??, August 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Albert:1997:TBR

[Alb97c] Jim Albert. Teaching Bayes’ rule: a data-oriented approach. The American Statistician, 51(3):247–253, August 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts/
albert.html;   http://www.jstor.org/stable/2684896.

Alexander:1991:MIM

[Ale91] William P. Alexander. MATRIX: An interpreted matrix language. The American Statistician, 45(2):154–155, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684384.

Allen:1997:HTU

[All97] David L. Allen. Hypothesis testing using an L1-distance bootstrap. The American Statistician, 51(2):145–150, May 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts/
allen.html;   http://www.jstor.org/stable/2685407.

Altman:1992:IKN

[Alt92] N. S. Altman. An introduction to kernel and nearest-neighbor nonparametric regression. The American Statistician, 46(3):175–185, August 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685209.

Amaratunga:1999:SRS

[Ama99] Dhammika Amaratunga. Searching for the right sample size. The American Statistician, 53(1):52–55, February 1999. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685652.

Ames:1993:BRB

[Ame93] Michael H. Ames. Book review: [untitled]. The American Statistician, 47(1):76–77, February 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684791.

Anderson-Sprecher:1994:MCR

[And94] Richard Anderson-Sprecher. Model comparisons and R^{2}. The American Statistician, 48(2):113–117, May 1994. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684259.

Anderson:1996:BRB

[And96] Robert J. Anderson. Book review: [untitled]. The American Statistician, 50(2):193, May 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684439.

Andrews:1999:BRB

[And99] Douglas M. Andrews. Book review: Statistical Reasoning and Methods. The American Statistician, 53(3):291, August 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2686113.

Angers:1989:NQS

[Ang89] Claude Angers. Note on quick simultaneous confidence intervals for multinomial proportions. The American Statistician, 43(2):91, ???? 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See correction [Ang90].

Angers:1990:CNQ

[Ang90] Claude Angers. Correction: “Note on quick simultaneous confidence intervals for multinomial proportions” [Amer. Statist. 43 (1989), no. 2, 91]. The American Statistician, 44(1):65, ???? 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [Ang89].

Angus:1992:MVU

[Ang92] John E. Angus. Minimum variance unbiased estimator of the probability of an odd number of successes. The American Statistician, 46(2):115–117, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684177.

Ankenman:1998:BRI

[Ank98] Bruce E. Ankenman. Book review: Introduction to the Design and Analysis of Experiments. The American Statistician, 52(4):368, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685445.

Anonymous:1990:ASCa

[Ano90a] Anonymous. ASA’s sesquicentennial celebration. The American Statistician, 44(2):113–??, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1990:ASCb

[Ano90b] Anonymous. ASA’s sesquicentennial celebration. The American Statistician, 44(3):200–??, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1990:BMa

[Ano90c] Anonymous. Back matter. The American Statistician, 44(1):??, February 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684966.

Anonymous:1990:BMb

[Ano90d] Anonymous. Back matter. The American Statistician, 44(2):??, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684167.

Anonymous:1990:BMc

[Ano90e] Anonymous. Back matter. The American Statistician, 44(3):??, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685356.

Anonymous:1990:BMd

[Ano90f] Anonymous. Back matter. The American Statistician, 44(4):??, November 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684363.

Anonymous:1990:CCS

[Ano90g] Anonymous. Correction: Computer selection of size-biased samples. The American Statistician, 44(2):194, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684166.

Anonymous:1990:CEP

[Ano90h] Anonymous. Corrections: An elementary proof of the orderstatistics characterization of the Poisson process. The American Statistician, 44(1):65, February 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684964.

Anonymous:1990:CNQ

[Ano90i] Anonymous. Corrections: Note on quick simultaneous confidence intervals for multinomial proportions. The American Statistician, 44(1):65, February 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684965.

Anonymous:1990:CPS

[Ano90j] Anonymous. Corrections: Power and sample size via MS/PC-DOS computers. The American Statistician, 44(3):264, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685354.

Anonymous:1990:CAS

[Ano90k] Anonymous. Corrections: The ASA section on statistical education: a historical note. The American Statistician, 44(3):264, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685355.

Anonymous:1990:EC

[Ano90l] Anonymous. Editorial collaborators. The American Statistician, 44(4):331–332, November 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684362.

Anonymous:1990:FMa

[Ano90m] Anonymous. Front matter. The American Statistician, 44(1):??, February 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684944.

Anonymous:1990:FMb

[Ano90n] Anonymous. Front matter. The American Statistician, 44(2):??, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684128.

Anonymous:1990:FMc

[Ano90o] Anonymous. Front matter. The American Statistician, 44(3):??, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685332.

Anonymous:1990:FMd

[Ano90p] Anonymous. Front matter. The American Statistician, 44(4):??, November 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684342.

Anonymous:1990:KPM

[Ano90q] Anonymous. N-KERNEL: a practical multivariate density estimation and nonlinear regression program. The American Statistician, 44(3):244, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685350.

Anonymous:1990:NDS

[Ano90r] Anonymous. New developments in statistical computing. The American Statistician, 44(3):242–??, August 1990. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1990:SCS

[Ano90s] Anonymous. Statistical computing software reviews. The American Statistician, 44(3):236–??, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1990:VI

[Ano90t] Anonymous. Volume information. The American Statistician, 44(4):333–335, November 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684341.

Anonymous:1991:ATMa

[Ano91a] Anonymous. Accent on teaching materials. The American Statistician, 45(1):64–??, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1991:ATMb

[Ano91b] Anonymous. Accent on teaching materials. The American Statistician, 45(2):141–??, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1991:BMa

[Ano91c] Anonymous. Back matter. The American Statistician, 45(1):??, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685250.

Anonymous:1991:BMb

[Ano91d] Anonymous. Back matter. The American Statistician, 45(2):??, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684398.

Anonymous:1991:BMc

[Ano91e] Anonymous. Back matter. The American Statistician, 45(3):??, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684307.

Anonymous:1991:BMd

[Ano91f] Anonymous. Back matter. The American Statistician, 45(4):??, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684477.

Anonymous:1991:CCS

[Ano91g] Anonymous. Correction: Computer selection of size-biased samples. The American Statistician, 45(2):172, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684397.

Anonymous:1991:CBH

[Ano91h] Anonymous. Corrections: “A Brief History of the American Statistical Association, 1839–1989”. The American Statistician, 45(1): 82, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685249.

Anonymous:1991:CPW

[Ano91i] Anonymous. Corrections: a property of well-formulated polynomial regression models. The American Statistician, 45(1):82, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685248.

Anonymous:1991:EC

[Ano91j] Anonymous. Editorial collaborators. The American Statistician, 45(4):349–350, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684476.

Anonymous:1991:FMa

[Ano91k] Anonymous. Front matter. The American Statistician, 45(1):??, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685228.

Anonymous:1991:FMb

[Ano91l] Anonymous. Front matter. The American Statistician, 45(2):??, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684364.

Anonymous:1991:FMc

[Ano91m] Anonymous. Front matter. The American Statistician, 45(3):??, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684283.

Anonymous:1991:FMd

[Ano91n] Anonymous. Front matter. The American Statistician, 45(4):??, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684448.

Anonymous:1991:NDSa

[Ano91o] Anonymous. New developments in statistical computing. The American Statistician, 45(1):74–??, February 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1991:NDSb

[Ano91p] Anonymous. New developments in statistical computing. The American Statistician, 45(2):154–??, May 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1991:SCSa

[Ano91q] Anonymous. Statistical computing software reviews. The American Statistician, 45(1):68–??, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1991:SCSb

[Ano91r] Anonymous. Statistical computing software reviews. The American Statistician, 45(2):146–??, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1991:VI

[Ano91s] Anonymous. Volume information. The American Statistician, 45(4):351–353, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684447.

Anonymous:1991:WRSa

[Ano91t] Anonymous. William R. Schucany. The American Statistician, 45(1):1–??, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1991:WRSb

[Ano91u] Anonymous. William R. Schucany. The American Statistician, 45 (4):263–??, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1992:ATM

[Ano92a] Anonymous. Accent on teaching materials. The American Statistician, 46(2):129–??, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1992:BMa

[Ano92b] Anonymous. Back matter. The American Statistician, 46(1):??, February 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684417.

Anonymous:1992:BMb

[Ano92c] Anonymous. Back matter. The American Statistician, 46(2):166, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684191.

Anonymous:1992:BMc

[Ano92d] Anonymous. Back matter. The American Statistician, 46(3):??, August 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685227.

Anonymous:1992:BMd

[Ano92e] Anonymous. Back matter. The American Statistician, 46(4):??, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685331.

Anonymous:1992:CG

[Ano92f] Anonymous. Correction: Genstat 5. The American Statistician, 46(1):77, February 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684416.

Anonymous:1992:CGM

[Ano92g] Anonymous. Correction: The Gauss–Markov theorem and random regressors. The American Statistician, 46(3):243, August 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685226.

Anonymous:1992:EC

[Ano92h] Anonymous. Editorial collaborators. The American Statistician, 46(4):338–339, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685330.

Anonymous:1992:EN

[Ano92i] Anonymous. Editor’s notes. The American Statistician, 46(4): 319–320, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685326.

Anonymous:1992:FMa

[Ano92j] Anonymous. Front matter. The American Statistician, 46(1):??, February 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684399.

Anonymous:1992:FMb

[Ano92k] Anonymous. Front matter. The American Statistician, 46(2):??, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684168.

Anonymous:1992:FMc

[Ano92l] Anonymous. Front matter. The American Statistician, 46(3):??, August 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685207.

Anonymous:1992:FMd

[Ano92m] Anonymous. Front matter. The American Statistician, 46(4):??, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685304.

Anonymous:1992:VI

[Ano92n] Anonymous. Volume information. The American Statistician, 46(4):340–341, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685303.

Anonymous:1993:ATM

[Ano93a] Anonymous. Accent on teaching materials. The American Statistician, 47(4):277–??, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1993:BMa

[Ano93b] Anonymous. Back matter. The American Statistician, 47(1):89, February 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684796.

Anonymous:1993:BMb

[Ano93c] Anonymous. Back matter. The American Statistician, 47(2):164, May 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685206.

Anonymous:1993:BMc

[Ano93d] Anonymous. Back matter. The American Statistician, 47(3):??, August 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684991.

Anonymous:1993:BMd

[Ano93e] Anonymous. Back matter. The American Statistician, 47(4):321, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685302.

Anonymous:1993:BRa

[Ano93f] Anonymous. Books received. The American Statistician, 47(1):78, February 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684793.

Anonymous:1993:BRb

[Ano93g] Anonymous. Books received. The American Statistician, 47(3): 236, August 1993. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2684988.

Anonymous:1993:CTM

[Ano93h] Anonymous. Correction: Testing means using hypothesisdependent variance estimates. The American Statistician, 47(4): 316, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685300.

Anonymous:1993:CWA

[Ano93i] Anonymous. Correction: What about the other intervals? The American Statistician, 47(3):238, August 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684990.

Anonymous:1993:EC

[Ano93j] Anonymous. Editorial collaborators. The American Statistician, 47(4):317–318, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685301.

Anonymous:1993:FMa

[Ano93k] Anonymous. Front matter. The American Statistician, 47(1):??, February 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684773.

Anonymous:1993:FMb

[Ano93l] Anonymous. Front matter. The American Statistician, 47(2):??, May 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685183.

Anonymous:1993:FMc

[Ano93m] Anonymous. Front matter. The American Statistician, 47(3):??, August 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684967.

Anonymous:1993:FMd

[Ano93n] Anonymous. Front matter. The American Statistician, 47(4):??, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685278.

Anonymous:1993:NDS

[Ano93o] Anonymous. New developments in statistical computing. The American Statistician, 47(2):146–147, May 1993. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685199.

Anonymous:1993:SCS

[Ano93p] Anonymous. Statistical computing software reviews. The American Statistician, 47(4):297–??, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1993:VI

[Ano93q] Anonymous. Volume information. The American Statistician, 47(4):319–320, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685277.

Anonymous:1994:BMa

[Ano94a] Anonymous. Back matter. The American Statistician, 48(1):??, February 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685095.

Anonymous:1994:BMb

[Ano94b] Anonymous. Back matter. The American Statistician, 48(2):??, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684282.

Anonymous:1994:BMc

[Ano94c] Anonymous. Back matter. The American Statistician, 48(3):??, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684736.

Anonymous:1994:BMd

[Ano94d] Anonymous. Back matter. The American Statistician, 48(4):??, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684861.

Anonymous:1994:BR

[Ano94e] Anonymous. Books received. The American Statistician, 48(1):59, February 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685093.

Anonymous:1994:CCT

[Ano94f] Anonymous. Correction: Comparing two proportions from the same survey. The American Statistician, 48(3):269, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684735.

Anonymous:1994:EC

[Ano94g] Anonymous. Editorial collaborators. The American Statistician, 48(4):355–357, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684860.

Anonymous:1994:FMa

[Ano94h] Anonymous. Front matter. The American Statistician, 48(1):??, February 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685075.

Anonymous:1994:FMb

[Ano94i] Anonymous. Front matter. The American Statistician, 48(2):??, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684239.

Anonymous:1994:FMc

[Ano94j] Anonymous. Front matter. The American Statistician, 48(3):??, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684712.

Anonymous:1994:FMd

[Ano94k] Anonymous. Front matter. The American Statistician, 48(4):??, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684830.

Anonymous:1994:VI

[Ano94l] Anonymous. Volume information. The American Statistician, 48(4):358–360, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684829.

Anonymous:1995:ATMa

[Ano95a] Anonymous. Accent on teaching materials. The American Statistician, 49(1):62–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1995:ATMb

[Ano95b] Anonymous. Accent on teaching materials. The American Statistician, 49(2):219–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1995:ATMc

[Ano95c] Anonymous. Accent on teaching materials. The American Statistician, 49(4):372–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1995:BMa

[Ano95d] Anonymous. Back matter. The American Statistician, 49(1):??, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684828.

Anonymous:1995:BMb

[Ano95e] Anonymous. Back matter. The American Statistician, 49(2):??, May 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684653.

Anonymous:1995:BMc

[Ano95f] Anonymous. Back matter. The American Statistician, 49(3):??, August 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684216.

Anonymous:1995:BMd

[Ano95g] Anonymous. Back matter. The American Statistician, 49(4):??, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684592.

Anonymous:1995:CSS

[Ano95h] Anonymous. Correction: Sample size determination for interval estimation of multinomial probabilities. The American Statistician, 49(4):403, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684590.

Anonymous:1995:EC

[Ano95i] Anonymous. Editorial collaborators. The American Statistician, 49(4):404–405, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684591.

Anonymous:1995:ENa

[Ano95j] Anonymous. Editor’s notes. The American Statistician, 49(1): 84–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1995:ENb

[Ano95k] Anonymous. Editor’s notes. The American Statistician, 49(2): 235–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1995:ENc

[Ano95l] Anonymous. Editor’s notes. The American Statistician, 49(4): 395–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1995:FMa

[Ano95m] Anonymous. Front matter. The American Statistician, 49(1):??, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684797.

Anonymous:1995:FMb

[Ano95n] Anonymous. Front matter. The American Statistician, 49(2):??, May 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684624.

Anonymous:1995:FMc

[Ano95o] Anonymous. Front matter. The American Statistician, 49(3):??, August 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684192.

Anonymous:1995:FMd

[Ano95p] Anonymous. Front matter. The American Statistician, 49(4):??, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684567.

Anonymous:1995:SCSa

[Ano95q] Anonymous. Statistical computing software reviews. The American Statistician, 49(1):84–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1995:SCSb

[Ano95r] Anonymous. Statistical computing software reviews. The American Statistician, 49(2):235–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1995:SCSc

[Ano95s] Anonymous. Statistical computing software reviews. The American Statistician, 49(4):395–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1995:VI

[Ano95t] Anonymous. Volume information. The American Statistician, 49(4):406–408, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684566.

Anonymous:1996:ATMa

[Ano96a] Anonymous. Accent on teaching materials. The American Statistician, 50(2):168–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1996:ATMb

[Ano96b] Anonymous. Accent on teaching materials. The American Statistician, 50(3):257–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1996:BMa

[Ano96c] Anonymous. Back matter. The American Statistician, 50(1):??, February 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685056.

Anonymous:1996:BMb

[Ano96d] Anonymous. Back matter. The American Statistician, 50(2):??, May 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684446.

Anonymous:1996:BMc

[Ano96e] Anonymous. Back matter. The American Statistician, 50(3):??, August 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684680.

Anonymous:1996:BMd

[Ano96f] Anonymous. Back matter. The American Statistician, 50(4):??, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684943.

Anonymous:1996:CEP

[Ano96g] Anonymous. Correction: Equating Poisson and normal probability functions to derive Stirling’s formula. The American Statistician, 50(2):201, May 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684445.

Anonymous:1996:EC

[Ano96h] Anonymous. Editorial collaborators. The American Statistician, 50(4):390–392, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684942.

Anonymous:1996:ENa

[Ano96i] Anonymous. Editor’s notes. The American Statistician, 50(1): 86–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1996:ENb

[Ano96j] Anonymous. Editor’s notes. The American Statistician, 50(2): 182–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1996:ENc

[Ano96k] Anonymous. Editor’s notes. The American Statistician, 50(3): 266–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1996:ENd

[Ano96l] Anonymous. Editor’s notes. The American Statistician, 50(4):366– ??, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1996:FMa

[Ano96m] Anonymous. Front matter. The American Statistician, 50(1):??, February 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685034.

Anonymous:1996:FMb

[Ano96n] Anonymous. Front matter. The American Statistician, 50(2):??, May 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684418.

Anonymous:1996:FMc

[Ano96o] Anonymous. Front matter. The American Statistician, 50(3):??, August 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684654.

Anonymous:1996:FMd

[Ano96p] Anonymous. Front matter. The American Statistician, 50(4):??, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684920.

Anonymous:1996:SCSa

[Ano96q] Anonymous. Statistical computing software reviews. The American Statistician, 50(1):86–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1996:SCSb

[Ano96r] Anonymous. Statistical computing software reviews. The American Statistician, 50(2):182–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1996:SCSc

[Ano96s] Anonymous. Statistical computing software reviews. The American Statistician, 50(3):266–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1996:SCSd

[Ano96t] Anonymous. Statistical computing software reviews. The American Statistician, 50(4):366–??, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1996:VI

[Ano96u] Anonymous. Volume information. The American Statistician, 50(4):393–395, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684919.

Anonymous:1997:EC

[Ano97a] Anonymous. 1997 editorial collaborators. The American Statistician, 51(4):376–377, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685911.

Anonymous:1997:BMa

[Ano97b] Anonymous. Back matter. The American Statistician, 51(1):??, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684711.

Anonymous:1997:BMb

[Ano97c] Anonymous. Back matter. The American Statistician, 51(2):??, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685426.

Anonymous:1997:BMc

[Ano97d] Anonymous. Back matter. The American Statistician, 51(3):??, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684918.

Anonymous:1997:BMd

[Ano97e] Anonymous. Back matter. The American Statistician, 51(4):??, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685912.

Anonymous:1997:CWF

[Ano97f] Anonymous. Correction: Who is the fastest man in the world? The American Statistician, 51(3):300, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684917.

Anonymous:1997:FMa

[Ano97g] Anonymous. Front matter. The American Statistician, 51(1):1, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684681.

Anonymous:1997:FMb

[Ano97h] Anonymous. Front matter. The American Statistician, 51(2):??, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685395.

Anonymous:1997:FMc

[Ano97i] Anonymous. Front matter. The American Statistician, 51(3):??, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684889.

Anonymous:1997:FMd

[Ano97j] Anonymous. Front matter. The American Statistician, 51(4):??, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685894.

Anonymous:1997:IV

[Ano97k] Anonymous. Index to volume 51. The American Statistician, 51 (4):378–??, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1997:LE

[Ano97l] Anonymous. Letters to the editor. The American Statistician, 51 (2):210–??, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1997:RBTa

[Ano97m] Anonymous. Reviews of books and teaching materials. The American Statistician, 51(1):93–??, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1997:RBTb

[Ano97n] Anonymous. Reviews of books and teaching materials. The American Statistician, 51(2):206–??, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1997:RBTc

[Ano97o] Anonymous. Reviews of books and teaching materials. The American Statistician, 51(3):293–??, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1997:RBTd

[Ano97p] Anonymous. Reviews of books and teaching materials. The American Statistician, 51(4):374–??, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1997:SCS

[Ano97q] Anonymous. Statistical computing software reviews. The American Statistician, 51(4):??, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1997:VI

[Ano97r] Anonymous. Volume information. The American Statistician, 51(4):378–380, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685893.

Anonymous:1998:EC

[Ano98a] Anonymous. 1998 editorial collaborators. The American Statistician, 52(4):372–373, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:/

/www.amstat.org/publications/tas/1998%20referees.html;
 http://www.jstor.org/stable/2685449.

Anonymous:1998:OSL

[Ano98b] Anonymous. [A one-semester, laboratory-based, quality-oriented statistics curriculum for engineering students]: Reply. The American Statistician, 52(3):243, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685935.

Anonymous:1998:ASE

[Ano98c] Anonymous. The American Statistician Editor’s report for 1997. The American Statistician, 52(3):288–??, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1998:BMa

[Ano98d] Anonymous. Back matter. The American Statistician, 52(1):??, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685582.

Anonymous:1998:BMb

[Ano98e] Anonymous. Back matter. The American Statistician, 52(2):??, May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685487.

Anonymous:1998:BMc

[Ano98f] Anonymous. Back matter. The American Statistician, 52(3):??, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685951.

Anonymous:1998:BMd

[Ano98g] Anonymous. Back matter. The American Statistician, 52(4):??, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685450.

Anonymous:1998:C

[Ano98h] Anonymous. Correction. The American Statistician, 52(4):369– ??, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1998:CRB

[Ano98i] Anonymous. Correction: Reviews of books and teaching materials. The American Statistician, 52(4):369, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685447.

Anonymous:1998:FMa

[Ano98j] Anonymous. Front matter. The American Statistician, 52(1):??, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685556.

Anonymous:1998:FMb

[Ano98k] Anonymous. Front matter. The American Statistician, 52(2):??, May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685464.

Anonymous:1998:FMc

[Ano98l] Anonymous. Front matter. The American Statistician, 52(3):??, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685924.

Anonymous:1998:FMd

[Ano98m] Anonymous. Front matter. The American Statistician, 52(4):??, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685428.

Anonymous:1998:IV

[Ano98n] Anonymous. Index to Volume 52. The American Statistician, 52 (4):374–??, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1998:LE

[Ano98o] Anonymous. Letters to the editor. The American Statistician, 52 (4):370–??, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1998:RBT

[Ano98p] Anonymous. Reviews of books and teaching materials. The American Statistician, 52(1):80–??, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1998:SEN

[Ano98q] Anonymous. Section Editor’s notes. The American Statistician, 52(3):266–??, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1998:SCSa

[Ano98r] Anonymous. Statistical computing software reviews. The American Statistician, 52(1):68–??, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1998:SCSb

[Ano98s] Anonymous. Statistical computing software reviews. The American Statistician, 52(1):70–??, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1998:TCR

[Ano98t] Anonymous. Teacher’s corner — reply — A one-semester, laboratory-based, quality-oriented statistics curriculum for engineering students. The American Statistician, 52(3):243–??, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1998:VI

[Ano98u] Anonymous. Volume information. The American Statistician, 52(4):374–376, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685427.

Anonymous:1999:EC

[Ano99a] Anonymous. 1999 editorial collaborators. The American Statistician, 53(4):394–395, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2686068.

Anonymous:1999:BMa

[Ano99b] Anonymous. Back matter. The American Statistician, 53(1):??, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685662.

Anonymous:1999:BMb

[Ano99c] Anonymous. Back matter. The American Statistician, 53(2):??, May 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685745.

Anonymous:1999:BMc

[Ano99d] Anonymous. Back matter. The American Statistician, 53(3):??, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2686120.

Anonymous:1999:BMd

[Ano99e] Anonymous. Back matter. The American Statistician, 53(4):??, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2686069.

Anonymous:1999:CLE

[Ano99f] Anonymous. Correction: Letters to the editor. The American Statistician, 53(4):393, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2686067.

Anonymous:1999:ER

[Ano99g] Anonymous. Editor’s report for 1998. The American Statistician, 53(2):175–??, May 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1999:FMa

[Ano99h] Anonymous. Front matter. The American Statistician, 53(1):??, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685640.

Anonymous:1999:FMb

[Ano99i] Anonymous. Front matter. The American Statistician, 53(2):??, May 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685722.

Anonymous:1999:FMc

[Ano99j] Anonymous. Front matter. The American Statistician, 53(3):??, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2686092.

Anonymous:1999:FMd

[Ano99k] Anonymous. Front matter. The American Statistician, 53(4):??, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2686045.

Anonymous:1999:IV

[Ano99l] Anonymous. Index of volume 53. The American Statistician, 53 (4):396–??, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1999:LE

[Ano99m] Anonymous. Letters to the editor. The American Statistician, 53 (4):393–??, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1999:R

[Ano99n] Anonymous. Reply. The American Statistician, 53(3):201–??, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1999:VI

[Ano99o] Anonymous. Volume information. The American Statistician, 53(4):396–397, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2686044.

Altenburg:1991:AQQ

[AR91] H.-P. Altenburg and G. Rosenkranz. Analysis of quantal and quantitative dose-response data using SAS. The American Statistician, 45(2):159, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684392.

Alper:1995:IVI

[AR95] Paul Alper and Robert L. Raymond. Imposing views, imposing shoes: a statistician as a sole model. The American Statistician, 49(3):317–319, August 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684207.

Abel:1994:SFQ

[AS94] Patricia S. Abel and Nozer D. Singpurwalla. To survive or to fail: That is the question. The American Statistician, 48(1):18–21, February 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685079.

Arnold:1998:JCS

[AS98] Barry C. Arnold and Robert M. Shavelle. Joint confidence sets for the mean and variance of a normal distribution. The American Statistician, 52(2):133–140, May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.amstat.org/publications/tas/abstracts_98/arnold.
html;   http://www.jstor.org/stable/2685471.

Anderson:1999:CSS

[AS99] Jon E. Anderson and Engin A. Sungur. Community service statistics projects. The American Statistician, 53(2):132–136, May 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.amstat.org/publications/tas/
anderson.pdf;   http://www.jstor.org/stable/2685732.

Assuncao:1996:LEE

[Ass96] Renato Assuncao. Likelihood elasticity and error bounds. The American Statistician, 50(2):165–167, May 1996. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684431.

Au:1999:TVU

[AT99] Chi Au and Judy Tam. Transforming variables using the Dirac generalized function. The American Statistician, 53(3):270–272, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2686109.

Aguirre-Torres:1994:LE

[ATHM^{+}94] Victor Aguirre-Torres, Gerald J. Hahn, William Q. Meeker, G. Trenkler, J. B. Douglas, Jorgen Hilden, Katherine L. Monti, and Jeff A. Sloan. Letters to the editor. The American Statistician, 48(1):60–63, February 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685094.

Atkinson:1994:TBS

[Atk94] A. C. Atkinson. Transforming both sides of a tree. The American Statistician, 48(4):307–313, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684841.

Atkinson:1995:IDG

[Atk95] E. Neely Atkinson. Interactive dynamic graphics for exploratory survival analysis. The American Statistician, 49(1):77–84, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684818.

Ames:1991:EAD

[AW91a] Michael H. Ames and John T. Webster. On estimating approximate degrees of freedom. The American Statistician, 45(1):45–50, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685239.

Arndt:1991:EBC

[AW91b] Stephan Arndt and Robert F. Woolson. Establishing a biostatistical core unit in a clinical research center. The American Statistician, 45(1):22–27, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685234.

Baglivo:1995:BRB

[Bag95] Jenny Baglivo. Book review: Computer Algebra Systems: Maple and Mathematica. The American Statistician, 49(1):86–92, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684820.

Bailey:1992:DIB

[Bai92] Ralph W. Bailey. Distributional identities of beta and chi-squared variates: a geometrical interpretation. The American Statistician, 46(2):117–120, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684178.

Bailar:1995:LP

[Bai95] John C. Bailar, III. A larger perspective. The American Statistician, 49(1):10–11, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684802.

Barabba:1990:WDF

[Bar90] Vincent P. Barabba. What does the future hold for statisticians? The American Statistician, 44(2):125–128, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684146.

Barrett:1997:BRB

[Bar97] Linda Barrett. Book review: [untitled]. The American Statistician, 51(3):296, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684912.

Bartelson:1998:BRW

[Bar98] Becki Bucher Bartelson. Book review: Workshop Statistics: Discovery with Data and the Graphing Calculator. The American Statistician, 52(4):367–368, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685444.

Bartelson:1999:BRP

[Bar99] Becki Bucher Bartelson. Book review: Probability Without Equations: Concepts for Clinicians. The American Statistician, 53(2): 171–172, May 1999. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685741.

Bassett:1991:EMB

[Bas91] Gilbert W. Bassett, Jr. Equivariant, monotonic, 50% breakdown estimators. The American Statistician, 45(2):135–137, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684377.

Bassett:1997:RSR

[Bas97] Gilbert W. Bassett, Jr. Robust sports ratings based on least absolute errors. The American Statistician, 51(2):99–105, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts/
bassett.html;   http://www.jstor.org/stable/2685396.

Baxter:1990:FS

[Bax90] Laurence A. Baxter. Futures of statistics. The American Statistician, 44(2):128–129, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684147.

Berger:1992:DGM

[BC92] Roger L. Berger and George Casella. Deriving generalized means as least squares and maximum likelihood estimates. The American Statistician, 46(4):279–282, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685312.

Bedrick:1997:BBR

[BCJ97] Edward J. Bedrick, Ronald Christensen, and Wesley Johnson. Bayesian binomial regression: Predicting survival at a trauma center. The American Statistician, 51(3):211–218, August 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts/
bedrick.html;   http://www.jstor.org/stable/2684890.

Blachman:1996:LE

[BCUF96] Nelson M. Blachman, Ronald Christensen, Jessica M. Utts, and David J. Finney. Letters to the editor. The American Statistician, 50(1):98–99, February 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685055.

Beckman:1996:PFF

[Bec96] Richard J. Beckman. Plotting p^{k} factorial or p^{n−k} fractional factorial data. The American Statistician, 50(2):170–174, May 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684433.

Behboodian:1990:EUD

[Beh90] Javad Behboodian. Examples of uncorrelated dependent random variables using a bivariate mixture. The American Statistician, 44(3):218, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685339.

Bennett:1993:DSJ

[Ben93] Jay Bennett. Did Shoeless Joe Jackson throw the 1919 World Series? The American Statistician, 47(4):241–250, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685280.

Berger:1996:MPT

[Ber96a] Roger L. Berger. More powerful tests from confidence interval p values. The American Statistician, 50(4):314–318, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684927.

Bergerud:1996:DFR

[Ber96b] Wendy A. Bergerud. Displaying factor relationships in experiments. The American Statistician, 50(3):228–233, August 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684660.

Bernard:1996:BIF

[Ber96c] Jean-Marc Bernard. Bayesian interpretation of frequentist procedures for a Bernoulli process. The American Statistician, 50 (1):7–13, February 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685036.

Berry:1997:BBS

[Ber97a] Donald A. Berry. [Bayes for beginners? some reasons to hesitate]: Reply. The American Statistician, 51(3):270–271, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684903.

Berry:1997:RBB

[Ber97b] Donald A. Berry. Reply — Bayes for beginners? some reasons to hesitate. The American Statistician, 51(3):270–??, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Berry:1997:TEB

[Ber97c] Donald A. Berry. Teaching elementary Bayesian statistics with real applications in science. The American Statistician, 51(3):241–246, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.amstat.org/publications/tas/abstracts/berry.htm;
http://www.jstor.org/stable/2684895.

Blom:1998:MP

[BES98] Gunnar Blom, Jan-Eric Englund, and Dennis Sandell. The Mississippi problem. The American Statistician, 52(1):49–50, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.amstat.org/publications/tas/
abstracts_98/blom.html;   http://www.jstor.org/stable/
2685567.

Bartels:1991:SCS

[BF91] Robert Bartels and Denzil G. Fiebig. A simple characterization of seemingly unrelated regressions models in which OLS is BLUE. The American Statistician, 45(2):137–140, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684378.

Brown:1993:HAH

[BfJTH93] Lawrence D. Brown and J. T. Gene Hwang (formerly Jiunn T. Hwang). How to approximate a histogram by a normal density. The American Statistician, 47(4):251–255, November 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685281.

Bruce:1992:LE

[BGD^{+}92] Peter C. Bruce, Peter Guttorp, Gerard E. Dallal, Lynn Roy LaMotte, Donald G. Watts, William Bell, M. Bhaskara Rao, John P. Wendell, Terry R. Ridgway, Richard G. Seymann, W. Edwards Deming, Paul R. McAllister, Dale A. Kopas, and Paul Switzer. Letters to the editor. The American Statistician, 46(3):239–243, August 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685225.

Barr:1998:SNA

[BGG98] Donald R. Barr, Maj. Andrew G. Glen, and Harvey F. Graf. The “Straightforward” nature of arrival rate estimation? The American Statistician, 52(4):346–350, November 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts_
98/barr.html;   http://www.jstor.org/stable/2685439.

Brown:1991:S

[BGS91] Barry W. Brown, David Gutierrez, and Martin F. Spears. STAT-TAB. The American Statistician, 45(2):157, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684389.

Bross:1994:LE

[BGS^{+}94] Irwin D. Bross, Bert Gunter, Ronald D. Snee, Roger L. Berger, Cyrus R. Mehta, Joan F. Hilton, Lynn Roy LaMotte, Roland D. Fisch, Günther A. Strehlau, Kosmas Ferentinos, Stavros Kourouklis, Robert Juola, and G. R. Dargahi-Noubary. Letters to the editor. The American Statistician, 48(2):174–177, May 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684281.

Bassett:1998:EAH

[BH98] G. W. Bassett and W. J. Hurley. The effects of alternative HOME-AWAY sequences in a best-of-seven playoff series. The American Statistician, 52(1):51–53, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/tas/abstracts_98/bassett.html;
 http://www.jstor.org/stable/2685568.

Boos:1998:ABT

[BHO98] Dennis D. Boos and Jacqueline M. Hughes-Oliver. Applications of Basu’s theorem. The American Statistician, 52(3):218–221, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.amstat.org/publications/tas/
abstracts_98/boos.html;   http://www.jstor.org/stable/
2685927.

Beachell:1994:LE

[BHR^{+}94] Eileen J. Beachell, Ruth Hubbard, Sheldon M. Ross, Ronald Christensen, Jessica Utts, Roger L. Berger, Dennis D. Boos, David A.

Sprott, Vernon T. Farewell, G. S. Watson, Oliver D. Anderson, and Rich Dodenhoff. Letters to the editor. The American Statistician, 48(3):267–269, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684734.

Bickel:1995:WAN

[Bic95] Peter J. Bickel. What academia needs. The American Statistician, 49(1):5–6, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684800.

Billard:1998:RSS

[Bil98] Lynne Billard. The role of statistics and the statistician. The American Statistician, 52(4):319–324, November 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts_
98/billard.html;   http://www.jstor.org/stable/2685434.

Birkes:1990:GLR

[Bir90] David Birkes. Generalized likelihood ratio tests and uniformly most powerful tests. The American Statistician, 44(2):163–166, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684159.

Birch:1995:TSE

[Bir95] Jeffrey B. Birch. Ten suggestions for effectively teaching short courses to heterogeneous groups. The American Statistician, 49(2): 190–195, May 1995. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2684637.

Bisgaard:1991:TSE

[Bis91] Søren Bisgaard. Teaching statistics to engineers. The American Statistician, 45(4):274–283, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684452.

Bisgaard:1995:BRB

[Bis95] Soren Bisgaard. Book review: [untitled]. The American Statistician, 49(3):320–321, August 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684208.

Bisgaard:1998:TCD

[Bis98] Søren Bisgaard. Teacher’s corner — discussion — A one-semester, laboratory-based, quality-oriented statistics curriculum for engineering students. The American Statistician, 52(3):238–239, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685931.

Boen:1990:LE

[BJK^{+}90] James R. Boen, Norman L. Johnson, Samuel Kotz, Leon Jay Gleser, Donald R. Ploch, W. D. Kaigh, Wolfgang J. Buhler, Donald A. Berry, James Berger, David A. Harville, Herbert F. Spirer, Michael D. Conerly, Oscar Kempthorne, Andrew P. Grieve, Lawrence L. Kupper, Kerry B. Hafner, R. W. Farebrother, Ronald Christensen, Simo Puntanen, George P. H. Styan, Fred Ederer, Daniel G. Seigel, Sunil K. Sapra, A. W. Marshall, and I. Olkin. Letters to the editor. The American Statistician, 44(2):184–194, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684165.

Becker:1992:LLS

[BK92] William Becker and Peter Kennedy. A lesson in least squares and R squared. The American Statistician, 46(4):282–283, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685313.

Boen:1993:CIS

[BK93] James R. Boen and Marcus O. Kjelsberg. Career information for statistics graduate students. The American Statistician, 47(2): 148–149, May 1993. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685200.

Bhaumik:1996:SEM

[BK96] D. K. Bhaumik and P. M. Kulkarni. A simple and exact method of constructing tolerance intervals for the one-way ANOVA with random effects. The American Statistician, 50(4):319–323, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684928.

Bolton:1999:CPC

[BK99] Richard J. Bolton and Wojtek J. Krzanowski. A characterization of principal components for projection pursuit. The American Statistician, 53(2):108–109, May 1999. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685727.

Becker:1996:PM

[BKM96] Richard A. Becker and Sallie Keller-McNulty. Presentation myths. The American Statistician, 50(2):112–115, May 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684420.

Bar-Lev:1994:DCC

[BL94] Shaul K. Bar-Lev. A derivation of conditional cumulants in exponential models. The American Statistician, 48(2):126–129, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684263.

Bliss:1990:CCP

[Bli90] J. M. Bliss. COMPACT: a computer package for clinical trials management. The American Statistician, 44(3):242, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685346.

Blyth:1993:REP

[Bly93] Colin R. Blyth. Restrict estimates to the possible values? The American Statistician, 47(1):73–75, February 1993. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684789.

Best:1991:SPD

[BM91] Al M. Best and David Morganstein. Statistics programs designed for the macintosh: Data desk, exstatix, fastat, JMP, StatView II, and SuperANOVA. The American Statistician, 45(4):318–338, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684465.

Boomsma:1994:FET

[BM94] Anne Boomsma and Ivo W. Molenaar. Four electronic tables for probability distributions. The American Statistician, 48(2):153– 162, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684276.

Barton:1998:OSL

[BN98] Russell R. Barton and Craig A. Nowack. A one-semester, laboratory-based, quality-oriented statistics curriculum for engineering students. The American Statistician, 52(3):233– 238, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.amstat.

org/publications/tas/abstracts_98/barton.html;   http:
//www.jstor.org/stable/2685930.

Bancroft:1994:LE

[BNL^{+}94] Diccon Bancroft, Heinz Neudecker, Shuangzhe Liu, N. R. Chaganty, A. K. Vaish, Peter Nemenyi, R. Murray Lindsay, Andrew S. C. Ehrenberg, Nico F. Laubscher, M. C. Jones, N. L. Hjort, Lawrence D. Brown, and J. T. Gene Hwang. Letters to the editor. The American Statistician, 48(4):351–354, November 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684856.

Bowerman:1992:IQC

[BO92] Bruce L. Bowerman and Richard T. O’Connell. Integrating quality control concepts into the introductory business statistics course. The American Statistician, 46(2):124–128, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684181.

Boardman:1994:SWC

[Boa94] Thomas J. Boardman. The statistician who changed the world: W. Edwards Deming, 1900–1993. The American Statistician, 48 (3):179–187, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684713.

Boen:1994:SVQ

[Boe94] James Boen. [Statistical values, quality, and certification]: Invited comment. The American Statistician, 48(2):73, May 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684244.

Bollen:1991:BRB

[Bol91] Kenneth A. Bollen. Book review: [untitled]. The American Statistician, 45(1):68–73, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685244.

Bond:1998:BRA

[Bon98] Marjorie E. Bond. Book reviews: Applied Statistical Methods for Business, Economics, and the Social Sciences and Statistics for Managers: Using Microsoft Excel. The American Statistician, 52(2):201–202, May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685484.

Boos:1992:GST

[Boo92] Dennis D. Boos. On generalized score tests. The American Statistician, 46(4):327–333, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685328.

Booth:1997:BRB

[Boo97] David E. Booth. Book review: [untitled]. The American Statistician, 51(2):209, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685424.

Borack:1997:TEP

[Bor97] Jules I. Borack. A technique for estimating the probability of detecting a nongaming drug user. The American Statistician, 51(2):134–136, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.amstat.org/publications/tas/abstracts/borack.html;
 http://www.jstor.org/stable/2685404.

Boronico:1999:MTP

[Bor99] Jess S. Boronico. Multi-tiered playoffs and their impact on professional baseball. The American Statistician, 53(1):56–61, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685653.

Bost:1990:H

[Bos90] R. H. ‘Butch’ Bost, Jr. 2120 hindsight. The American Statistician, 44(2):130–131, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684148.

Boyer:1994:BRB

[Boy94] John E. Boyer, Jr. Book review: [untitled]. The American Statistician, 48(1):58, February 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685089.

Balakrishnan:1997:NBL

[BR97] N. Balakrishnan and C. R. Rao. A note on the best linear unbiased estimation based on order statistics. The American Statistician, 51(2):181–185, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.amstat.org/

publications/tas/abstracts/balakrishnan.html;   http:/
/www.jstor.org/stable/2685416.

Brailsford:1990:FSN

[Bra90] Thomas W. Brailsford. The future of statistics in the next N years (25 ¡ N ¡ 150). The American Statistician, 44(2):131, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684149.

Bradstreet:1996:TIS

[Bra96a] Thomas E. Bradstreet. Teaching introductory statistics courses so that nonstatisticians experience statistical reasoning. The American Statistician, 50(1):69–78, February 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685047.

Brant:1996:DLR

[Bra96b] Rollin Brant. Digesting logistic regression results. The American Statistician, 50(2):117–119, May 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684422.

Bross:1991:LE

[BRG^{+}91] Irwin D. Bross, Marion R. Reynolds, Jr., Timothy G. Gregoire, Stanley S. Schor, Beth Dawson-Saunders, Paul K. Jones, Steven J. Verhulst, Burt Holland, Per Lea, Dave Saville, Eugene Garfield, I. J. Good, Daniel Gianola, Jean-Louis Foulley, Laurence A. Baxter, Sherman Chottiner, Paul A. Herzberg, Judith D. Singer, John B. Willett, Ronald A. Thisted, Chap T. Le, Jean D. Gibbons, B. M. Pőtscher, C. M. Hurvich, and C. L. Tsai. Letters to the editor. The American Statistician, 45(2):163–172, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684396.

Bring:1994:HSR

[Bri94] Johan Bring. How to standardize regression coefficients. The American Statistician, 48(3):209–213, August 1994. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684719.

Bring:1996:GAC

[Bri96] Johan Bring. A geometric approach to compare variables in a regression model. The American Statistician, 50(1):57–62, February 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685045.

Bromaghin:1993:SSD

[Bro93] Jeffrey F. Bromaghin. Sample size determination for interval estimation of multinomial probabilities. The American Statistician, 47(3):203–206, August 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684978.

Bromaghin:1994:CSS

[Bro94] Jeffrey F. Bromaghin. Correction: Sample size determination for interval estimation of multinomial probabilities. The American Statistician, 48(4):354, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684859.

Brunelle:1991:BS

[Bru91] Rocco L. Brunelle. Breeze/STAT. The American Statistician, 45 (3):246, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2684301.

Bryce:1999:D

[Bry99a] G. Rex Bryce. Discussion. The American Statistician, 53(1):21– ??, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bryce:1999:LUC

[Bry99b] G. Rex Bryce. [Let’s use CQI in our statistics programs]: Discussion. The American Statistician, 53(1):21–23, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685645.

Balakrishnan:1995:SSA

[BS95] N. Balakrishnan and R. A. Sandhu. A simple simulational algorithm for generating progressive Type-II censored samples. The American Statistician, 49(2):229–230, May 1995. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684646.

Booth:1998:MCA

[BS98] James G. Booth and Somnath Sarkar. Monte Carlo approximation of bootstrap variances. The American Statistician, 52 (4):354–357, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.amstat.

org/publications/tas/abstracts_98/booth.html;   http:/
/www.jstor.org/stable/2685441.

Barr:1999:MVT

[BS99] Donald R. Barr and E. Todd Sherrill. Mean and variance of truncated normal distributions. The American Statistician, 53 (4):357–361, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2686057.

Boullion:1992:MDP

[BSY92] Thomas L. Boullion, John W. Seaman, Jr., and Dean M. Young. Moments of discrete probability distributions derived using a differential operator. The American Statistician, 46(1):22–24, February 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684404.

Blom:1990:DRP

[BTV90] Gunnar Blom, Daniel Thorburn, and Theodore A. Vessey. The distribution of the record position and its applications. The American Statistician, 44(2):151–153, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684157.

Butz:1990:EYC

[But90] William P. Butz. An extrapolation to the year 2100 career possibilities for statisticians. The American Statistician, 44(2):132, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684150.

Buzas:1997:FEJ

[Buz97] J. S. Buzas. Fast estimators of the jackknife. The American Statistician, 51(3):235–240, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:/

/www.amstat.org/publications/tas/abstracts/buzas.html;
 http://www.jstor.org/stable/2684894.

Berger:1998:QSA

[BWM98] Paul D. Berger, Chen Wang, and James P. Monahan. Quantifying a statistical aspect of segmented selection/quota systems. The American Statistician, 52(3):228–232, August 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts_
98/berger.html;   http://www.jstor.org/stable/2685929.

Chen:1997:ID

[CA97] Gemai Chen and A. Adatia. Independence and t distribution. The American Statistician, 51(2):176–177, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.amstat.org/publications/tas/abstracts/chen.
html;   http://www.jstor.org/stable/2685414.

Cain:1994:MGF

[Cai94] Michael Cain. The moment-generating function of the minimum of bivariate normal random variables. The American Statistician, 48(2):124–125, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684262.

Cannon:1998:BRU

[Can98] Ann R. Cannon. Book review: Understanding Data: Principles and Practice of Statistics. The American Statistician, 52(4):367, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685443.

Carlson:1993:BRB

[Car93] Richard R. Carlson. Book review: [untitled]. The American Statistician, 47(4):308–309, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685295.

Carlin:1996:INB

[Car96] Bradley P. Carlin. Improved NCAA basketball tournament modeling via point spread and team strength information. The American Statistician, 50(1):39–43, February 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685042.

Coe:1995:OSS

[CB95a] Paul R. Coe and William Butterworth. Optimal stopping in “The showcase showdown”. The American Statistician, 49(3):271–275, August 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684199.

Cook:1995:BSU

[CB95b] Peyton Cook and Lyle D. Broemeling. Bayesian statistics using Mathematica. The American Statistician, 49(1):70–76, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684817.

Cockerill:1991:IPA

[CF91] Rhonda Cockerill and Bruce Fried. Increasing public awareness of statistics as a science and a profession — reinforcing the message in universities. The American Statistician, 45(3):174–178, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684285.

Casella:1992:EGS

[CG92] George Casella and Edward I. George. Explaining the Gibbs sampler. The American Statistician, 46(3):167–174, August 1992. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685208.

Charnes:1995:UCC

[CG95a] John M. Charnes and Howard S. Gitlow. Using control charts to corroborate bribery in jai alai. The American Statistician, 49(4):386–389, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684581.

Chib:1995:UMH

[CG95b] Siddhartha Chib and Edward Greenberg. Understanding the Metropolis–Hastings algorithm. The American Statistician, 49 (4):327–335, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684568.

Choi:1991:LPL

[CGO91] Sungsub Choi, K. Ruben Gabriel, and Charles L. Odoroff. LO-RANK: a package for lower rank approximation of matrices. The American Statistician, 45(3):246–247, August 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684302.

Christensen:1996:NAR

[CH96] Ronald Christensen and Aparna V. Huzurbazar. A note on augmenting resolution III designs. The American Statistician, 50(2): 175–177, May 1996. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2684434.

Chattamvelli:1995:NNB

[Cha95] R. Chattamvelli. A note on the noncentral beta distribution function. The American Statistician, 49(2):231–234, May 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684647.

Chambers:1999:CDC

[Cha99] John Chambers. Computing with data: Concepts and challenges. The American Statistician, 53(1):73–84, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685657.

Campbell:1995:LE

[CHD^{+}95] J. M. Campbell, James J. Higgins, Gerard E. Dallal, Chamont Wang, Arthur Silverberg, Peter B. Imrey, Oliver D. Anderson, Bruno D. Zumbo, and Jorgen Hilden. Letters to the editor. The American Statistician, 49(1):115–119, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684827.

Chernick:1994:BRBb

[Che94a] Michael Chernick. Book review: [untitled]. The American Statistician, 48(4):347–348, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684848.

Chernick:1994:BRBa

[Che94b] Michael R. Chernick. Book review: [untitled]. The American Statistician, 48(1):58–59, February 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685090.

Chou:1997:BRB

[Cho97] Youn-Min Chou. Book review: [untitled]. The American Statistician, 51(1):95, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684705.

Choudhury:1999:UTS

[CHS99] Askar H. Choudhury, Robert Hubata, and Robert D. St. Louis. Understanding time-series regression estimators. The American Statistician, 53(4):342–348, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2686054.

Chun:1999:IEA

[Chu99] Young H. Chun. On the information economics approach to the generalized game show problem. The American Statistician, 53 (1):43–51, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685651.

Christensen:1995:LE

[CIM^{+}95] Ronald Christensen, Henry F. Inman, Xiao-Li Meng, Ananda Sen, Joseph B. Keller, Kevin Donegan, and Peter Nemenyi. Letters to the editor. The American Statistician, 49(4):400–403, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684589.

Cohen:1992:ASE

[CJ92] Steven B. Cohen and Ayah E. Johnson. Address sample effects in the National Medical Expenditure Survey. The American Statistician, 46(2):100–105, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684173.

Clason:1993:BRB

[Cla93] Dennis Clason. Book review: [untitled]. The American Statistician, 47(4):310, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685298.

Cleary:1999:BRBa

[Cle99a] Richard Cleary. Book review: [untitled]. The American Statistician, 53(3):292, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2686115.

Cleary:1999:BRBb

[Cle99b] Richard Cleary. Book review: [untitled]. The American Statistician, 53(4):388–389, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2686062.

Cleary:1999:BRW

[Cle99c] Richard Cleary. Book review: Workshop Statistics: Discovery with Data and Minitab, Data Analysis with Microsoft Excel, and Statistical Laboratory Exercises Using Excel: A Guide to Understanding Data. The American Statistician, 53(4):388–??, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cleary:1999:SME

[Cle99d] Richard Cleary. Statistical methods for engineers. The American Statistician, 53(3):292–??, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chang:1992:TSS

[CLM92] Ted C. Chang, Sharon L. Lohr, and C. Graham McLaren. Teaching survey sampling using simulation. The American Statistician, 46 (3):232–237, August 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685223.

Clason:1992:EFV

[CM92] Dennis L. Clason and Leigh W. Murray. Exponential families and variance component models. The American Statistician, 46 (1):29–31, February 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684407.

Cobb:1999:D

[Cob99a] George W. Cobb. Discussion. The American Statistician, 53(1): 16–??, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cobb:1999:LUC

[Cob99b] George W. Cobb. [Let’s use CQI in our statistics programs]: Discussion. The American Statistician, 53(1):16–21, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685644.

Cohen:1991:DVS

[Coh91] Ayala Cohen. Dummy variables in stepwise regression. The American Statistician, 45(3):226–228, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684296.

Cohen:1997:EAP

[Coh97] Steven B. Cohen. An evaluation of alternative PC-Based software packages developed for the analysis of complex survey data. The American Statistician, 51(3):285–292, August 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts/
cohen.html;   http://www.jstor.org/stable/2684908.

Cook:1991:BRB

[Coo91] Richard L. Cook. Book review: [untitled]. The American Statistician, 45(2):146–149, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684381.

Cotton:1998:BRA

[Cot98] John W. Cotton. Book review: Analyzing Within-Subjects Experiments. The American Statistician, 52(4):368–??, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Clemons:1999:BN

[CP99] Traci Clemons and Marcello Pagano. Are babies normal? The American Statistician, 53(4):298–302, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2686047.

Chambers:1990:ASC

[CR90] John M. Chambers and Barbara F. Ryan. The ASA statistical computing section: a history. The American Statistician, 44(2):87– 89, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684135.

Cameron:1995:CE

[CR95] Joseph M. Cameron and Joan R. Rosenblatt. Churchill Eisenhart, 1913–1994. The American Statistician, 49(3):243–244, August 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684193.

Carroll:1996:UMO

[CR96] R. J. Carroll and David Ruppert. The use and misuse of orthogonal regression in linear errors-in-variables models. The American Statistician, 50(1):1–6, February 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685035.

Cornell:1995:CTS

[CRV95] John A. Cornell, Ron H. Randles, and G. Geoff Vining. Challenges in teaching short courses by an academic department: The University of Florida experience. The American Statistician, 49(2):195– 200, May 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684638.

Cryer:1991:BRB

[Cry91] Jonathan D. Cryer. Book review: TIMESLAB: a Times Series Laboratory. The American Statistician, 45(3):243–244, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684299.

Cryer:1997:BRB

[Cry97] Jonathan D. Cryer. Book review: [untitled]. The American Statistician, 51(1):95–96, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684706.

Chevan:1991:HP

[CS91] Albert Chevan and Michael Sutherland. Hierarchical partitioning. The American Statistician, 45(2):90–96, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684366.

Chen:1994:QAF

[CS94] Wenjun Chen and Tim Swartz. Quantitative aspects of five-pin bowling. The American Statistician, 48(2):92–98, May 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684254.

Chib:1991:EBA

[CT91] Siddhartha Chib and Ram C. Tiwari. Extreme bounds analysis in the Kalman filter. The American Statistician, 45(2):113–114, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684370.

Christensen:1992:BRE

[CU92] Ronald Christensen and Jessica Utts. Bayesian resolution of the “Exchange paradox”. The American Statistician, 46(4):274–276, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685310.

Cuadras:1993:IIM

[Cua93] Carles M. Cuadras. Interpreting an inequality in multiple regression. The American Statistician, 47(4):256–258, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685282.

Cox:1992:CCD

[CW92] D. R. Cox and Nanny Wermuth. A comment on the coefficient of determination for binary responses. The American Statistician, 46(1):1–4, February 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684400.

Coyle:1993:WBG

[CW93] Cynthia A. Coyle and Chamont Wang. Wanna bet? on gambling strategies that may or may not work in a casino. The American Statistician, 47(2):108–111, May 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685188.

Conerly:1994:BRB

[CW94] Michael D. Conerly and William H. Woodall. Book review: [untitled]. The American Statistician, 48(4):347, November 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684847.

Cook:1999:GSA

[CW99] R. Dennis Cook and Sanford Weisberg. Graphs in statistical analysis: Is the medium the message? The American Statistician, 53 (1):29–37, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685649.

Chatterjee:1999:NEM

[CY99] Sangit Chatterjee and Mustafa R. Yilmaz. The NBA as an evolving multivariate system. The American Statistician, 53(3):257–262, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2686106.

Czitrom:1998:TCD

[Czi98] Veronica Czitrom. Teacher’s corner — discussion — A onesemester, laboratory-based, quality-oriented statistics curriculum for engineering students. The American Statistician, 52(3):240, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685932.

Czitrom:1999:OFT

[Czi99] Veronica Czitrom. One-factor-at-a-time versus designed experiments. The American Statistician, 53(2):126–131, May 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.amstat.org/publications/tas/
czitrom.pdf;   http://www.jstor.org/stable/2685731.

DAgostino:1996:BRB

[D’A96] Ralph B. D’Agostino. Book review: [untitled]. The American Statistician, 50(2):193–194, May 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684440.

DeMets:1994:TNG

[DAF^{+}94] David L. DeMets, Dan Anbar, William Fairweather, Thomas A. Louis, and Robert T. O’Neill. Training the next generation of biostatisticians. The American Statistician, 48(4):280–284, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684833.

Dallal:1990:EIC

[Dal90a] Gerard E. Dallal. Editor’s invited column: Statistical computing packages: Dare we abandon their teaching to others? The American Statistician, 44(4):265–266, November 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684343.

Dallal:1990:PSC

[Dal90b] Gerard E. Dallal. PC-SIZE: Consultant — a program for samplesize determinations. The American Statistician, 44(3):243, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685347.

Dallal:1990:SCP

[Dal90c] Gerard E. Dallal. Statistical computing packages: Dare we abandon their teaching to others? The American Statistician, 44 (4):265–??, November 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dallal:1991:LLM

[Dal91] Gerard E. Dallal. LMS: Least median of squares regression. The American Statistician, 45(1):74, February 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685245.

Danaher:1992:SSM

[Dan92] Peter J. Danaher. Some statistical modeling problems in the advertising industry: a look at media exposure distributions. The American Statistician, 46(4):254–260, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685307.

Daniel:1994:FOF

[Dan94] Cuthbert Daniel. Factorial one-factor-at-a-time experiments. The American Statistician, 48(2):132–135, May 1994. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684266.

David:1993:NOS

[Dav93] H. A. David. A note on order statistics for dependent variates. The American Statistician, 47(3):198–199, August 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684976.

David:1995:FOC

[Dav95] H. A. David. First (?) occurrence of common terms in mathematical statistics. The American Statistician, 49(2):121–133, May 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684625.

David:1998:FOC

[Dav98a] H. A. David. First (?) occurrence of common terms in probability and statistics — A second list, with corrections. The American Statistician, 52(1):36–40, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/tas/abstracts_98/david.html;
http://www.jstor.org/stable/2685564.

David:1998:SOM

[Dav98b] H. A. David. Stronger opposition may make it easier to win. The American Statistician, 52(4):351–353, November 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts_
98/david2.html;   http://www.jstor.org/stable/2685440.

Dawkins:1991:SPC

[Daw91] Brian Dawkins. Siobhan’s problem: The coupon collector revisited. The American Statistician, 45(1):76–82, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685247.

Dawson:1996:HML

[Daw96] Robert J. MacG. Dawson. How many light bulbs does it take to generate a data set? The American Statistician, 50(3):247–249, August 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684666.

Dayton:1998:ICP

[Day98] C. Mitchell Dayton. Information criteria for the pairedcomparisons problem. The American Statistician, 52(2):144–151, May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.amstat.org/publications/tas/
abstracts_98/dayton.html;   http://www.jstor.org/stable/
2685473.

Doganaksoy:1997:UPB

[DB97] Necip Doganaksoy and N. Balakrishnan. A useful property of best linear unbiased predictors with applications to life-testing. The American Statistician, 51(1):22–28, February 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts/
doganaksoy.html;   http://www.jstor.org/stable/2684687.

DAgostino:1990:SUP

[DBD90] Ralph B. D’Agostino, Albert Belanger, and Ralph B. D’Agostino, Jr. A suggestion for using powerful and informative tests of normality. The American Statistician, 44(4):316–321, November 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684359.

Deng:1990:RRE

[DC90] Lih-Yuan Deng and Raj S. Chhikara. On the ratio and regression estimation in finite population sampling. The American Statistician, 44(4):282–284, November 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684347.

Drane:1993:LFP

[DCWP93] J. Wanzer Drane, Su Hua Cao, Lixia Wang, and T. Postelnicu. Limiting forms of probability mass functions via recurrence formulas. The American Statistician, 47(4):269–274, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685285.

DeLong:1994:IRD

[DeL94] David DeLong. Invited response to Dallal, G. E. (1992), “The Computer Analysis of Factorial Experiments with Nested Factors,” The American Statistician, 46, 240. The American Statistician, 48(2):141, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684270.

Delaney:1995:BRB

[Del95] Nancy Jo Delaney. Book review: [untitled]. The American Statistician, 49(4):397, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684584.

DelCastillo:1997:BRB

[Del97] Enrique Del Castillo. Book review: [untitled]. The American Statistician, 51(4):374, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685908.

Derr:1993:BCI

[Der93] Janice Derr. Biostatistics cores: Improving the chances for funding. The American Statistician, 47(2):99–101, May 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685185.

David:1997:PTU

[DG97] H. A. David and Jason L. Gunnink. The paired t test under artificial pairing. The American Statistician, 51(1):9–12, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts/
david.html;   http://www.jstor.org/stable/2684684.

Dawson:1997:TGT

[DGC97] Kathryn S. Dawson, Chris Gennings, and Walter H. Carter. Two graphical techniques useful in detecting correlation structure in repeated measures data. The American Statistician, 51(3):275–283, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.amstat.org/publications/tas/abstracts/dawson.html;
 http://www.jstor.org/stable/2684906.

Dietz:1993:CLA

[Die93] E. Jacquelin Dietz. A cooperative learning activity on methods of selecting a sample. The American Statistician, 47(2):104–108, May 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685187.

Dietz:1997:BRB

[Die97] E. Jacquelin Dietz. Book review: [untitled]. The American Statistician, 51(2):207–208, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685422.

Driscoll:1995:APC

[DK95] Michael F. Driscoll and Barbara Krasnicka. An accessible proof of Craig’s theorem in the general case. The American Statistician, 49(1):59–62, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684814.

Dueker:1992:BRB

[DN92] Marilynn S. Dueker and Michael Nabel. Book review: [untitled]. The American Statistician, 46(3):226–227, August 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685221.

Dinh:1994:APM

[DN94a] K. T. Dinh and Truc T. Nguyen. Acknowledgement of priority: “Maximum likelihood estimators of binomial parameters under an order restriction” [Amer. Statist. 48 (1994), no. 1, 29–30]. The American Statistician, 48(4):354, ???? 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [DN94c].

Dinh:1994:CML

[DN94b] K. T. Dinh and Truc T. Nguyen. Correction: Maximum likelihood estimators of binomial parameters under an order restriction. The American Statistician, 48(4):354, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684858.

Dinh:1994:MLE

[DN94c] Khoan T. Dinh and Truc T. Nguyen. Maximum likelihood estimators of binomial parameters under an order restriction. The American Statistician, 48(1):29–30, February 1994. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685081. See acknowledgement of pri-

ority [DN94a].

Dargahi-Noubary:1998:RMT

[DNG98] G. R. Dargahi-Noubary and Jo Anne S. Growney. Risk — A motivating theme for an introductory statistics course. The American Statistician, 52(1):44–48, February 1998. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/tas/abstracts_98/noubary.html;
 http://www.jstor.org/stable/2685566.

Dobler:1993:BRBa

[Dob93a] Carolyn Pillers Dobler. Book review: [untitled]. The American Statistician, 47(4):307–308, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685294.

Dobler:1993:BRBb

[Dob93b] Carolyn Pillers Dobler. Book review: [untitled]. The American Statistician, 47(4):309–310, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685297.

Dobler:1998:BRB

[Dob98] Carolyn Pillers Dobler. Book review: [untitled]. The American Statistician, 52(3):284–285, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685948.

Donohue:1993:BRB

[Don93] Joan M. Donohue. Book review: [untitled]. The American Statistician, 47(4):309, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685296.

Donohue:1996:BRB

[Don96] Joan M. Donohue. Book review: [untitled]. The American Statistician, 50(4):379, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684938.

Donahue:1999:NIS

[Don99] Rafe M. J. Donahue. A note on information seldom reported via the P value. The American Statistician, 53(4):303–306, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2686048.

Dorfman:1996:PSR

[Dor96] Jeffrey H. Dorfman. Pseudorandom sampling has a real effect on test size. The American Statistician, 50(2):151–155, May 1996.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684427.

Douglas:1993:CRP

[Dou93] J. B. Douglas. Confidence regions for parameter pairs. The American Statistician, 47(1):43–45, February 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684784.

Dougherty:1997:BRB

[Dou97] Anne M. Dougherty. Book review: [untitled]. The American Statistician, 51(1):94–95, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684704.

Dupont:1991:PCP

[DP91] William D. Dupont and Walton D. Plummer, Jr. POWER.FOR: a computer program for power and sample size calculations. The American Statistician, 45(2):158, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684390.

Dodge:1999:CFC

[DR99] Yadolah Dodge and Valentin Rousson. The complications of the fourth central moment. The American Statistician, 53(3):267–269, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2686108.

Driscoll:1999:IRR

[Dri99] Michael F. Driscoll. An improved result relating quadratic forms and chi-square distributions. The American Statistician, 53(3): 273–275, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2686110.

Dueker:1991:BRB

[DS91] Marilynn S. Dueker and Herbert F. Spirer. Book review: [untitled]. The American Statistician, 45(1):64–65, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685242.

DuarteSilva:1998:NTG

[DS98] A. Pedro Duarte Silva and Antonie Stam. Nonparametric twogroup classification: Concepts and a SAS-based software package. The American Statistician, 52(2):183–?? (or 185–??), May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.amstat.org/publications/tas/
abstracts_98/silva.html.

Dawson-Saunders:1990:HST

[DSJV90] Beth Dawson-Saunders, Paul K. Jones, and Steven J. Verhulst. The history of the subsection on teaching of statistics in the health sciences. The American Statistician, 44(2):101–103, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684140.

Dallal:1992:LE

[DSS^{+}92] Gerard E. Dallal, John Stufken, Michael Sutherland, Martin Fox, I. S. Alalouf, Jonathan J. Shuster, Stephen M. Stigler, Malay Ghosh, Peter W. Zehna, Ronald Christensen, Richard K. Burdick, Michael H. Ames, Dudley M. Harde, Bruce Levin, S. R. Searle, R. K. Elswick, Jr., Y. H. Wang, and Victor Perez-Abreu. Letters to the editor. The American Statistician, 46(1):70–77, February 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684415.

DuMouchel:1999:BDMa

[DuM99a] William DuMouchel. Bayesian data mining in large frequency tables, with an application to the FDA spontaneous reporting system. The American Statistician, 53(3):177–190, August 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2686093.

DuMouchel:1999:BDMb

[DuM99b] William DuMouchel. [Bayesian data mining in large frequency tables, with an application to the FDA spontaneous reporting system]: Reply. The American Statistician, 53(3):201–202, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2686097.

Dunn:1991:PSP

[Dun91] Charles L. Dunn. Precise simulated percentiles in a pinch. The American Statistician, 45(3):207–211, August 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684291.

Dunlop:1994:RLD

[Dun94] Dorothy D. Dunlop. Regression for longitudinal data: a bridge from least squares regression. The American Statistician, 48(4): 299–303, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684838.

Dunn:1999:TTI

[Dun99] Peter K. Dunn. Three tools for interactively visualizing some distribution theory concepts. The American Statistician, 53(2):137– 139, May 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.amstat.org/publications/
tas/dunn.pdf;   http://www.jstor.org/stable/2685733.

DeMets:1998:WJS

[DWBQ98] David L. DeMets, Robert Woolson, Curtis Brooks, and Roger Qu. Where the jobs are: a study of Amstat News job advertisements. The American Statistician, 52(4):303–307, November 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685431.

Edelman:1990:CIC

[Ede90a] David Edelman. A confidence interval for the center of an unknown unimodal distribution based on a sample of size 1. The American Statistician, 44(4):285–287, November 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684348.

Edelman:1990:NUM

[Ede90b] David Edelman. A note on uniformly most powerful two-sided tests. The American Statistician, 44(3):219–220, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685340.

Edlefsen:1991:ATG

[Edl91] Lee E. Edlefsen. Axum technical graphics and data analysis. The American Statistician, 45(2):157, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684388.

Edwards:1996:DET

[Edw96] Christopher T. Edwards. Double-elimination tournaments: Counting and calculating. The American Statistician, 50(1):27–33, February 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685040.

Elswick:1991:SAF

[EGCD91] R. K. Elswick, Jr., Chris Gennings, Vernon M. Chinchilli, and Kathryn S. Dawson. A simple approach for finding estimable functions in linear models. The American Statistician, 45(1):51–53, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685240.

Ernst:1996:SUP

[EGS96] Michael D. Ernst, Rudy Guerra, and William R. Schucany. Scatterplots for unordered pairs. The American Statistician, 50(3):260– 265, August 1996. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2684670.

Ehrenberg:1990:EIC

[Ehr90] A. S. C. Ehrenberg. Editor’s invited column: a hope for the future of statistics: MSOD. The American Statistician, 44(3):195–196, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685333.

Eldridge:1990:SAC

[Eld90] Marie D. Eldridge. The status of advisory committees to the Federal Statistical Agencies. The American Statistician, 44(2):154– 162, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684158.

Eltinge:1996:BRB

[Elt96] Elizabeth M. Eltinge. Book review: [untitled]. The American Statistician, 50(1):96–97, February 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685053.

Escobar:1993:NUR

[EM93] Luis A. Escobar and E. Barry Moser. A note on the updating of regression estimates. The American Statistician, 47(3):192–194, August 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684974.

Emerson:1996:BRB

[Eme96] Scott S. Emerson. Book review: Statistical Packages for Group Sequential Methods. The American Statistician, 50(2):183–192, May 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684438.

Engeman:1993:SCP

[Eng93] Richard M. Engeman. Short courses as part of statistical consulting in a foreign setting. The American Statistician, 47(2):122–125, May 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685193.

Emrich:1991:MGH

[EP91] Lawrence J. Emrich and Marion R. Piedmonte. A method for generating high-dimensional multivariate binary variates. The American Statistician, 45(4):302–304, November 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684460.

Epps:1993:CFT

[Epp93] T. W. Epps. Characteristic functions and their empirical counterparts: Geometrical interpretations and applications to statistical inference. The American Statistician, 47(1):33–38, February 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684780.

Engeman:1993:AWS

[ES93] Richard M. Engeman and Stephen A. Shumake. Animal welfare and the statistical consultant. The American Statistician, 47 (3):229–233, August 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684983.

Eubank:1994:SSS

[Eub94] R. L. Eubank. A simple smoothing spline. The American Statistician, 48(2):103–106, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684256.

Farnum:1990:IRE

[Far90] Nicholas R. Farnum. Improving the relative error of estimation. The American Statistician, 44(4):288–289, November 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684349.

Fawcett:1990:AMD

[Faw90] Richard F. Fawcett. The alignment method for displaying and analyzing treatments in blocking designs. The American Statistician, 44(3):204–209, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685336.

Finn:1991:MP

[FB91] Jeremy D. Finn and R. Darrell Bock. Multivariance PC. The American Statistician, 45(3):247, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684303.

Fears:1996:RFW

[FBG96] Thomas R. Fears, Jacques Benichou, and Mitchell H. Gail. A reminder of the fallibility of the Wald statistic. The American Statistician, 50(3):226–227, August 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684659.

Ferentinos:1990:SCI

[Fer90] Kosmas K. Ferentinos. Shortest confidence intervals for families of distributions involving truncation parameters. The American Statistician, 44(2):167–168, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684160.

Fienberg:1994:SVQ

[Fie94] Stephen E. Fienberg. [Statistical values, quality, and certification]: Invited comment. The American Statistician, 48(2):71–72, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684243.

Fisher:1990:LE

[FJH^{+}90] N. I. Fisher, Ian T. Jolliffe, Chien-Pai Han, Jean D. Gibbons, Julia Lane, Russ Ray, Dennis Glennon, Sunil K. Sapra, Roy T. St. Laurent, James W. Mergerson, Colin R. Blyth, Vedula N. Murty, Jeffrey P. Benedict, and Tommy Wright. Letters to the editor. The American Statistician, 44(4):327–330, November 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684361.

Frey:1997:PCC

[FK97] Michael R. Frey and John C. Kern. The Pitman closeness of a class of scaled estimators. The American Statistician, 51(2):151–154, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.amstat.org/publications/tas/abstracts/frey.html;
http://www.jstor.org/stable/2685408.

Fecso:1996:TSS

[FKL^{+}96a] Ronald S. Fecso, William D. Kalsbeek, Sharon L. Lohr, Richard L. Scheaffer, Fritz J. Scheuren, and Elizabeth A. Stasny. Teaching survey sampling. The American Statistician, 50(4):328–340, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684930.

Fesco:1996:TSS

[FKL^{+}96b] Ronald S. Fesco, William D. Kalsbeek, Sharon L. Lohr, Richard L. Scheaffer, Fritz J. Scheuren, and Elizabeth A. Stasny. Teaching survey sampling. The American Statistician, 50(4):328–??, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fleury:1994:IRD

[Fle94] Patrick J. Fleury. Invited response to Dallal, G. E. (1992), “The Computer Analysis of Factorial Experiments with Nested Factors,” The American Statistician, 46, 240. The American Statistician, 48(2):145–147, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684273.

Federer:1992:CAS

[FM92] Walter T. Federer and Michael P. Meredith. Covariance analysis for split-plot and split-block designs. The American Statistician, 46(2):155–162, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684189.

Flury:1992:MAM

[FN92] Bernard D. Flury and A. Narayanan. A mixture approach to multivariate analysis of variance. The American Statistician, 46 (1):31–34, February 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684408.

Franklin:1992:GIM

[Fra92] LeRoy A. Franklin. Graphical insight into multiple regression concepts. The American Statistician, 46(4):284–288, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685314.

Frankowski:1997:BRB

[Fra97] Ralph F. Frankowski. Book review: [untitled]. The American Statistician, 51(3):297, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684914.

Free:1990:SHB

[Fre90] Spencer M. Free, Jr. Some history for the biopharmaceutical section. The American Statistician, 44(2):80–82, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684132.

Freeny:1994:BRB

[Fre94] Anne Freeny. Book review: [untitled]. The American Statistician, 48(4):349, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684851.

Freedman:1997:DBB

[Fre97] David A. Freedman. Discussion — Bayes for beginners? some reasons to hesitate. The American Statistician, 51(3):266–268, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684901.

Friendly:1995:CVM

[Fri95] Michael Friendly. Conceptual and visual models for categorical data. The American Statistician, 49(2):153–160, May 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684630.

Fricker:1996:BRB

[Fri96] Ronald D. Fricker, Jr. Book review: [untitled]. The American Statistician, 50(3):278–279, August 1996. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684678.

Fisch:1993:SAC

[FS93] Roland D. Fisch and Günther A. Strehlau. A simplified approach to calibration confidence sets. The American Statistician, 47(3): 168–171, August 1993. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2684969.

Flury:1993:RLC

[FT93] Bernard D. Flury and Thaddeus Tarpey. Representing a large collection of curves: a case for principal points. The American Statistician, 47(4):304–306, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685292.

Gabriel:1991:UAA

[Gab91] K. Ruben Gabriel. Using the “Against all odds” video series in an introductory statistics course. The American Statistician, 45(2):142–145, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684380.

Gange:1995:GMC

[Gan95] Stephen J. Gange. Generating multivariate categorical variates using the iterative proportional fitting algorithm. The American Statistician, 49(2):134–138, May 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684626.

Gregoire:1992:MSE

[GAR92] Timothy G. Gregoire, Alexandros A. Arabatzis, and Marion R. Reynolds, Jr. Mean squared error performance of simple linear regression conditional upon the outcome of pretesting the intercept. The American Statistician, 46(2):89–93, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684171.

Garfield:1995:MSP

[Gar95] Joan B. Garfield. [Modernizing statistics Ph.D. programs]: Respondent. The American Statistician, 49(1):18–20, February 1995.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684804.

Garfield:1997:BRB

[Gar97] Joan Garfield. Book review: [untitled]. The American Statistician, 51(4):375, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685910.

Gastwirth:1992:SRL

[Gas92] Joseph L. Gastwirth. Statistical reasoning in the legal setting. The American Statistician, 46(1):55–69, February 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684414.

Gastwirth:1997:BRB

[Gas97] Joseph L. Gastwirth. Book review: [untitled]. The American Statistician, 51(3):297–299, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684916.

Gates:1990:NE

[Gat90] Charles E. Gates. New editor. The American Statistician, 44(1):1, February 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684945.

Gates:1995:WRE

[Gat95] Charles E. Gates. What really is experimental error in block designs? The American Statistician, 49(4):362–363, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684574.

Goutis:1999:ESA

[GC99] Constantino Goutis and George Casella. Explaining the saddlepoint approximation. The American Statistician, 53(3):216–224, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2686100.

Gelber:1995:CTU

[GCGG95] Richard D. Gelber, Bernard F. Cole, Shari Gelber, and Aron Goldhirsch. Comparing treatments using quality-adjusted survival: The Q-TWiST method. The American Statistician, 49(2):161–169, May 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684631.

Gelman:1998:SCP

[Gel98] Andrew Gelman. Some class-participation demonstrations for decision theory and Bayesian statistics. The American Statistician, 52(2):167–174, May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/tas/abstracts_98/gelman2.html;
 http://www.jstor.org/stable/2685476.

Gaudard:1991:UCA

[GH91] Marie Gaudard and Gerald J. Hahn. An undergraduate concentration in applied statistics for mathematics majors. The American Statistician, 45(2):115–120, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684371.

Ghosh:1996:WDI

[Gho96] Malay Ghosh. Wishart distribution via induction. The American Statistician, 50(3):243–246, August 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684664.

Gibbons:1990:UIR

[Gib90] Jean D. Gibbons. U.S. institutional representation on editorial boards of U.S. statistics journals. The American Statistician, 44 (3):210–213, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685337.

Gitlow:1994:CJT

[Git94] Howard S. Gitlow. A comparison of Japanese total quality control and Deming’s theory of management. The American Statistician, 48(3):197–203, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684716.

Gautam:1999:SRM

[GK99] Shiva Gautam and George Kimeldorf. Some results on the maximal correlation in 2×k contingency tables. The American Statistician, 53(4):336–341, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2686053.

Gabriel:1991:BRB

[GKM^{+}91] K. Ruben Gabriel, Thomas R. Knapp, James McNamara, Donna Rose, and Carl Wetzstein. Book review: [untitled]. The American Statistician, 45(2):141–142, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684379.

Gastwirth:1994:HCA

[GKR94] Joseph Gastwirth, Abba Krieger, and Paul Rosenbaum. How a court accepted an impossible explanation. The American Statistician, 48(4):313–315, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684842. See comment [PG97].

Gastwirth:1997:RHH

[GKR97a] Joseph L. Gastwirth, Abba M. Krieger, and Paul R. Rosenbaum. Reply: Hypotheticals and hypotheses. The American Statistician, 51(2):120–121, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685401.

Gastwirth:1997:RUI

[GKR97b] Joseph L. Gastwirth, Abba M. Krieger, and Paul R. Rosenbaum. Reply: Unquestionably impossible. The American Statistician, 51(2):115–116, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685399.

Gan:1991:PPD

[GKT91] F. F. Gan, Kenneth J. Koehler, and John C. Thompson. Probability plots and distribution curves for assessing the fit of probability models. The American Statistician, 45(1):14–21, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685233.

Goodall:1998:CUR

[GKT98] Colin R. Goodall, Karen Kafadar, and John W. Tukey. Computing and using rural versus urban measures in statistical applications. The American Statistician, 52(2):101–111, May 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts_
98/goodall.html;   http://www.jstor.org/stable/2685467.

Gelman:1991:NBD

[GM91] Andrew Gelman and Xiao-Li Meng. A note on bivariate distributions that are conditionally normal. The American Statistician, 45(2):125–126, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684374.

Gunst:1996:IOD

[GM96] Richard F. Gunst and Gary C. McDonald. The importance of outcome dynamics, simple geometry, and pragmatic statistical arguments in exposing deficiencies of experimental design strategies. The American Statistician, 50(1):44–50, February 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685043.

Gnanadesikan:1990:LAC

[Gna90] R. Gnanadesikan. Looking ahead: Cross-disciplinary opportunities for statistics. The American Statistician, 44(2):121–125, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684145.

Gelman:1998:SPS

[GNM^{+}98] Andrew Gelman, Deborah Nolan, Anna Men, Steve Warmerdam, and Michelle Bautista. Student projects on statistical literacy and the media. The American Statistician, 52(2):160–166, May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685475.

Goddard:1991:CSC

[God91] M. J. Goddard. Constructing some categorical anomalies. The American Statistician, 45(2):129–134, May 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684376.

Goldfield:1990:HSS

[Gol90] Edwin D. Goldfield. History of the social statistics section. The American Statistician, 44(2):84–87, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684134.

Goldstein:1991:EN

[Gol91] Richard Goldstein. Editor’s notes. The American Statistician, 45(4):304–305, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684461.

Goldman:1992:EVS

[Gol92a] Anne I. Goldman. EVENTCHARTS: Visualizing survival and other timed-events data. The American Statistician, 46(1):13–18, February 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684402.

Goldstein:1992:ENa

[Gol92b] Richard Goldstein. Editor’s notes. The American Statistician, 46 (1):48–49, February 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684411.

Goldstein:1992:ENb

[Gol92c] Richard Goldstein. Editor’s notes. The American Statistician, 46 (2):154, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2684188.

Goldstein:1992:ENc

[Gol92d] Richard Goldstein. Editor’s notes. The American Statistician, 46 (3):237–238, August 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685224.

Goldstein:1993:ENb

[Gol93a] Richard Goldstein. Editor’s notes. The American Statistician, 47(2):134–136, May 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685197.

Goldstein:1993:ENc

[Gol93b] Richard Goldstein. Editor’s notes. The American Statistician, 47(3):216, August 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684981.

Goldstein:1993:ENd

[Gol93c] Richard Goldstein. Editor’s notes. The American Statistician, 47(4):297–298, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685290.

Goldstein:1994:ENa

[Gol94a] Richard Goldstein. Editor’s notes. The American Statistician, 48 (1):43–44, February 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685086.

Goldstein:1994:ENb

[Gol94b] Richard Goldstein. Editor’s notes. The American Statistician, 48(2):138–139, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684268.

Goldstein:1994:ENc

[Gol94c] Richard Goldstein. Editor’s notes. The American Statistician, 48 (3):254–255, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684731.

Goldstein:1995:ENa

[Gol95a] Richard Goldstein. Editor’s notes. The American Statistician, 49 (1):84–86, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684819.

Goldstein:1995:ENb

[Gol95b] Richard Goldstein. Editor’s notes. The American Statistician, 49(2):235–236, May 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684648.

Goldstein:1995:ENc

[Gol95c] Richard Goldstein. Editor’s notes. The American Statistician, 49 (3):296–297, August 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684204.

Goldstein:1995:ENd

[Gol95d] Richard Goldstein. Editor’s notes. The American Statistician, 49(4):395–396, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684583.

Goldstein:1996:ENa

[Gol96a] Richard Goldstein. Editor’s notes. The American Statistician, 50 (1):86–87, February 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685049.

Goldstein:1996:ENb

[Gol96b] Richard Goldstein. Editor’s notes. The American Statistician, 50(2):182–183, May 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684437.

Goldstein:1996:ENc

[Gol96c] Richard Goldstein. Editor’s notes. The American Statistician, 50 (3):266–267, August 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684671.

Goldstein:1996:ENd

[Gol96d] Richard Goldstein. Editor’s notes. The American Statistician, 50(4):366–367, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684936.

Goldstein:1997:SENa

[Gol97a] Richard Goldstein. Section Editor’s notes. The American Statistician, 51(1):65–66, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684698.

Goldstein:1997:SENb

[Gol97b] Richard Goldstein. Section Editor’s notes. The American Statistician, 51(2):193, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685418.

Goldman:1998:BRI

[Gol98] Robert N. Goldman. Book review: Interactive Statistics: Preliminary Edition. The American Statistician, 52(3):283, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685946.

Goldstein:1999:BRS

[Gol99] Richard Goldstein. Book review: Statistics with Stata 5 . The American Statistician, 53(4):392, November 1999. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2686065.

Gonzalez:1994:IDQ

[Gon94] Maria Elena Gonzalez. Improving data quality awareness in the United States Federal Statistical Agencies. The American Statistician, 48(1):12–17, February 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685078.

Good:1990:ASC

[Goo90] I. J. Good. Abstract of “Speculations concerning the future of statistics”. The American Statistician, 44(2):132–133, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684151.

Gorman:1996:BRB

[Gor96] Bernard S. Gorman. Book review: [untitled]. The American Statistician, 50(3):257–258, August 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684668.

Goel:1997:CAT

[GPA97] Prem K. Goel, Mario Peruggia, and Baoshe An. Computer-aided teaching of probabilistic modeling for biological phenomena. The American Statistician, 51(2):164–169, May 1997. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.amstat.org/publications/tas/abstracts/goel.
html;   http://www.jstor.org/stable/2685411.

Grimmett:1996:EVD

[GR96] D. R. Grimmett and J. R. Ridenhour. The effect of a variable data point on hypothesis tests for means. The American Statistician, 50(2):145–150, May 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684426.

Greenland:1991:LJC

[Gre91] Sander Greenland. On the logical justification of conditional tests for two-by-two contingency tables. The American Statistician, 45 (3):248–251, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684304.

Greenland:1996:LBC

[Gre96] Sander Greenland. A lower bound for the correlation of exponentiated bivariate normal pairs. The American Statistician, 50(2):163– 164, May 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684430.

Groeneveld:1990:RTL

[Gro90] Richard A. Groeneveld. Ranking teams in a league with two divisions of t teams. The American Statistician, 44(4):277–281, November 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684346.

Groeneveld:1991:IFA

[Gro91] Richard A. Groeneveld. An influence function approach to describing the skewness of a distribution. The American Statistician, 45(2):97–102, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684367.

Groshen:1994:SVQ

[Gro94] Susan Groshen. [Statistical values, quality, and certification]: Invited comment. The American Statistician, 48(2):74, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684246.

Groeneveld:1998:CQM

[Gro98] Richard A. Groeneveld. A class of quantile measures for kurtosis. The American Statistician, 52(4):325–329, November 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts_
98/groen.html;   http://www.jstor.org/stable/2685435.

Groeneveld:1999:D

[Gro99a] Richard A. Groeneveld. Discussion. The American Statistician, 53(1):23–??, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Groeneveld:1999:LUC

[Gro99b] Richard A. Groeneveld. [Let’s use CQI in our statistics programs]: Discussion. The American Statistician, 53(1):23–25, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685646.

Gbur:1995:KWP

[GT95] Edward E. Gbur, Jr. and Bruce E. Trumbo. Key words and phrases — the key to scholarly visibility and efficiency in an information explosion. The American Statistician, 49(1):29–33, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684807.

Gunter:1993:TFS

[Gun93] Bert Gunter. Through a funnel slowly with ball bearing and insight to teach experimental design. The American Statistician, 47(4):265–269, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685284.

Gunter:1994:SVQ

[Gun94] Bert Gunter. [Statistical values, quality, and certification]: Invited comment. The American Statistician, 48(2):77–78, May 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684249.

Gunawardena:1997:BRB

[Gun97] K. L. D. Gunawardena. Book review: [untitled]. The American Statistician, 51(1):96, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684707.

Gunter:1998:BRB

[Gun98] Berton Gunter. Book review: Visual Revelations: Graphical Tales of Fate and Deception from Napoleon Bonaparte to Ross Perot. The American Statistician, 52(1):83–84, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685574.

Gupta:1994:BRB

[Gup94] Ramesh C. Gupta. Book review: [untitled]. The American Statistician, 48(4):350, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684854.

Goldstein:1993:ENa

[GV93] Richard Goldstein and Paul F. Velleman. Editor’s notes. The American Statistician, 47(1):46–47, February 1993. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684785.

Gray:1994:MSS

[GW94] J. Brian Gray and William H. Woodall. The maximum size of standardized and internally studentized residuals in regression analysis. The American Statistician, 48(2):111–113, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684258.

Gan:1989:EPO

[GY89] Gaoxiong Gan and Shie-Shien Yang. An elementary proof of the order-statistics characterization of the Poisson process. The American Statistician, 43(1):45–46, ???? 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See correction [GY90].

Gan:1990:CEP

[GY90] Gaoxiong Gan and Shie-Shien Yang. Correction: “An elementary proof of the order-statistics characterization of the Poisson process” [Amer. Statist. 43 (1989), no. 1, 45–46]. The American Statistician, 44(1):65, ???? 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [GY89].

Haaland:1990:HSA

[Haa90] Perry D. Haaland. Historical sketch of the ASA section on statistical graphics. The American Statistician, 44(2):96–98, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684138.

Hahn:1995:DII

[Hah95] Gerald J. Hahn. Deming’s impact on industrial statistics: Some reflections. The American Statistician, 49(4):336–341, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684569.

Hallahan:1992:BRB

[Hal92] Charles Hallahan. Book review: DBMS/COPY and DBMS/COPY Plus — (Version 2.0). The American Statistician, 46(1):49–52, February 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684412.

Halvorsen:1994:BRBa

[Hal94a] Katherine Taylor Halvorsen. Book review: [untitled]. The American Statistician, 48(4):350, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684853.

Halvorsen:1994:BRBb

[Hal94b] Katherine Taylor Halvorsen. Book review: [untitled]. The American Statistician, 48(4):350, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684855.

Halvorsen:1995:BRBa

[Hal95a] Katherine Halvorsen. Book review: [untitled]. The American Statistician, 49(3):321, August 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684209.

Halvorsen:1995:BRBb

[Hal95b] Katherine Halvorsen. Book review: [untitled]. The American Statistician, 49(3):322, August 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684211.

Halvorsen:1995:BRBc

[Hal95c] Katherine Taylor Halvorsen. Book review: [untitled]. The American Statistician, 49(4):398, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684586.

Hamming:1990:FS

[Ham90] Richard W. Hamming. The future of statistics. The American Statistician, 44(2):133–135, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684152.

Hammer:1997:SES

[Ham97] Heather Hammer. Symbolic exclusion in statistical literature: The impact of gendered language. The American Statistician, 51(1):13–19, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.amstat.org/publications/tas/abstracts/hammer.html;
 http://www.jstor.org/stable/2684685.

Hand:1992:CTT

[Han92a] David J. Hand. On comparing two treatments. The American Statistician, 46(3):190–192, August 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685211.

Hanley:1992:JCD

[Han92b] James A. Hanley. Jumping to coincidences: Defying odds in the realm of the preposterous. The American Statistician, 46(3):197– 202, August 1992. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685213.

Handcock:1995:BRB

[Han95] Mark S. Handcock. Book review: [untitled]. The American Statistician, 49(2):238, May 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684650.

Hand:1998:DMS

[Han98] David J. Hand. Data mining: Statistics and more? The American Statistician, 52(2):112–118, May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts_
98/hand.html;   http://www.jstor.org/stable/2685468.

Hartlaub:1995:BRB

[Har95] Bradley A. Hartlaub. Book review: [untitled]. The American Statistician, 49(4):397–398, November 1995. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684585.

Hare:1997:BRB

[Har97] Lynne B. Hare. Book review: [untitled]. The American Statistician, 51(3):295–296, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684911.

Haughton:1997:PEF

[Hau97] Dominique Haughton. Packages for estimating finite mixtures: a review. The American Statistician, 51(2):194–205, May 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685419.

Hawkins:1991:FFI

[Haw91] Douglas M. Hawkins. FIRM: Formal inference-based recursive modeling. The American Statistician, 45(2):155, May 1991. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684385.

Hayden:1994:BRB

[Hay94] Robert W. Hayden. Book review: [untitled]. The American Statistician, 48(4):305–306, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684840.

Hayden:1996:BRB

[Hay96] Robert W. Hayden. Book review: Review of Two Collections of Data for Use in a First Course in Statistics. The American Statistician, 50(2):168–169, May 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684432.

Hayden:1999:BRI

[Hay99] Robert W. Hayden. Book review: Introduction to Statistical Reasoning. The American Statistician, 53(1):86, February 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685659.

Heitjan:1996:DMR

[HB96] Daniel F. Heitjan and Srabashi Basu. Distinguishing “missing at random” and “missing completely at random”. The American Statistician, 50(3):207–213, August 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684656.

Haslett:1991:DGE

[HBC^{+}91] John Haslett, Ronan Bradley, Peter Craig, Antony Unwin, and Graham Wills. Dynamic graphics for exploring spatial data with application to locating global and local anomalies. The American Statistician, 45(3):234–242, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684298.

Hahn:1990:HSQ

[HBF90] Gerald J. Hahn, Thomas J. Boardman, and David Fluharty. History of the section on quality and productivity of the American Statistical Association. The American Statistician, 44(3):197–199, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685334.

Hillis:1994:SJI

[HD94] Stephen L. Hillis and Charles S. Davis. A simple justification of the iterative fitting procedure for generalized linear models. The American Statistician, 48(4):288–289, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684835.

He:1997:QCC

[He97] Xuming He. Quantile curves without crossing. The American Statistician, 51(2):186–192, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/tas/abstracts/he.html;   http:
//www.jstor.org/stable/2685417.

Heiberger:1991:BRB

[Hei91] Richard M. Heiberger. Book review: [untitled]. The American Statistician, 45(4):311–315, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684463.

Helland:1995:SCA

[Hel95] Inge S. Helland. Simple counterexamples against the conditionality principle. The American Statistician, 49(4):351–356, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684572.

Helberg:1999:BRB

[Hel99] Clay Helberg. Book review: Finding Statistics Online: How to Locate the Elusive Numbers You Need. The American Statistician, 53(3):293–294, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2686117.

Hengartner:1999:NML

[Hen99] Nicolas W. Hengartner. A note on maximum likelihood estimation. The American Statistician, 53(2):123–125, May 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685730.

Herson:1990:B

[Her90] Jay Herson. Biostatistics 2089. The American Statistician, 44(2): 136–137, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2684153.

Herrin:1995:BRB

[Her95] Gary D. Herrin. Book review: [untitled]. The American Statistician, 49(2):238–239, May 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684651.

Hess:1990:HSS

[Hes90] Irene Hess. History of the section on survey research methods. The American Statistician, 44(2):98–100, May 1990. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684139.

Hyndman:1996:SQS

[HF96] Rob J. Hyndman and Yanan Fan. Sample quantiles in statistical packages. The American Statistician, 50(4):361–365, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684934.

Harrell:1997:SMS

[HG97] Frank E. Harrell, Jr. and Richard Goldstein. A survey of microcomputer survival analysis software: The need for an integrated framework. The American Statistician, 51(4):360–373, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.amstat.org/publications/tas/
abstracts/harrell.html;   http://www.jstor.org/stable/
2685907.

Hines:1995:ECS

[HH95] R. J. O’Hara Hines and W. G. S. Hines. Exploring Cook’s statistic graphically. The American Statistician, 49(4):389–394, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684582.

Hayes:1999:SGL

[HH99] Kevin Hayes and John Haslett. Simplifying general least squares. The American Statistician, 53(4):376–381, November 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2686060.

Hill:1991:LRA

[HHB91] L. Robert Hill, William G. Hammond, and John R. Benfield. Logistic regression analysis in experimental bronchial carcinogenesis. The American Statistician, 45(3):184–186, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684287.

Hahn:1999:ISS

[HHHZ99] Gerald J. Hahn, William J. Hill, Roger W. Hoerl, and Stephen A. Zinkgraf. The impact of six sigma improvement—A glimpse into the future of statistics. The American Statistician, 53(3):208–215, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2686099.

Hoerl:1993:SIS

[HHJL93] Roger W. Hoerl, Jeffrey H. Hooper, Peter J. Jacobs, and James M. Lucas. Skills for industrial statisticians to survive and prosper in the emerging quality environment. The American Statistician, 47(4):280–292, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685288.

Higgins:1999:LUC

[Hig99a] James J. Higgins. [Let’s use CQI in our statistics programs]: Reply. The American Statistician, 53(1):26–27, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685647.

Higgins:1999:LE

[Hig99b] James J. Higgins. Letters to the editor. The American Statistician, 53(2):170–??, 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Higgins:1999:NSN

[Hig99c] James J. Higgins. Nonmathematical statistics: a new direction for the undergraduate discipline. The American Statistician, 53(1):1– 6, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685641.

Higgins:1999:R

[Hig99d] James J. Higgins. Reply. The American Statistician, 53(1):26– ??, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:1993:BRB

[Hil93] Joseph M. Hilbe. Book review: Generalized Additive Models Software. The American Statistician, 47(1):59–64, February 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684787.

Hilbe:1994:BRB

[Hil94] Joseph M. Hilbe. Book review: Generalized Linear Models. The American Statistician, 48(3):255–265, August 1994. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684732.

Hilbe:1996:BRB

[Hil96a] Joseph Hilbe. Book review: Windows File Conversion Software. The American Statistician, 50(3):268–270, August 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684672.

Hillmer:1996:PSA

[Hil96b] Steven C. Hillmer. A problem-solving approach to teaching business statistics. The American Statistician, 50(3):249–256, August 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684667.

Hilbe:1997:SENa

[Hil97a] Joseph Hilbe. Section Editor’s notes. The American Statistician, 51(3):284, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684907.

Hilbe:1997:SENb

[Hil97b] Joseph Hilbe. Section Editor’s notes. The American Statistician, 51(4):359, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.amstat.org/

publications/tas/abstracts/hilbe.html;   http://www.jstor.
org/stable/2685906.

Hilbe:1998:SENa

[Hil98a] Joseph Hilbe. Section Editor’s notes. The American Statistician, 52(1):70, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685571.

Hilbe:1998:SENb

[Hil98b] Joseph Hilbe. Section Editor’s notes. The American Statistician, 52(3):266, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685941.

Hilbe:1999:SEN

[Hil99] Joseph Hilbe. Section Editor’s notes. The American Statistician, 53(2):147–148, May 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685735.

Hines:1996:PPA

[Hin96] W. G. S. Hines. Pragmatics of pooling in ANOVA tables. The American Statistician, 50(2):127–139, May 1996. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684424.

Hayden:1992:PHS

[HK92] Robert W. Hayden and Farid Kianifard. Preparing high school mathematics teachers to teach statistics. The American Statistician, 46(4):290–295, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685316.

Hayter:1996:NCP

[HL96] A. J. Hayter and W. Liu. A note on the calculation of Pr(X_{1} < X_{2} ⋯ < X_{k}). The American Statistician, 50(4):365, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684935.

Hadi:1998:SCN

[HL98] Ali S. Hadi and Robert F. Ling. Some cautionary notes on the use of principal components regression. The American Statistician, 52(1):15–19, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.amstat.org/

publications/tas/abstracts_98/hadi.html;   http://www.
jstor.org/stable/2685559.

Horton:1999:RSF

[HL99] Nicholas J. Horton and Stuart R. Lipsitz. Review of software to fit generalized estimating equation regression models. The American Statistician, 53(2):160–169, May 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/horton.pdf;
 http://www.jstor.org/stable/2685737.

Hahn:1993:ASI

[HM93] Gerald J. Hahn and William Q. Meeker. Assumptions for statistical inference. The American Statistician, 47(1):1–11, February 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684774.

Hintze:1998:VPB

[HN98] Jerry L. Hintze and Ray D. Nelson. Violin plots: a box plot-density trace synergism. The American Statistician, 52 (2):181–184, May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.amstat.

org/publications/tas/abstracts_98/hintze.html;   http:
//www.jstor.org/stable/2685478.

Hodges:1994:SVQ

[Hod94] James S. Hodges. [Statistical values, quality, and certification]: Invited comment. The American Statistician, 48(2):75, May 1994.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684247.

Hoeting:1997:BRB

[Hoe97] Jennifer Hoeting. Book review: [untitled]. The American Statistician, 51(1):93–94, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684702.

Hoerl:1998:BRB

[Hoe98] Roger Hoerl. Book review: The Pocket Statistician: a Practical Guide to Quality Improvement. The American Statistician, 52 (1):84–85, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685575.

Hoffman:1991:BRB

[Hof91] Donna L. Hoffman. Book review: [untitled]. The American Statistician, 45(4):305–311, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684462.

Hogg:1991:SEI

[Hog91] Robert V. Hogg. Statistical education: Improvements are badly needed. The American Statistician, 45(4):342–343, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684473.

Hogan:1992:PES

[Hog92a] Howard Hogan. The 1990 post-enumeration survey: An overview. The American Statistician, 46(4):261–269, November 1992. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685308.

Hogg:1992:SCS

[Hog92b] Robert V. Hogg. [Statistical consulting is scholarship]: Invited comment. The American Statistician, 46(4):301–302, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685320.

Hogg:1994:CSE

[Hog94] Robert V. Hogg. A core in statistics for engineering students. The American Statistician, 48(4):285–287, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684834.

Hogg:1999:LUCa

[Hog99a] Robert V. Hogg. Let’s use CQI in our statistics programs. The American Statistician, 53(1):7–14, February 1999. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685642.

Hogg:1999:LUCb

[Hog99b] Robert V. Hogg. [Let’s use CQI in our statistics programs]: Reply. The American Statistician, 53(1):27–28, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685648.

Hogg:1999:R

[Hog99c] Robert V. Hogg. Reply. The American Statistician, 53(1):27–??, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Holmes:1995:BRB

[Hol95] Peter Holmes. Book review: [untitled]. The American Statistician, 49(3):276, August 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684200.

Holcomb:1999:BRB

[Hol99] John P. Holcomb. Book reviews: Applied Regression Analysis, 3rd edition and Applied Regression Analysis: A Research Tool, 2nd edition. The American Statistician, 53(2):170–171, May 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685739.

Hombas:1997:WTE

[Hom97] V. C. Hombas. Waiting time and expected waiting time: Paradoxical situations. The American Statistician, 51(2):130–133, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.amstat.org/publications/tas/
abstracts/hombas.html;   http://www.jstor.org/stable/2685403.

Hombas:1998:LE

[Hom98] V. C. Hombas. Letters to the editor. The American Statistician, 52(2):203, May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685486.

Hosking:1992:MME

[Hos92] J. R. M. Hosking. Moments or L moments? an example comparing two measures of distributional shape. The American Statistician, 46(3):186–189, August 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685210.

Houghton:1997:PEF

[Hou97] Dominique Houghton. Packages for estimating finite mixtures: a review. The American Statistician, 51(2):194–??, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.amstat.org/publications/tas/
abstracts/haughton.html.

Hedayat:1998:EUS

[HR98] A. S. Hedayat and Weining Z. Robieson. Exclusion of an undesirable sample from the support of a simple random sample. The American Statistician, 52(1):41–43, February 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts_
98/hedayat.html;   http://www.jstor.org/stable/2685565.

Haunsperger:1991:LCS

[HS91] Deanna B. Haunsperger and Donald G. Saari. The lack of consistency for statistical decision procedures. The American Statistician, 45(3):252–255, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684305.

Hettmansperger:1992:CNM

[HS92] Thomas P. Hettmansperger and Simon J. Sheather. A cautionary note on the method of least median squares. The American Statistician, 46(2):79–83, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684169.

Harville:1994:HCA

[HS94] David A. Harville and Michael H. Smith. The home-court advantage: How large is it and does it vary from team to team? The American Statistician, 48(1):22–28, February 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685080.

Hunter:1994:SVQ

[HSF^{+}94] J. Stuart Hunter, Stephen M. Stigler, Stephen E. Fienberg, James Boen, and Joseph B. Kadane. [Statistical values, quality, and certification]: Invited comment. The American Statistician, 48(2):70, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684241.

Hsieh:1991:SSS

[Hsi91] F. Y. Hsieh. SSIZE: a sample size program for clinical and epidemiologic studies. The American Statistician, 45(4):338, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684466.

Hsieh:1995:CNA

[Hsi95] Frank Y. Hsieh. A cautionary note on the analysis of extreme data with Cox regression. The American Statistician, 49(2):226–228, May 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684645.

Hurvich:1990:IMS

[HT90] Clifford M. Hurvich and Chih-Ling Tsai. The impact of model selection on inference in linear regression. The American Statistician, 44(3):214–217, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685338.

Hu:1992:EST

[Hu92] Inchi Hu. An elementary solution to a traffic problem. The American Statistician, 46(2):120–121, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684179.

Hunter:1998:CDI

[Hun98] J. Stuart Hunter. Cuthbert Daniel: Industrial statistician. The American Statistician, 52(2):89–92, May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685465.

Hutchinson:1993:SRF

[Hut93] T. P. Hutchinson. The seventh-root formula for a trivariate normal probability. The American Statistician, 47(2):102–103, May 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685186.

Huzurbazar:1999:PSA

[Huz99] S. Huzurbazar. Practical saddlepoint approximations. The American Statistician, 53(3):225–232, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2686101.

Hoaglin:1995:CLS

[HV95] David C. Hoaglin and Paul F. Velleman. A critical look at some analyses of major league baseball salaries. The American Statistician, 49(3):277–285, August 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684201.

Hyndman:1996:CGH

[Hyn96] Rob J. Hyndman. Computing and graphing highest density regions. The American Statistician, 50(2):120–126, May 1996. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684423.

Ibrahim:1997:PPP

[Ibr97] Joseph G. Ibrahim. On properties of predictive priors in linear models. The American Statistician, 51(4):333–337, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.amstat.org/publications/tas/
abstracts/ibrahim.html;   http://www.jstor.org/stable/
2685901.

Iglewicz:1998:SIS

[Igl98] Boris Iglewicz. Selected information on the statistics profession. The American Statistician, 52(4):289–294, November 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts_
98/iglewicz.html;   http://www.jstor.org/stable/2685429.

Imrey:1994:R

[Imr94a] Peter B. Imrey. Reply. The American Statistician, 48(2):82–??, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Imrey:1994:SVQa

[Imr94b] Peter B. Imrey. Statistical values, quality, and certification. The American Statistician, 48(2):65–70, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684240.

Imrey:1994:SVQb

[Imr94c] Peter B. Imrey. [Statistical values, quality, and certification]: Reply. The American Statistician, 48(2):82–87, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684252.

Inman:1990:ASS

[Inm90] Henry F. Inman. The ASA section on statistical education: a historical note. The American Statistician, 44(2):90–93, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684136.

Inman:1994:KPR

[Inm94] Henry F. Inman. Karl Pearson and R. A. Fisher on statistical tests: a 1935 exchange from Nature. The American Statistician, 48(1):2–11, February 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Jammalamadaka:1999:BRA

[Jam99] Sreenivasa Rao Jammalamadaka. Book review: Against the Gods: The Remarkable Story of Risk. The American Statistician, 53(2): 171, May 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685740.

Jeske:1994:IGM

[Jes94] Daniel R. Jeske. Illustrating the Gauss–Markov theorem. The American Statistician, 48(3):237–238, August 1994. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684724.

Jolliffe:1996:BBD

[JH96] Ian T. Jolliffe and Peter B. Hope. Bounded bivariate distributions with nearly normal marginals. The American Statistician, 50(1):17–20, February 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685038.

Johnson:1990:RWA

[JK90] Norman L. Johnson and Samuel Kotz. Randomly weighted averages: Some aspects and extensions. The American Statistician, 44 (3):245–249, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685351.

Jong:1999:RBP

[JK99] Jyh-Cherng Jong and Samuel Kotz. On a relation between principal components and regression analysis. The American Statistician, 53(4):349–351, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2686055.

Jovanovic:1997:LRT

[JL97] B. D. Jovanovic and P. S. Levy. A look at the rule of three. The American Statistician, 51(2):137–139, May 1997. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.amstat.org/publications/tas/abstracts/jovanovic.
html;   http://www.jstor.org/stable/2685405.

Johnson:1996:BRB

[Joh96] Thomas Johnson. Book review: [untitled]. The American Statistician, 50(1):95, February 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685051.

Johnson:1999:BRB

[Joh99] Roger W. Johnson. Book review: Applied Regression Analysis and Multivariable Methods, 3rd edition. The American Statistician, 53(3):292–293, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2686116.

Jolliffe:1995:SSC

[Jol95] Ian T. Jolliffe. Sample sizes and the central limit theorem: The Poisson distribution as an illustration. The American Statistician, 49(3):269, August 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684197.

Jones:1999:DRA

[Jon99] M. C. Jones. Distributional relationships arising from simple trigonometric formulas. The American Statistician, 53(2):99–102, May 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685725.

Jorgensen:1994:TFI

[Jor94] Murray A. Jorgensen. Tail functions and iterative weights in binary regression. The American Statistician, 48(3):230–234, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684722.

Joseph:1993:BRB

[Jos93] Lawrence Joseph. Book review: [untitled]. The American Statistician, 47(4):307, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685293.

Jones:1992:DIF

[JR92] M. C. Jones and John A. Rice. Displaying the important features of large collections of similar curves. The American Statistician, 46(2):140–145, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684184.

Johnston:1995:SAR

[JSM95] Roger G. Johnston, Shayla D. Schroder, and A. Rajika Mallawaaratchy. Statistical artifacts in the ratio of discrete quantities. The American Statistician, 49(3):285–291, August 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684202.

Johnson:1996:CIC

[JT96] Bruce R. Johnson and Benjamin J. Tilly. On the construction of independence counterexamples. The American Statistician, 50 (1):14–16, February 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685037.

Juola:1993:MSC

[Juo93] R. C. Juola. More on shortest confidence intervals. The American Statistician, 47(2):117–119, May 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685191.

Koziol:1997:GTD

[KAB^{+}97] James A. Koziol, Joel E. Alexander, Lance O. Bauer, Samuel Kuperman, Sandra Morzorati, Sean J. O’Connor, John Rohrbaugh, Bernice Porjesz, Henri Begleiter, and John Polich. A graphical technique for displaying correlation matrices. The American Statistician, 51(4):301–304, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685895.

Kadane:1994:SVQ

[Kad94] Joseph B. Kadane. [Statistical values, quality, and certification]: Invited comment. The American Statistician, 48(2):73–74, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684245.

Kafadar:1994:SVQ

[Kaf94] Karen Kafadar. [Statistical values, quality, and certification]: Invited comment. The American Statistician, 48(2):79–82, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684251.

Kafadar:1997:BRB

[Kaf97] Karen Kafadar. Book review: [untitled]. The American Statistician, 51(4):374–375, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685909.

Kafadar:1998:BRB

[Kaf98] Karen Kafadar. Book review: Statistics and Public Policy. The American Statistician, 52(1):86, February 1998. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685578.

Kafadar:1999:BRBb

[Kaf99a] Karen Kafadar. Book review: [untitled]. The American Statistician, 53(2):170, May 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685738.

Kafadar:1999:BRS

[Kaf99b] Karen Kafadar. Book review: Statistics: Concepts and Controversies, 4th edition. The American Statistician, 53(2):170, May 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kahn:1991:LSA

[Kah91] Jeffrey A. Kahn. @Library — statistical applications — technical description (for Lotus 1-2-3 and Symphony). The American Statistician, 45(2):156, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684387.

Kasparek:1994:BRB

[Kas94] Linda Kasparek. Book review: [untitled]. The American Statistician, 48(1):59, February 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685092.

Kastner:1998:BRB

[Kas98] William Kastner. Book review: Statistical Quality Control Using the SAS System. The American Statistician, 52(1):85, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685576.

Kazempour:1995:ISI

[Kaz95] Kazem Kazempour. Impact of stratification imbalance on probability of type I error. The American Statistician, 49(2):170–174, May 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684632.

Koenker:1993:CFP

[KB93] Roger Koenker and Gib Bassett. Correction: Four (pathological) examples in asymptotic statistics. The American Statistician, 47(1):88, February 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684795.

Kass:1998:MCM

[KCGN98] Robert E. Kass, Bradley P. Carlin, Andrew Gelman, and Radford M. Neal. Markov chain Monte Carlo in practice: A roundtable discussion. The American Statistician, 52(2):93–100, May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts_
98/kass.html;   http://www.jstor.org/stable/2685466.

Kreft:1994:BRB

[KDV94] Ita G. G. Kreft, Jan De Leeuw, and Rien Van Der Leeden. Book review: Review of Five Multilevel Analysis Programs: BMDP-5V, GENMOD, HLM, ML3, VARCL. The American Statistician, 48 (4):324–335, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684845.

Keller:1994:CPD

[Kel94] Joseph B. Keller. A characterization of the Poisson distribution and the probability of winning a game. The American Statistician, 48(4):294–298, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684837.

Kettenring:1995:WIN

[Ket95] Jon R. Kettenring. What industry needs. The American Statistician, 49(1):2–4, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684799.

Korn:1990:STR

[KG90] Edward L. Korn and Barry I. Graubard. Simultaneous testing of regression coefficients with complex survey data: Use of Bonferroni t statistics. The American Statistician, 44(4):270–276, November 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684345.

Korn:1995:EDW

[KG95] Edward L. Korn and Barry I. Graubard. Examples of differing weighted and unweighted estimates from a sample survey. The American Statistician, 49(3):291–295, August 1995. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684203.

Korn:1998:SSD

[KG98] Edward L. Korn and Barry I. Graubard. Scatterplots with survey data. The American Statistician, 52(1):58–69, February 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts_
98/korn.html;   http://www.jstor.org/stable/2685570.

Khamis:1991:MCT

[Kha91] H. J. Khamis. Manual computations — A tool for reinforcing concepts and techniques. The American Statistician, 45(4):294–299, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684458.

Khattree:1992:CEP

[Kha92] Ravindra Khattree. Comparing estimators for population variance using Pitman nearness. The American Statistician, 46(3):214–217, August 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685218.

Khuri:1993:NSC

[Khu93] André I. Khuri. A note on Scheffé’s confidence intervals. The American Statistician, 47(3):176–178, August 1993. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684971.

King:1995:BRB

[Kin95] Bruce King. Book review: [untitled]. The American Statistician, 49(1):113, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684824.

Kirk:1991:SCU

[Kir91] Roger E. Kirk. Statistical consulting in a university: Dealing with people and other challenges. The American Statistician, 45 (1):28–34, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685235.

Kedem:1991:LE

[KJM^{+}91] Benjamin Kedem, Lyle V. Jones, Vedula N. Murty, H. A. David, Colin Mallows, Sudhakar Dharmadhikari, Colm O’Cinneide, Daniel J. Gans, Carol A. Markowski, Edward P. Markowski, Hatem A. Howlader, Richard A. Groeneveld, Donald T. Searls, Gilbert W. Basset, Jr., Xuming He, Nicholas R. Farnum, Neil S. Fleming, Gregory Piepel, Bulent Uyar, and Orhan Erdem. Letters to the editor. The American Statistician, 45(3):256–262, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684306.

Kemp:1991:WDD

[KK91] Adrienne W. Kemp and C. David Kemp. Weldon’s dice data revisited. The American Statistician, 45(3):216–222, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684294.

Kafadar:1999:BRBa

[KK99a] Karen Kafadar and James R. Koehler. Book review: [untitled]. The American Statistician, 53(1):86–87, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685660.

Kafadar:1999:BRM

[KK99b] Karen Kafadar and James R. Koehler. Book review: Modern Applied Statistics with S-Plus, 2nd edition. The American Statistician, 53(1):86–??, 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kleinbaum:1995:TSC

[Kle95] David G. Kleinbaum. Teaching short courses in biostatistics and epidemiology. The American Statistician, 49(2):200–203, May 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684639.

Koch:1990:ACS

[KLM90] Gary G. Koch, Fred C. Leone, and Robert L. Mason. ASA celebrates sesquicentennial. The American Statistician, 44(2):113–115, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684143.

Kopas:1992:PIE

[KM92] Dale A. Kopas and Paul R. McAllister. Process improvement exercises for the chemical industry. The American Statistician, 46 (1):34–41, February 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684409.

Khamis:1994:OUS

[KM94] Harry J. Khamis and Barbara L. Mann. Outreach at a University Statistical Consulting Center. The American Statistician, 48 (3):204–207, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684717.

Knusel:1994:PPD

[Knü94] Leo Knüsel. The prediction problem as the dual form of the twosample problem with applications to the Poisson and the binomial distribution. The American Statistician, 48(3):214–219, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684720.

Koehler:1997:BRB

[Koe97a] James R. Koehler. Book review: [untitled]. The American Statistician, 51(2):206–207, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685421.

Koehn:1997:BRB

[Koe97b] Uwe Koehn. Book review: [untitled]. The American Statistician, 51(3):295, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684910.

Konold:1995:CCF

[Kon95] Clifford Konold. Confessions of a coin flipper and would-be instructor. The American Statistician, 49(2):203–209, May 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684640.

Koonce:1990:BRB

[Koo90] Kenneth L. Koonce. Book review: [untitled]. The American Statistician, 44(1):49–52, February 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684959.

Koopmans:1993:NUM

[Koo93] Lambert H. Koopmans. A note on using the moment generating function to teach the laws of large numbers. The American Statistician, 47(3):199–202, August 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684977.

Kooperberg:1997:BRB

[Koo97] Charles Kooperberg. Book review: [untitled]. The American Statistician, 51(1):98, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684710.

Korn:1990:PPP

[Kor90] Edward L. Korn. Projecting power from a previous study: Maximum likelihood estimation. The American Statistician, 44(4): 290–292, November 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684350.

Kott:1991:MBL

[Kot91] Phillip S. Kott. A model-based look at linear regression with survey data. The American Statistician, 45(2):107–112, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684369.

Kuczmarski:1999:QPP

[KR99] James G. Kuczmarski and Paul R. Rosenbaum. Quantile plots, partial orders, and financial risk. The American Statistician, 53 (3):239–246, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2686103.

Korn:1991:ERV

[KS91] Edward L. Korn and Richard Simon. Explained residual variation, explained risk, and goodness of fit. The American Statistician, 45(3):201–206, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684290.

Kim:1994:BRB

[KT94] Junghun Kim and Pravin K. Trivedi. Book review: Econometric Time Series Analysis Software: A Review. The American Statistician, 48(4):336–346, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684846.

Kunte:1996:EMP

[KU96] S. Kunte and K. S. Upadhya. Estimating multinomial probabilities. The American Statistician, 50(3):214–216, August 1996. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684657.

Kutner:1994:IRD

[Kut94] Michael Kutner. Invited response to Dallal, G. E. (1992), “The Computer Analysis of Factorial Experiments with Nested Factors,” The American Statistician, 46, 240. The American Statistician, 48(2):147, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684274.

Kutner:1995:I

[Kut95] Michael H. Kutner. Introduction. The American Statistician, 49 (2):190, May 1995. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2684636.

Kvam:1996:UES

[Kva96] Paul H. Kvam. Using exam scores to estimate the prevalence of classroom cheating. The American Statistician, 50(3):238–242, August 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684662.

Lackritz:1990:SEP

[Lac90] James R. Lackritz. Salary evaluation for professional baseball players. The American Statistician, 44(1):4–8, February 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684947.

Lamble:1990:FP

[Lam90] Andrea P. Lamble. Fig. P. The American Statistician, 44(3):243, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685348.

Lam:1994:IRD

[Lam94a] Ming-Long Lam. Invited response to Dallal, G. E. (1992), “The Computer Analysis of Factorial Experiments with Nested Factors,” The American Statistician, 46, 240. The American Statistician, 48(2):142–144, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684271.

LaMotte:1994:NRI

[LaM94b] Lynn Roy LaMotte. A note on the role of independence in t statistics constructed from linear statistics in regression models. The American Statistician, 48(3):238–240, August 1994. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684725.

Landwehr:1997:ER

[Lan97] James M. Landwehr. Editor’s report. The American Statistician, 51(1):2, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.amstat.org/

ed_rept.html;   http://www.jstor.org/stable/2684682.

Landwehr:1998:ASE

[Lan98] James Landwehr. The American Statistician Editor’s Report for 1997. The American Statistician, 52(3):288, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685950.

Landwehr:1999:ASE

[Lan99a] James M. Landwehr. The American Statistician, Editor’s Report for 1998. The American Statistician, 53(2):175, May 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685744.

Landwehr:1999:ER

[Lan99b] James M. Landwehr. Editor’s report. The American Statistician, 53(4):297, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2686046.

Larson:1992:AVJ

[Lar92] David A. Larson. Analysis of variance with just summary statistics as input. The American Statistician, 46(2):151–152, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684186.

Larntz:1994:ER

[Lar94] Kinley Larntz. Editor’s report. The American Statistician, 48(1):1, February 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685076.

Larntz:1996:ER

[Lar96] Kinley Larntz. Editor’s report. The American Statistician, 50 (4):283, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684921.

Lawrance:1996:DEW

[Law96] A. J. Lawrance. A design of experiments workshop as an introduction to statistics. The American Statistician, 50(2):156–158, May 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684428.

Ledolter:1996:EIM

[LC96] Johannes Ledolter and K. S. Chan. Evaluating the impact of the 65 mph maximum speed limit on Iowa rural interstates. The American Statistician, 50(1):79–85, February 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685048.

Levin:1999:OHC

[LC99] Bruce Levin and Xun Chen. Is the one-half continuity correction used once or twice to derive a well-known approximate sample size formula to compare two independent binomial distributions? The American Statistician, 53(1):62–66, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685654.

Lane:1994:SCI

[LD94] Thomas P. Lane and William H. DuMouchel. Simultaneous confidence intervals in multiple regression. The American Statistician, 48(4):315–321, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684843.

Ludbrook:1998:WPT

[LD98] John Ludbrook and Hugh Dudley. Why permutation tests are superior to t and F tests in biomedical research. The American Statistician, 52(2):127–132, May 1998. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.amstat.org/publications/tas/abstracts_98/ludbrook.
html;   http://www.jstor.org/stable/2685470.

Lazaro:1994:IRD

[LDL^{+}94] Carlos G. Lazaro, David DeLong, Ming-Long Lam, William H. Rogers, and Patrick J. Fleury. Invited responses to Dallal, G. E. (1992), “The Computer Analysis of Factorial Experiments With Nested Factors,” The American Statistician, 46, 240. The American Statistician, 48(2):140–141, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684269.

Lindsay:1993:DRS

[LE93] R. Murray Lindsay and A. S. C. Ehrenberg. The design of replicated studies. The American Statistician, 47(3):217–228, August 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684982.

Ledolter:1995:PIS

[Led95] Johannes Ledolter. Projects in introductory statistics courses. The American Statistician, 49(4):364–367, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684575.

Lee:1993:GRB

[Lee93] A. J. Lee. Generating random binary deviates having fixed marginal distributions and specified degrees of association. The American Statistician, 47(3):209–215, August 1993. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684980.

Lee:1995:GDO

[Lee95] Yoong-Sin Lee. Graphical demonstration of an optimality property of the median. The American Statistician, 49(4):369–372, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684577.

Lehoczky:1995:MSP

[Leh95] John Lehoczky. Modernizing statistics Ph.D. programs. The American Statistician, 49(1):12–17, February 1995. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684803.

Lenth:1998:BRB

[Len98] Russell V. Lenth. Book review: [untitled]. The American Statistician, 52(4):368–369, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685446.

Lesser:1995:BRB

[Les95] Virginia M. Lesser. Book review: [untitled]. The American Statistician, 49(4):398–399, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684587.

Levine:1992:BRB

[Lev92] David M. Levine. Book review: [untitled]. The American Statistician, 46(3):226, August 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685220.

Levin:1993:IMR

[Lev93] Joel R. Levin. An improved modification of a regression-toward-the-mean demonstration. The American Statistician, 47(1):24–26, February 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684777.

Levin:1997:LBM

[Lev97] Bruce Levin. A lower bound for the Mantel–Haenszel one degree of freedom chi-squared statistic in 1: m matched samples. The American Statistician, 51(4):318–320, November 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts/
levin.html;   http://www.jstor.org/stable/2685898.

Ley:1996:PDU

[Ley96] Eduardo Ley. On the peculiar distribution of the U.S. stock indexes’ digits. The American Statistician, 50(4):311–313, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684926.

Loyer:1993:LE

[LFL^{+}93] Milton W. Loyer, Jeremy D. Finn, Richard S. Lehman, David A. Larson, Tim Kramer, Alan E. Gelfand, Adrian F. M. Smith, N. R. Chaganty, Peter Nemenyi, David J. Hand, David A. Binder, Ronald Christensen, Jessica Utts, Gautam M. Appa, Ailsa H. Land, Thomas P. Hettmansperger, Simon J. Sheather, and Peter J. Rousseeuw. Letters to the editor. The American Statistician, 47(2):157–163, May 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685205.

Lichtenberg:1990:AVL

[Lic90] Frank R. Lichtenberg. Aggregation of variables in least-squares regression. The American Statistician, 44(2):169–171, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684161.

Ling:1992:JSN

[Lin92] Robert F. Ling. Just say no to binomial (and other discrete distributions) tables. The American Statistician, 46(1):53–54, February 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684413.

Link:1993:STU

[Lin93] William A. Link. Simulation testing of unbiasedness of variance estimators. The American Statistician, 47(2):132–134, May 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685196.

Lindley:1997:DBB

[Lin97] D. V. Lindley. Discussion — Bayes for beginners? some reasons to hesitate. The American Statistician, 51(3):265–266, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684900.

Lyles:1999:NCI

[LK99] Robert H. Lyles and Lawrence L. Kupper. A note on confidence interval estimation in measurement error adjustment. The American Statistician, 53(3):247–253, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2686104.

Lesser:1995:FPG

[LKNP95] Martin L. Lesser, Nina E. Kohn, Barbara A. Napolitano, and Savita Pahwa. The FU-PLOT: a graphical method for visualizing the timing of follow-up in longitudinal studies. The American Statistician, 49(2):139–143, May 1995. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684627.

Lin:1997:CBT

[LL97] Dennis K. J. Lin and Amy W. Lam. Connections between twolevel factorials and Venn diagrams. The American Statistician, 51(1):49–51, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.amstat.

org/publications/tas/abstracts/lin.html;   http://www.
jstor.org/stable/2684694.

Lynch:1991:BRB

[LLL91] James C. Lynch, J. Richard Landis, and A. Russell Localio. Book review: [untitled]. The American Statistician, 45(2):151–154, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684383.

Lock:1993:BRB

[Loc93] Robin H. Lock. Book review: A Comparison of Five Student Versions of Statistics Packages. The American Statistician, 47(2):136– 145, May 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685198.

Lohr:1994:SVQ

[Loh94] Sharon Lohr. [Statistical values, quality, and certification]: Invited comment. The American Statistician, 48(2):76–77, May 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684248.

Lohr:1995:HDS

[Loh95] Sharon L. Lohr. Hasse diagrams in statistical consulting and teaching. The American Statistician, 49(4):376–381, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684579.

Looney:1995:HUT

[Loo95] Stephen W. Looney. How to use tests for univariate normality to assess multivariate normality. The American Statistician, 49 (1):64–70, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684816.

Loosen:1997:CST

[Loo97] Franz Loosen. A concrete strategy for teaching hypothesis testing. The American Statistician, 51(2):158–163, May 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts/
loosen.html;   http://www.jstor.org/stable/2685410.

Lorenzen:1990:UAC

[Lor90] Gunter Lorenzen. A unified approach to the calculation of growth rates. The American Statistician, 44(2):148–150, May 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684156.

Lindley:1996:LE

[LOS^{+}96] Dennis V. Lindley, W. Scott Overton, Stephen V. Stehman, A. S. Hedayat, Bruno D. Zumbo, Donald W. Zimmerman, Joseph W. McKean, Thomas J. Vidmar, Dan C. Stone, Robert B. Ilderton, Lee R. Abramson, Paul Gottfried, and Clifford Konold. Letters to the editor. The American Statistician, 50(2):197–200, May 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684444.

Low:1998:BRU

[Low98] Elizabeth Low. Book review: Understanding Basic Statistics. The American Statistician, 52(2):198, May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685480.

Lane:1990:WPR

[LRG90] Julia Lane, Russ Ray, and Dennis Glennon. Work profiles of research statisticians. The American Statistician, 44(1):9–13, February 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684948.

Liggett:1990:HSS

[LS90] Walter Liggett and Douglas Splitstone. History of the section on statistics and the environment. The American Statistician, 44(2): 94–95, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2684137.

Lackritz:1997:LE

[LS97a] James R. Lackritz and Oscar Sheynin. Letters to the editor. The American Statistician, 51(2):210, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685425.

Li:1997:BIN

[LS97b] Heng Li and Hal S. Stern. Bayesian inference for nested designs based on Jeffreys’s prior. The American Statistician, 51(3):219–224, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/tas/abstracts/li.html;   http:
//www.jstor.org/stable/2684891.

Lavine:1999:BFW

[LS99a] Michael Lavine and Mark J. Schervish. Bayes factors: What they are and what they are not. The American Statistician, 53(2):119– 122, May 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685729.

Louis:1999:BDM

[LS99b] Thomas A. Louis and Wei Shen. [Bayesian data mining in large frequency tables, with an application to the FDA spontaneous reporting system]: Discussion. The American Statistician, 53 (3):196–198, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2686095.

Leemis:1996:CAI

[LT96] Lawrence M. Leemis and Kishor S. Trivedi. A comparison of approximate interval estimators for the Bernoulli parameter. The American Statistician, 50(1):63–68, February 1996. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685046.

Lee:1997:VOD

[LT97] J. Jack Lee and Z. Nora Tu. A versatile one-dimensional distribution plot: The BLiP plot. The American Statistician, 51 (4):353–358, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685905.

Luceno:1997:FES

[Luc97] Alberto Luceño. Further evidence supporting the numerical usefulness of characteristic functions. The American Statistician, 51(3):233–234, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.amstat.org/publications/tas/abstracts/luceno.html;
 http://www.jstor.org/stable/2684893.

Lurie:1995:BRB

[Lur95] Philip M. Lurie. Book review: A Review of Five Statistical Packages for Windows. The American Statistician, 49(1):99–107, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684822.

Lan:1990:LRT

[LW90] Kuang-Kuo Gordon Lan and Janet Wittes. Linear rank tests for survival data: Equivalence of two formulations. The American Statistician, 44(1):23–26, February 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684951.

Loftsgaarden:1998:STC

[LW98] Don O. Loftsgaarden and Ann E. Watkins. Statistics teaching in colleges and universities: Courses, instructors, and degrees in fall 1995. The American Statistician, 52(4):308–314, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts_
98/loft.html;   http://www.jstor.org/stable/2685432.

Lyerla:1996:BRB

[Lye96] Rob Lyerla. Book review: [untitled]. The American Statistician, 50(4):379–380, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684939.

Maatta:1999:BRA

[Maa99] Jon Maatta. Book review: ActivStats 2.0 and The Active Practice of Statistics. The American Statistician, 53(1):85–86, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685658.

MacLennan:1993:IRS

[Mac93] Richard N. MacLennan. Interrater reliability with SPSS for Windows 5.0. The American Statistician, 47(4):292–296, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685289.

Macky:1999:BRC

[Mac99] David W. Macky. Book review: A Course in Mathematical Statistics, 2nd edition. The American Statistician, 53(1):87, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685661.

Madsen:1991:PPP

[Mad91] Richard Madsen. On the probability of a perfect progression. The American Statistician, 45(3):214–216, August 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684293.

Madgett:1998:SUD

[Mad98] Alan Madgett. Some uses for distribution-fitting software in teaching statistics. The American Statistician, 52(3):253–256, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.amstat.org/publications/
tas/abstracts_98/madgett.html;   http://www.jstor.org/
stable/2685938.

Madigan:1999:BDM

[Mad99] David Madigan. [Bayesian data mining in large frequency tables, with an application to the FDA spontaneous reporting system]: Discussion. The American Statistician, 53(3):198–200, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2686096.

Magee:1990:MBW

[Mag90] Lonnie Magee. R^{2} measures based on Wald and likelihood ratio joint significance tests. The American Statistician, 44(3):250–253, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685352.

Magel:1996:ISP

[Mag96] Rhonda C. Magel. Increasing student participation in large introductory statistics classes. The American Statistician, 50(1):51–56, February 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685044.

Magee:1998:NBP

[Mag98] Lonnie Magee. Nonlocal behavior in polynomial regressions. The American Statistician, 52(1):20–22, February 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts_
98/magee.html;   http://www.jstor.org/stable/2685560.

Malone:1994:BRB

[Mal94] Linda C. Malone. Book review: [untitled]. The American Statistician, 48(4):349–350, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684852.

Malone:1997:BRB

[Mal97] Linda C. Malone. Book review: [untitled]. The American Statistician, 51(1):94, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684703.

Mallows:1998:ZP

[Mal98] Colin Mallows. The zeroth problem. The American Statistician, 52(1):1–9, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/tas/abstracts_98/mallows.html;
 http://www.jstor.org/stable/2685557.

Mann:1991:BRB

[Man91] Barbara L. Mann. Book review: [untitled]. The American Statistician, 45(2):149–151, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684382.

Marsh:1994:BRB

[Mar94] Paul L. Marsh. Book review: A Review and Comparison of Five Graphics Programs for PC-DOS and MS-DOS Computers. The American Statistician, 48(1):44–51, February 1994. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685087.

Matthews:1999:EPS

[Mat99] J. N. S. Matthews. Effect of prior specification on Bayesian design for two-sample comparison of a binary outcome. The American Statistician, 53(3):254–256, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2686105.

MacEachern:1994:SGS

[MB94] Steven N. MacEachern and L. Mark Berliner. Subsampling the Gibbs sampler. The American Statistician, 48(3):188–190, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684714.

Minton:1990:MCA

[MBB90] Paul D. Minton, Dwight B. Brock, and Mary B. Brock. Musical celebration of ASA’s sesquicentennial — a personal account. The American Statistician, 44(3):200–203, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685335.

Mee:1991:RTM

[MC91] Robert W. Mee and Tin Chiu Chua. Regression toward the mean and the paired sample t test. The American Statistician, 45(1): 39–42, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685237.

Murdoch:1996:GDL

[MC96] D. J. Murdoch and E. D. Chow. A graphical display of large correlation matrices. The American Statistician, 50(2):178–180, May 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684435.

McConnell:1990:LEP

[McC90] C. R. McConnell. Letter to the Editor: Pocket computer approximation for areas under the standard normal curve. The American Statistician, 44(1):63, February 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684963.

McCullagh:1994:DMG

[McC94] Peter McCullagh. Does the moment-generating function characterize a distribution? The American Statistician, 48(3):208, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684718.

Mccormick:1997:BRB

[Mcc97] David Mccormick. Book review: [untitled]. The American Statistician, 51(3):296–297, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684913.

McCulloch:1998:ARS

[McC98a] B. D. McCulloch. Assessing the reliability of statistical software: Part I. The American Statistician, 52(4):358–??, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.amstat.org/publications/tas/
abstracts_98/mccullogh.html.

McCullough:1998:ARS

[McC98b] B. D. McCullough. Assessing the reliability of statistical software: Part I. The American Statistician, 52(4):358–366, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685442.

McCullough:1999:ARS

[McC99] B. D. McCullough. Assessing the reliability of statistical software: Part II. The American Statistician, 53(2):149–159, May 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/mccull.pdf;
 http://www.jstor.org/stable/2685736.

McDonald:1999:SSS

[McD99] Gary C. McDonald. Shaping statistics for success in the 21st Century: The needs of industry. The American Statistician, 53 (3):203–207, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2686098.

Morgan:1991:LMDa

[MCDD91a] J. P. Morgan, N. R. Chaganty, R. C. Dahiya, and M. J. Doviak. Let’s make a deal: The Player’s dilemma. The American Statistician, 45(4):284–287, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www. jstor.org/stable/2684453. With comment by R. G. Seymann (pp. 287–288) and rejoinder (p. 289).

Morgan:1991:LMDb

[MCDD91b] J. P. Morgan, N. R. Chaganty, R. C. Dahiya, and M. J. Doviak. [Let’s make a deal: The Player’s dilemma]: Rejoinder. The American Statistician, 45(4):289, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684455.

Moore:1995:SEF

[MCGM95] David S. Moore, George W. Cobb, Joan Garfield, and William Q. Meeker. Statistics education fin de siècle. The American Statistician, 49(3):250–260, August 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684195.

McLeod:1999:NSC

[McL99] A. Ian McLeod. Necessary and sufficient condition for nonsingular Fisher information matrix in ARMA and fractional ARIMA models. The American Statistician, 53(1):71–72, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685656.

McNeil:1992:GPD

[McN92] Don McNeil. On graphing paired data. The American Statistician, 46(4):307–311, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685323.

Mac:1996:HML

[MD96] Robert J. Mac and G. Dawson. How many light bulbs does it take to generate a data set? The American Statistician, 50(3):247–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mintz:1997:OOC

[MD97] Jim Mintz and Wilfrid Dixon. Objection overruled: a comment on Gastwirth, Krieger, and Rosenbaum. The American Statistician, 51(2):117–119, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:/

/www.amstat.org/publications/tas/abstracts/mintz.html;
 http://www.jstor.org/stable/2685400.

Muranaka:1999:LE

[MDCW99] Ken Muranaka, Norman R. Draper, Dennis Cook, and Sanford Weisberg. Letters to the editor. The American Statistician, 53 (3):295–296, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2686119.

Meeker:1995:TAA

[ME95] William Q. Meeker and Luis A. Escobar. Teaching about approximate confidence regions based on maximum likelihood estimation. The American Statistician, 49(1):48–53, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684811.

Mehta:1991:SSP

[Meh91] Cyrus R. Mehta. StatXact: a statistical package for exact nonparametric inference. The American Statistician, 45(1):74–75, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685246.

Mayo:1997:ESB

[MG97] Matthew S. Mayo and J. Brian Gray. Elemental subsets: The building blocks of regression. The American Statistician, 51(2):122–129, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.amstat.org/publications/tas/abstracts/mayo.html;
http://www.jstor.org/stable/2685402.

Molenberghs:1999:NMC

[MGLK99] Geert Molenberghs, Els J. T. Goetghebeur, Stuart R. Lipsitz, and Michael G. Kenward. Nonrandom missingness in categorical data: Strengths and limitations. The American Statistician, 53(2):110– 118, May 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685728.

Mehta:1993:EPC

[MH93] Cyrus R. Mehta and Joan F. Hilton. Exact power of conditional and unconditional tests: Going beyond the 2 × 2 contingency table. The American Statistician, 47(2):91–98, May 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685184.

Makuch:1990:SCC

[MHT90] William M. Makuch, Gerald J. Hahn, and William T. Tucker. A statistical computing curriculum to meet industrial needs. The American Statistician, 44(1):42–49, February 1990. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684958.

Midha:1996:BRB

[Mid96] Chand Midha. Book review: [untitled]. The American Statistician, 50(1):97, February 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685054.

Miller:1996:CES

[Mil96] Alan J. Miller. The convergence of Efroymson’s stepwise regression algorithm. The American Statistician, 50(2):180–181, May 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684436.

Minton:1992:SCS

[Min92] Paul D. Minton. [Statistical consulting is scholarship]: Invited comment. The American Statistician, 46(4):300–301, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685319.

McDonald:1999:LE

[MJ99] Gary C. McDonald and M. C. Jones. Letters to the editor. The American Statistician, 53(4):393, November 1999. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2686066.

Myers:1992:RSA

[MKV92] Raymond H. Myers, Andre I. Khuri, and Geoffrey Vining. Response surface alternatives to the Taguchi Robust Parameter Design Approach. The American Statistician, 46(2):131–139, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684183.

Moser:1992:ECF

[ML92a] Barry Kurt Moser and Yuh-Ren Lin. Equivalence of the corrected F test and the weighted least squares procedure. The American Statistician, 46(2):122–124, May 1992. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684180.

Moser:1992:ECT

[ML92b] Barry Kurt Moser and Yuh-Ren Lin. Equivalence of the corrected F test and the weighted least squares procedure. The American Statistician, 46(2):122–124, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Meng:1999:LE

[ML99] Xiao-Li Meng and Willard C. Losinger. Letters to the editor. The American Statistician, 53(2):173–174, May 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685742.

Markowski:1990:CEP

[MM90] Carol A. Markowski and Edward P. Markowski. Conditions for the effectiveness of a preliminary test of variance. The American Statistician, 44(4):322–326, November 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684360.

Marasinghe:1996:UGS

[MMCsS96] Mervyn G. Marasinghe, William Q. Meeker, Dianne Cook, and Tae sung Shin. Using graphics and simulation to teach statistical concepts. The American Statistician, 50(4):342–351, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684932.

Mason:1990:BHA

[MMR90] R. L. Mason, J. D. McKenzie, Jr., and S. J. Ruberg. A brief history of the American Statistical Association, 1839–1989. The American Statistician, 44(2):68–73, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684130.

Mason:1990:CAS

[MMS90] R. L. Mason, J. D. McKenzie, Jr., and R. K. Spoeri. The chapters of the American statistical association. The American Statistician, 44(2):103–106, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684141.

Mathew:1997:IMD

[MN97] Thomas Mathew and Kenneth Nordström. An inequality for a measure of deviation in linear models. The American Statistician, 51(4):344–349, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685903.

Molnar:1991:LLL

[Mol91] László D. Molnár. LOGLINDI: Log-linear discriminant analysis. The American Statistician, 45(4):339, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684467.

Monti:1995:FED

[Mon95] Katherine L. Monti. Folded empirical distribution function curves — mountain plots. The American Statistician, 49(4):342–345, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684570.

Monahan:1997:BRB

[Mon97a] John F. Monahan. Book review: [untitled]. The American Statistician, 51(3):297, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684915.

Monti:1997:MST

[Mon97b] Katherine L. Monti. Making statistical tables in windows: a comparison of Data Muncher and SAS (PROC TABULATE and PROC REPORT). The American Statistician, 51(1):83–92, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.amstat.org/publications/tas/
abstracts/monti.html;   http://www.jstor.org/stable/2684700.

Moore:1993:PVN

[Moo93] David S. Moore. The place of video in new styles of teaching and learning statistics. The American Statistician, 47(3):172–176, August 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684970.

Moore:1996:BRB

[Moo96] Thomas L. Moore. Book review: [untitled]. The American Statistician, 50(4):380–381, November 1996. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684940.

Moore:1997:BBSa

[Moo97a] David S. Moore. Bayes for beginners? some reasons to hesitate. The American Statistician, 51(3):254–261, August 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts/
moore.html;   http://www.jstor.org/stable/2684897.

Moore:1997:BBSb

[Moo97b] David S. Moore. [Bayes for beginners? some reasons to hesitate]: Reply. The American Statistician, 51(3):272–274, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684905.

Moore:1997:RBB

[Moo97c] David S. Moore. Reply — Bayes for beginners? some reasons to hesitate. The American Statistician, 51(3):272–??, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Morganstein:1991:BRB

[Mor91] David Morganstein. Book review: [untitled]. The American Statistician, 45(4):316–317, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684464.

Morris:1995:MSP

[Mor95a] Carl N. Morris. [Modernizing statistics Ph.D. programs]: Respondent. The American Statistician, 49(1):21–23, February 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684805.

Morris:1995:R

[Mor95b] Carl N. Morris. Respondent. The American Statistician, 49(1): 21–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Morgan:1998:RES

[Mor98] Walter T. Morgan. A review of eight statistics software packages for general use. The American Statistician, 52(1):70–82, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685572.

Mosteller:1997:LSS

[Mos97] Frederick Mosteller. Lessons from sports statistics. The American Statistician, 51(4):305–310, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685896.

Miller:1998:LE

[MOS^{+}98] Dave P. Miller, Doug Oman, Jeffrey S. Simonoff, Carles M. Cuadras, Ali S. Hadi, Robert F. Ling, and H. A. David. Letters to the editor. The American Statistician, 52(4):370–371, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685448.

Morrell:1997:LTL

[MPB97] Christopher H. Morrell, Jay D. Pearson, and Larry J. Brant. Linear transformations of linear mixed-effects models. The American Statistician, 51(4):338–343, November 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts/
morrell.html;   http://www.jstor.org/stable/2685902.

Moser:1992:HVT

[MS92] Barry K. Moser and Gary R. Stevens. Homogeneity of variance in the two-sample means test. The American Statistician, 46 (1):19–21, February 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684403.

Moschopoulos:1994:EGC

[MS94] Panagis Moschopoulos and Joan G. Staniswalis. Estimation given conditionals from an exponential family. The American Statistician, 48(4):271–275, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684831.

McCulloch:1995:IDU

[MS95] C. E. McCulloch and S. R. Searle. On an identity derived from unbiasedness in linear models. The American Statistician, 49 (1):39–42, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684809.

Moser:1998:ASS

[MS98] Barry Kurt Moser and Julia K. Sawyer. Algorithms for sums of squares and covariance matrices using Kronecker products. The American Statistician, 52(1):54–57, February 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts_
98/moser.html;   http://www.jstor.org/stable/2685569.

McLean:1991:UAM

[MSS91] Robert A. McLean, William L. Sanders, and Walter W. Stroup. A unified approach to mixed linear models. The American Statistician, 45(1):54–64, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685241.

Murray:1994:DHD

[Mur94] Michael P. Murray. A drunk and her dog: An illustration of cointegration and error correction. The American Statistician, 48 (1):37–39, February 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685084.

Murphy:1997:HRS

[Mur97] James R. Murphy. How to read the statistical methods literature: A guide for students. The American Statistician, 51(2):155–157, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.amstat.org/publications/tas/abstracts/murphy.html;
 http://www.jstor.org/stable/2685409.

McKean:1994:CTR

[MV94] Joseph W. McKean and Thomas J. Vidmar. A comparison of two rank-based methods for the analysis of linear models. The American Statistician, 48(3):220–229, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684721.

Mehta:1992:CEM

[MW92] Cyrus R. Mehta and Stephen J. Walsh. Comparison of exact, mid-p, and Mantel–Haenszel confidence intervals for the common odds ratio across several 2 × 2 contingency tables. The American Statistician, 46(2):146–150, May 1992. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684185.

MW:1995:BRB

[MW95] MW. Book review: [untitled]. The American Statistician, 49(3): 322, August 1995. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2684213.

Martz:1992:RCF

[MZ92] H. F. Martz and W. J. Zimmer. The risk of catastrophic failure of the solid rocket boosters on the Space Shuttle. The American Statistician, 46(1):42–47, February 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684410.

Nabel:1997:BRB

[Nab97] Michael Nabel. Book review: [untitled]. The American Statistician, 51(1):97, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684708.

Naghshpour:1990:BRB

[Nag90] Shahdad Naghshpour. Book review: [untitled]. The American Statistician, 44(1):52–54, February 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684960.

Nagler:1998:BRS

[Nag98] Jonathan Nagler. Book review: Statistical Computing Environments for Social Research. The American Statistician, 52(2):198– 199, May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685482.

Nash:1992:SSI

[Nas92] John C. Nash. Statistical shareware: Illustrations from regression techniques. The American Statistician, 46(4):311–318, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685324.

Nash:1994:EN

[Nas94a] John Nash. Editor’s notes. The American Statistician, 48(4): 322–323, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684844.

Nash:1994:TIS

[Nas94b] John C. Nash. Tools for including statistical graphics in application programs. The American Statistician, 48(1):52–57, February 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685088.

Nash:1995:BRB

[Nas95] John C. Nash. Book review: Computer Algebra Systems: DE-RIVE(R). The American Statistician, 49(1):93–99, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684821.

Naveau:1997:CBC

[Nav97a] Philippe Naveau. Comparison between the Chernoff and factorial moment bounds for discrete random variables. The American Statistician, 51(1):40–41, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.amstat.org/publications/tas/abstracts/naveau.html;
 http://www.jstor.org/stable/2684691.

Navidi:1997:GIE

[Nav97b] William Navidi. A graphical illustration of the EM algorithm. The American Statistician, 51(1):29–31, February 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts/
navidi.html;   http://www.jstor.org/stable/2684688.

Nayak:1994:PNC

[Nay94] Tapan K. Nayak. Pitman nearness comparison of some estimators of population variance. The American Statistician, 48(2):99–102, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684255.

Neill:1993:BRB

[Nei93] James W. Neill. Book review: [untitled]. The American Statistician, 47(3):234, August 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684984.

Neidert:1996:BRB

[Nei96] Sharon Neidert. Book review: [untitled]. The American Statistician, 50(2):195–196, May 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684443.

Nelder:1990:NPL

[Nel90] J. A. Nelder. Nearly parallel lines in residual plots. The American Statistician, 44(3):221–222, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685341.

Nelder:1998:STR

[Nel98a] John A. Nelder. The selection of terms in response-surface models: How strong is the weak-heredity principle? The American Statistician, 52(4):315–318, November 1998. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.amstat.org/publications/tas/abstracts_98/nelder.
html;   http://www.jstor.org/stable/2685433.

Nelsen:1998:CRL

[Nel98b] Roger B. Nelsen. Correlation, regression lines, and moments of inertia. The American Statistician, 52(4):343–345, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts_
98/nelsen.html;   http://www.jstor.org/stable/2685438.

Neuts:1992:IPP

[Neu92] Marcel F. Neuts. An invariance property of the Poisson process. The American Statistician, 46(4):276–278, November 1992. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685311.

Newton:1992:NDSa

[New92a] H. Joseph Newton. New developments in statistical computing. The American Statistician, 46(2):153–154, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684187.

Newton:1992:NDSb

[New92b] H. Joseph Newton. New developments in statistical computing. The American Statistician, 46(4):318–319, November 1992. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685325.

Newton:1999:LUC

[New99] H. Joseph Newton. [Let’s use CQI in our statistics programs]: Discussion. The American Statistician, 53(1):15–16, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685643.

Nick:1994:BRB

[Nic94a] Todd G. Nick. Book review: [untitled]. The American Statistician, 48(3):266, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684733.

Nickerson:1994:CCC

[Nic94b] David M. Nickerson. Construction of a conservative confidence region from projections of an exact confidence region in multiple linear regression. The American Statistician, 48(2):120–124, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684261.

Nickerson:1995:BRB

[Nic95] David M. Nickerson. Book review: [untitled]. The American Statistician, 49(3):321, August 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684210.

Naik:1996:ROT

[NK96] Dayanand N. Naik and Ravindra Khattree. Revisiting Olympic track records: Some practical considerations in the principal component analysis. The American Statistician, 50(2):140–144, May 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684425.

Nelder:1995:CAF

[NL95] John A. Nelder and Peter W. Lane. The computer analysis of factorial experiments: In memorian — Frank Yates. The American Statistician, 49(4):382–385, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684580.

Nash:1998:BRT

[NO98] John C. Nash and Chris Olsen. Book review: The Texas Instruments TI-83 Graphing Calculator. The American Statistician, 52 (3):285–287, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.amstat.org/

publications/tas/ti-83.html;   http://www.jstor.org/stable/
2685949.

Niland:1995:SBC

[NOMLT95] Joyce C. Niland, Tamara L. Odom-Maryon, Jennifer Lee, and Barbara C. Tilley. A survey of biostatistical consulting units throughout North America. The American Statistician, 49(2):183–189, May 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684635.

Norwood:1990:EIC

[Nor90] Janet L. Norwood. Editor’s invited column: ASA sesquicentennial. The American Statistician, 44(2):67, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684129.

Normand:1995:BRB

[Nor95] Sharon-Lise T. Normand. Book review: Meta-Analysis Software: A Comparative Review. The American Statistician, 49(3):298–309, August 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684205.

Nicewander:1997:CCC

[NP97] W. Alan Nicewander and James M. Price. A consonance criterion for choosing sample size. The American Statistician, 51(4):311–317, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/tas/abstracts/nicewander.html;
 http://www.jstor.org/stable/2685897.

Nash:1996:SMK

[NQ96] John C. Nash and Tony K. Quon. Software for modeling kinetic phenomena. The American Statistician, 50(4):368–378, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684937.

Neath:1997:EBI

[NS97] Andrew A. Neath and Francisco J. Samaniego. On the efficacy of Bayesian inference for nonidentifiable models. The American Statistician, 51(3):225–232, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:/

/www.amstat.org/publications/tas/abstracts/neath.html;
 http://www.jstor.org/stable/2684892.

Nolan:1999:TST

[NS99] D. Nolan and T. P. Speed. Teaching statistics theory through applications. The American Statistician, 53(4):370–375, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2686059.

Nunnikhoven:1992:BPS

[Nun92] Thomas S. Nunnikhoven. A birthday problem solution for nonuniform birth frequencies. The American Statistician, 46(4):270–274, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685309.

Nash:1990:NES

[NWS90] John C. Nash and Mary Walker-Smith. NLEX: Examples and software extensions for nonlinear modeling and robust nonlinear estimation. The American Statistician, 44(1):55, February 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684962.

Oakes:1993:NKM

[Oak93] David Oakes. A note on the Kaplan–Meier estimator. The American Statistician, 47(1):39–40, February 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684781.

Obenchain:1991:RRS

[Obe91] Bob Obenchain. Ridge regression systems for MS-DOS personal computers. The American Statistician, 45(3):245–246, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684300.

Obremski:1998:BRB

[Obr98] Tom Obremski. Book review: Statistics for Management, 7th ed. The American Statistician, 52(2):198, May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685481.

OCinneide:1990:MWO

[O’C90] Colm Art O’Cinneide. The mean is within one standard deviation of any median. The American Statistician, 44(4):292–293, November 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684351.

Okunade:1993:CAR

[OCE93] Albert A. Okunade, Cyril F. Chang, and Richard D. Evans. Comparative analysis of regression output summary statistics in common statistical packages. The American Statistician, 47(4):298– 303, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685291.

Oehlert:1992:NDM

[Oeh92] Gary W. Oehlert. A note on the delta method. The American Statistician, 46(1):27–29, February 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684406.

OHaraHines:1995:ECS

[OH95] R. J. O’Hara Hines and W. G. S. Hines. Exploring Cook’s statistic graphically. The American Statistician, 49(4):389–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Okolo:1999:NCP

[Oko99] Abraham Okolo. The Nigerian census: Problems and prospects. The American Statistician, 53(4):321–325, November 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2686050.

Oksanen:1991:SAT

[Oks91] E. H. Oksanen. A simple approach to teaching generalized least squares theory. The American Statistician, 45(3):229–233, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684297.

Ord:1996:BRB

[OL96] Keith Ord and Sam Lowe. Book review: Automatic Forecasting. The American Statistician, 50(1):88–94, February 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685050.

Oldford:1995:PDD

[Old95] R. W. Oldford. A physical device for demonstrating confounding, blocking, and the role of randomization in uncovering a causal relationship. The American Statistician, 49(2):210–216, May 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684641.

Olkin:1992:MFH

[Olk92] Ingram Olkin. A matrix formulation on how deviant an observation can be. The American Statistician, 46(3):205–209, August 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685215.

Overton:1995:HTT

[OS95] W. Scott Overton and Stephen V. Stehman. The Horvitz– Thompson theorem as a unifying perspective for probability sampling: With examples from natural resource sampling. The American Statistician, 49(3):261–268, August 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684196.

ONeill:1999:BDM

[OS99] Robert T. O’Neill and Ana Szarfman. [Bayesian data mining in large frequency tables, with an application to the FDA spontaneous reporting system]: Discussion. The American Statistician, 53(3):190–196, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2686094.

Oster:1998:EFS

[Ost98] Robert A. Oster. An examination of five statistical software packages for epidemiology. The American Statistician, 52(3):267–280, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.amstat.

org/publications/tas/abstracts_98/oster.html;   http:/
/www.jstor.org/stable/2685942.

Ottenbacher:1996:PRR

[Ott96] Kenneth J. Ottenbacher. The power of replications and replications of power. The American Statistician, 50(3):271–275, August 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684673.

OGorman:1991:VSD

[OW91] Thomas W. O’Gorman and Robert F. Woolson. Variable selection to discriminate between two groups: Stepwise logistic regression or stepwise discriminant analysis? The American Statistician, 45(3):187–193, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684288.

Perez-Abreu:1991:PAP

[PA91] Víctor Pérez-Abreu. Poisson approximation to power series distributions. The American Statistician, 45(1):42–45, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685238.

Palacios:1999:RPE

[Pal99] Jose Luis Palacios. The ruin problem via electric networks. The American Statistician, 53(1):67–70, February 1999. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685655.

Parr:1998:BRB

[Par98a] William C. Parr. Book review: [untitled]. The American Statistician, 52(3):283–284, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685947.

Parr:1998:BRP

[Par98b] William C. Parr. Book review: Practical Guide to Experimental Design. The American Statistician, 52(3):283–??, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Paulson:1992:ULG

[Pau92] Richard A. Paulson. Using lottery games to illustrate statistical concepts and abuses. The American Statistician, 46(3):202–204, August 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685214.

Picard:1990:DS

[PB90] Richard R. Picard and Kenneth N. Berk. Data splitting. The American Statistician, 44(2):140–147, May 1990. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684155.

Pal:1992:IML

[PB92] Nabendu Pal and J. Calvin Berry. On invariance and maximum likelihood estimation. The American Statistician, 46(3):209–212, August 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685216.

Posten:1993:BRB

[PD93] Harry O. Posten and Dipak K. Dey. Book review: [untitled]. The American Statistician, 47(4):277–279, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685287.

Peck:1997:BRB

[Pec97] Roxy L. Peck. Book review: [untitled]. The American Statistician, 51(2):208–209, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685423.

Peixoto:1990:PWF

[Pei90] Julio L. Peixoto. A property of well-formulated polynomial regression models. The American Statistician, 44(1):26–30, February 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684952.

Peixoto:1993:IFR

[Pei93] Julio L. Peixoto. Incorrect F formulas in regression analysis. The American Statistician, 47(3):194–197, August 1993. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684975.

Pena:1997:CIS

[Peñ97] Daniel Peña. Combining information in statistical modeling. The American Statistician, 51(4):326–332, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.amstat.org/publications/tas/abstracts/pena.
html;   http://www.jstor.org/stable/2685900.

Percival:1993:TCP

[Per93] Donald B. Percival. Three curious properties of the sample variance and autocovariance for stationary processes with unknown mean. The American Statistician, 47(4):274–276, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685286.

Perry:1998:SPE

[Per98] Robert J. Perry. Statistical practice — estimating strength of the Williamsburg Bridge suspension cables. The American Statistician, 52(3):211–217, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.amstat.

org/publications/tas/abstracts_98/perry.html;   http:/
/www.jstor.org/stable/2685926.

Peskun:1990:NGM

[Pes90] Peter H. Peskun. A note on a general method for obtaining confidence intervals from samples from discrete distributions. The American Statistician, 44(1):31–35, February 1990. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684953.

Peterson:1999:BRB

[Pet99] William P. Peterson. Book review: Interpreting Basic Statistics: a Guide and Workbook Based on Excerpts from Journal Articles. The American Statistician, 53(3):294, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2686118.

Proschan:1995:MCC

[PF95] Michael A. Proschan and Dean A. Follmann. Multiple comparisons with control in a single experiment versus separate experiments: Why do we feel differently? The American Statistician, 49(2):144– 149, May 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684628.

Pearson:1994:KPR

[PFI94] Karl Pearson, R. A. Fisher, and Henry F. Inman. Karl Pearson and R. A. Fisher on statistical tests: a 1935 exchange from Nature. The American Statistician, 48(1):2–11, February 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685077.

Poole:1997:HCA

[PG97] Charles Poole and Sander Greenland. How a court accepted a possible explanation: A comment on Gastwirth, Krieger, and Rosenbaum. The American Statistician, 51(2):112–114, May 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts/
poole.html;   http://www.jstor.org/stable/2685398. See

[GKR94].

Pickle:1997:BRB

[Pic97] Linda Williams Pickle. Book review: [untitled]. The American Statistician, 51(2):206, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685420.

Piepel:1990:MSD

[Pie90] Gregory F. Piepel. MIXSOFT^{T M}: Software for the design and analysis of mixture experiments. The American Statistician, 44 (3):243–244, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685349.

Piegorsch:1992:CLR

[Pie92] Walter W. Piegorsch. Complementary log regression for generalized linear models. The American Statistician, 46(2):94–99, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684172.

PillersDobler:1998:BRM

[Pil98] Carolyn Pillers Dobler. Book reviews: Modern Elementary Statistics (9th ed.), A First Course in Statistics (6th ed.). The American Statistician, 52(3):284–??, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Philips:1995:MBT

[PN95] Thomas K. Philips and Randolph Nelson. The moment bound is tighter than Chernoff’s bound for positive tail probabilities. The American Statistician, 49(2):175–178, May 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684633.

Polansky:1999:UBT

[Pol99] Alan M. Polansky. Upper bounds on the true coverage of bootstrap percentile type confidence intervals. The American Statistician, 53(4):362–369, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2686058.

Posten:1990:RAT

[Pos90] Harry O. Posten. Review of audio tape: The Statistician at Work. The American Statistician, 44(4):302, November 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684355.

Posten:1993:EAN

[Pos93] Harry O. Posten. An effective algorithm for the noncentral beta distribution function. The American Statistician, 47(2):129–131, May 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685195.

Posten:1994:BRB

[Pos94] Harry O. Posten. Book review: [untitled]. The American Statistician, 48(2):129, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684264.

Posten:1996:BRB

[Pos96] Harry O. Posten. Book review: [untitled]. The American Statistician, 50(3):258–259, August 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684669.

Proschan:1998:EUC

[PP98] Michael A. Proschan and Brett Presnell. Expect the unexpected from conditional expectation. The American Statistician, 52 (3):248–252, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.amstat.org/

publications/tas/abstracts_98/proschan.html;   http://
www.jstor.org/stable/2685937.

Park:1996:SMG

[PPS96] Chul Gyu Park, Taesung Park, and Dong Wan Shin. A simple method for generating correlated binary variates. The American Statistician, 50(4):306–310, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684925.

Pena:1994:SCA

[PR94] Edsel A. Peña and Vijay Rohatgi. Some comments about sufficiency and unbiased estimation. The American Statistician, 48 (3):242–243, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684727.

Piepel:1998:MEA

[PR98] Greg Piepel and Trish Redgate. A mixture experiment analysis of the Hald cement data. The American Statistician, 52 (1):23–30, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.amstat.

org/publications/tas/abstracts_98/piepel.html;   http:
//www.jstor.org/stable/2685561.

Prosser:1991:MST

[Pro91] Bob Prosser. ML3: Software for three-level analysis. The American Statistician, 45(2):155–156, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684386.

Petras:1994:BRB

[PRSE94] Michael F. Petras, Ronald R. Remke, Paul R. Saunders, and Philip J. Ernst. Book review: [untitled]. The American Statistician, 48(4):348, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684850.

Pontius:1994:GAT

[PS94] Jeffrey S. Pontius and Rhoda M. Schantz. Graphical analyses of a two-period crossover design. The American Statistician, 48 (3):249–253, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684730.

Parr:1998:DCB

[PS98a] William C. Parr and Marlene A. Smith. Developing case-based business statistics courses. The American Statistician, 52(4): 330–337, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.amstat.org/

publications/tas/abstracts_98/parr.html;   http://www.
jstor.org/stable/2685436.

Pratt:1998:NIS

[PS98b] John W. Pratt and Robert Schlaifer. A new interpretation of the F statistic. The American Statistician, 52(2):141–143, May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts_
98/pratt.html;   http://www.jstor.org/stable/2685472.

Pena:1991:NLE

[PT91] Daniel Peña and George C. Tiao. A note on likelihood estimation of missing values in time series. The American Statistician, 45 (3):212–213, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684292.

Pukelsheim:1994:TSR

[Puk94] Friedrich Pukelsheim. The three sigma rule. The American Statistician, 48(2):88–91, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684253.

Pulskamp:1990:NEB

[Pul90] Richard J. Pulskamp. A note on the estimation of binomial probabilities. The American Statistician, 44(4):293–295, November 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684352.

Penner:1991:MI

[PW91] Ralph Penner and Donald G. Watts. Mining information. The American Statistician, 45(1):4–9, February 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685231.

Ridgway:1993:LEb

[RCU^{+}93] Terry Ridgway, Ronald Christensen, Jessica Utts, Sunil K. Sapra, Shigeru Mase, Thomas Nunnikhoven, Harry M. Schey, Jonathan J. Shuster, Alan B. Cantor, David Oakes, Mark Krailo, Carl J. Huberty, David J. Hand, Milton Keynes, Paul F. Velleman, Leland Wilkinson, Thomas J. Lorenzen, and Richard K. Burdick. Letters to the editor. The American Statistician, 47(4):311–316, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685299.

Read:1998:LAC

[Rea98] K. L. Q. Read. A lognormal approximation for the Collector’s Problem. The American Statistician, 52(2):175–180, May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts_
98/read.html;   http://www.jstor.org/stable/2685477.

Reece:1994:BRB

[Ree94] Jack E. Reece. Book review: [untitled]. The American Statistician, 48(2):172, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684279.

Reiser:1994:BRB

[Rei94] Benjamin Reiser. Book review: [untitled]. The American Statistician, 48(4):348, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684849.

Rencher:1992:ICD

[Ren92] Alvin C. Rencher. Interpretation of canonical discriminant functions, canonical variates, and principal components. The American Statistician, 46(3):217–225, August 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685219.

Riedwyl:1990:RWC

[Rie90] Hans Riedwyl. Rudolf Wolf’s contribution to the Buffon Needle Problem (an early Monte Carlo experiment) and application of least squares. The American Statistician, 44(2):138–139, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684154.

Rappaport:1998:PIS

[RKW98] Karen D. Rappaport, Nouna Kettaneh, and Svante Wold. Perspectives on implementing statistical modeling and design (SMD) in an industrial/chemical environment. The American Statistician, 52(2):152–159, May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.amstat.org/

publications/tas/abstracts_98/rappaport.html;   http:/
/www.jstor.org/stable/2685474.

Ringel:1991:ASD

[RM91] Emanuel Ringel and Charles Mode. APL software for discrete semi-Markov and population process computations and Monte Carlo simulation. The American Statistician, 45(2):160, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684393.

Rousseeuw:1994:SCM

[RM94] Peter J. Rousseeuw and Geert Molenberghs. The shape of correlation matrices. The American Statistician, 48(4):276–279, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684832.

Rosner:1996:MPP

[RMY96] Bernard Rosner, Frederick Mosteller, and Cleo Youtz. Modeling pitcher performance and the distribution of runs per inning in major league baseball. The American Statistician, 50(4):352–360, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684933.

Randall:1996:LE

[RNB^{+}96] John H. Randall, John A. Nelder, James Berger, Inge S. Helland, Murray Aitkin, D. V. Lindley, Edward J. Stanek, III, Charles E. Gates, Susan R. Wilson, Judea Pearl, R. W. Oldford, Y. S. Sathe, Jerome H. Klotz, Lawrence Leemis, Kishor S. Trivedi, and Sunil K. Sapra. Letters to the editor. The American Statistician, 50(4): 382–389, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684941.

Rogers:1994:IRD

[Rog94] William H. Rogers. Invited response to Dallal, G. E. (1992), “The Computer Analysis of Factorial Experiments with Nested Factors,” The American Statistician, 46, 240. The American Statistician, 48(2):144–145, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684272.

Ridgway:1993:LEa

[RON93] Terry R. Ridgway, Cornel G. Ormsby, and Thomas S. Nunnikhoven. Letters to the editor. The American Statistician, 47 (3):237–238, August 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684989.

Rossi:1993:BRB

[Ros93] Richard J. Rossi. Book review: [untitled]. The American Statistician, 47(1):76, February 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684790.

Rossi:1994:BRB

[Ros94] Richard J. Rossi. Book review: [untitled]. The American Statistician, 48(2):172, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684278.

Ross:1995:WGN

[Ros95a] N. Phillip Ross. What the government needs. The American Statistician, 49(1):7–9, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684801.

Rossman:1995:BRB

[Ros95b] Allan J. Rossman. Book review: [untitled]. The American Statistician, 49(2):237–238, May 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684649.

Ross:1996:BSR

[Ros96a] Sheldon M. Ross. Bayesians should not resample a prior sample to learn about the posterior. The American Statistician, 50(2):116, May 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684421.

Rossman:1996:BRB

[Ros96b] Allan J. Rossman. Book review: [untitled]. The American Statistician, 50(3):277–278, August 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684676.

Rotondo:1990:BHS

[Rot90] John A. Rotondo. A brief history of the subsection on statistics in marketing. The American Statistician, 44(2):82–83, May 1990.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684133.

Roberts:1992:BRB

[RR92] Rosemary A. Roberts and David L. Roberts. Book review: [untitled]. The American Statistician, 46(2):129–130, May 1992. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684182.

Roberts:1993:BRB

[RR93] Rosemary A. Roberts and David L. Roberts. Book review: [untitled]. The American Statistician, 47(1):40–41, February 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684782.

Rousseeuw:1999:BBB

[RRT99] Peter J. Rousseeuw, Ida Ruts, and John W. Tukey. The bagplot: a bivariate boxplot. The American Statistician, 53(4):382–387, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2686061.

Radke-Sharpe:1991:WCS

[RS91a] Norean Radke-Sharpe. Writing as a component of statistics education. The American Statistician, 45(4):292–293, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684457.

Rae:1991:ASR

[RS91b] Gordon Rae and John E. Spencer. Average Spearman’s rho, concordance, and the matching problem: Some relationships. The American Statistician, 45(2):161–162, May 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684395.

Raftery:1993:IAR

[RS93] Adrian E. Raftery and Tore Schweder. Inference about the ratio of two parameters, with application to whale censusing. The American Statistician, 47(4):259–264, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685283.

RS:1995:BRB

[RS95] RS. Book review: [untitled]. The American Statistician, 49(3):322, August 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684214.

Roberts:1999:APS

[RSW99] Rosemary Roberts, Richard Scheaffer, and Ann Watkins. Advanced placement statistics—past, present, and future. The American Statistician, 53(4):307–320, November 1999. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2686049.

Rumsey:1998:BRB

[Rum98] Deborah J. Rumsey. Book review: Business Statistics: Decision Making With Data. The American Statistician, 52(1):85–86, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685577.

Rumsey:1999:BRB

[Rum99] Deborah J. Rumsey. Book review: Exploring Statistics: a Modern Introduction to Data Analysis and Inference. The American Statistician, 53(3):291–292, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2686114.

Russell:1991:EVS

[Rus91] K. G. Russell. Estimating the value of e by simulation. The American Statistician, 45(1):66–68, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685243.

Russek:1998:BRA

[Rus98] Bernadette Russek. Book review: The Assessment Challenge in Statistics Education. The American Statistician, 52(3):281–282, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685944.

Rovine:1997:WLC

[RvE97] Michael J. Rovine and Alexander von Eye. A 14th way to look at a correlation coefficient: Correlation as the proportion of matches. The American Statistician, 51(1):42–46, February 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts/
rovine.html;   http://www.jstor.org/stable/2684692.

Stinnett:1990:WSS

[S^{+}90] Sandra Stinnett et al. Women in statistics: Sesquicentennial activities. The American Statistician, 44(2):74–80, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684131.

Sall:1990:LPG

[Sal90] John Sall. Leverage plots for general linear hypotheses. The American Statistician, 44(4):308–315, November 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684358.

Salzberg:1999:RSB

[Sal99] Alan J. Salzberg. Removable selection bias in quasi-experiments. The American Statistician, 53(2):103–107, May 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685726.

Samuels:1991:SRT

[Sam91] Myra L. Samuels. Statistical reversion toward the mean: More universal than regression toward the mean. The American Statistician, 45(4):344–346, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684474.

Samaniego:1992:EDG

[Sam92] Francisco J. Samaniego. Elementary derivations of geometric moments. The American Statistician, 46(2):108–109, May 1992. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684175.

Sapra:1991:CBL

[Sap91] Sunil K. Sapra. A connection between the logit model, normal discriminant analysis, and multivariate normal mixtures. The American Statistician, 45(4):265–268, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684450.

Saphire:1995:BRB

[Sap95] Diane G. Saphire. Book review: [untitled]. The American Statistician, 49(3):322, August 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684212.

Solomon:1994:EPS

[SATV94] Daniel L. Solomon, J. Tim Arnold, Bruce E. Trumbo, and Paul F. Velleman. Electronic publications in statistics — ready or not, here they come. The American Statistician, 48(3):191–196, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684715.

Saville:1990:MCP

[Sav90] D. J. Saville. Multiple comparison procedures: The practical solution. The American Statistician, 44(2):174–180, May 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684163.

Saw:1992:ASS

[Saw92] Sutaip L. C. Saw. ANOVA sums of squares as quadratic forms. The American Statistician, 46(4):288–290, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685315.

Sawyer:1998:MRE

[Saw98] Hon. Thomas C. Sawyer. Measuring race and ethnicity: Meeting public policy goals. The American Statistician, 52(1): 34–35, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.amstat.

org/publications/tas/abstracts_98/sawyer.html;   http:
//www.jstor.org/stable/2685563.

Saxena:1994:BRB

[Sax94] K. M. Lal Saxena. Book review: [untitled]. The American Statistician, 48(1):59, February 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685091.

Strand:1998:SPM

[SB98] Matthew Strand and Duane Boes. Statistical practice — modeling road racing times of competitive recreational runners using extreme value theory. The American Statistician, 52(3):205–210, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.amstat.

org/publications/tas/abstracts_98/strand.html;   http:
//www.jstor.org/stable/2685925.

Seely:1997:CSS

[SBL97] Justus F. Seely, David Birkes, and Youngjo Lee. Characterizing sums of squares by their distributions. The American Statistician, 51(1):55–58, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:/

/www.amstat.org/publications/tas/abstracts/seely.html;
 http://www.jstor.org/stable/2684696.

Stanek:1995:LE

[SBS^{+}95] Edward J. Stanek, III, John P. Buonaccorsi, Carl J. Schwarz, David M. Nickerson, Thomas N. Herzog, Thomas J. Boardman, Richard W. Katz, Roman Viveros, K. Balasubramanian, and N. Balakrishnan. Letters to the editor. The American Statistician, 49(3):323–326, August 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684215.

Samuel-Cahn:1994:CUE

[SC94] Ester Samuel-Cahn. Combining unbiased estimators. The American Statistician, 48(1):34–36, February 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685083.

Silver:1997:PSD

[SC97] Edward A. Silver and Daniel Costa. A property of symmetric distributions and a related order statistic result. The American Statistician, 51(1):32–33, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.amstat.org/publications/tas/abstracts/silver.html;
 http://www.jstor.org/stable/2684689.

Shmueli:1999:ADH

[SC99] Galit Shmueli and Ayala Cohen. Analysis and display of hierarchical life-time data. The American Statistician, 53(2):140–146, May 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685734.

Scheaffer:1990:EIC

[Sch90] Richard L. Scheaffer. Editor’s invited column: Toward a more quantitatively literate citizenry. The American Statistician, 44(1): 2–3, February 1990. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2684946.

Schechtman:1991:EAV

[Sch91a] Edna Schechtman. On estimating the asymptotic variance of a function of U statistics. The American Statistician, 45(2):103–106, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684368.

Schucany:1991:E

[Sch91b] William R. Schucany. Editorial. The American Statistician, 45 (4):263–264, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684449.

Schucany:1991:EIC

[Sch91c] William R. Schucany. Editor’s invited column. The American Statistician, 45(1):1, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685229.

Schucany:1992:ER

[Sch92] William R. Schucany. Editor’s report. The American Statistician, 46(4):245, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685305.

Schey:1993:RBM

[Sch93a] Harry M. Schey. The relationship between the magnitudes of SSR(x_{2} ) and SSR(x_{2}|x_{1} ): a geometric description. The American Statistician, 47(1):26–30, February 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684778.

Schucany:1993:ER

[Sch93b] William R. Schucany. Editor’s report. The American Statistician, 47(4):239–240, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685279.

Schumm:1993:BRB

[Sch93c] Walter R. Schumm. Book review: [untitled]. The American Statistician, 47(2):155–156, May 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685203.

Schwarz:1993:MMA

[Sch93d] Carl J. Schwarz. The mixed-model ANOVA: The truth, the computer packages, the books. Part I: Balanced data. The American Statistician, 47(1):48–59, February 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684786.

Schwenke:1993:BRBa

[Sch93e] James R. Schwenke. Book review: [untitled]. The American Statistician, 47(1):77, February 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684792.

Schwenke:1993:BRBb

[Sch93f] James R. Schwenke. Book review: [untitled]. The American Statistician, 47(2):156, May 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685204.

Schervish:1996:VWT

[Sch96a] Mark J. Schervish. P values: What they are and what they are not. The American Statistician, 50(3):203–206, August 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684655.

Scheuren:1996:BRB

[Sch96b] Fritz Scheuren. Book review: [untitled]. The American Statistician, 50(3):276, August 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684674.

Schwertman:1996:CBQ

[Sch96c] Neil C. Schwertman. A connection between quadratic-type confidence limits and fiducial limits. The American Statistician, 50 (3):242–243, August 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684663.

Scheaffer:1997:DBB

[Sch97] Richard L. Scheaffer. Discussion — Bayes for beginners? some reasons to hesitate. The American Statistician, 51(3):268–270, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684902.

Schou:1999:BRB

[Sch99] Sue B. Schou. Book review: Data, Statistics, and Decision Models with Excel. The American Statistician, 53(4):389, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2686063.

Sapra:1995:LE

[SCHI95] Sunil K. Sapra, Michael Cain, Werner E. Helm, and Peter B. Imrey. Letters to the editor. The American Statistician, 49(2):240–241, May 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684652.

Searle:1994:AVC

[Sea94] Shayle R. Searle. Analysis of variance computing package output for unbalanced data from fixed-effects models with nested factors. The American Statistician, 48(2):148–153, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684275.

Searle:1997:BRB

[Sea97] Shayle R. Searle. Built-in restrictions on best linear unbiased predictors (BLUP) of random effects in mixed models. The American Statistician, 51(1):19–21, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.amstat.org/publications/tas/abstracts/searle.html;
 http://www.jstor.org/stable/2684686.

Senn:1990:RNM

[Sen90] Stephen Senn. Regression: a new mode for an old meaning? The American Statistician, 44(2):181–183, May 1990. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684164.

Seymann:1991:LMD

[Sey91] Richard G. Seymann. [Let’s make a deal: The Player’s Dilemma]: Comment. The American Statistician, 45(4):287–288, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684454.

Sprott:1993:DBT

[SF93] D. A. Sprott and V. T. Farewell. The difference between two normal means. The American Statistician, 47(2):126–128, May 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685194.

Schott:1990:LE

[SFP^{+}90] Brian Schott, R. W. Farebrother, Victor R. Prybutok, Steve M. Bajgier, MaryAnne Atkinson, Grace Wahba, Noel Cressie, Richard Goldstein, Richard A. Lockhart, John J. Spinelli, Robert Kinnison, A. S. C. Ehrenberg, James D. Knoke, Roderick J. A. Little, W. E. Sharp, Haim Shore, Robert M. Norton, P. A. V. B. Swamy, Sastry G. Pantula, Marcia Gumpertz, Jay H. Beder, T. D. Stanley, Clint Cummins, and H. D. Vinod. Letters to the editor. The American Statistician, 44(3):254–264, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685353.

Sampson:1991:PTT

[SG91] Paul D. Sampson and Peter Guttorp. Power transformations and tests of environmental impact as interaction effects. The American Statistician, 45(2):83–89, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684365.

Smith:1992:BST

[SG92] A. F. M. Smith and A. E. Gelfand. Bayesian statistics without tears: A sampling–resampling perspective. The American Statistician, 46(2):84–88, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684170.

Spence:1993:RS

[SG93] Ian Spence and Robert F. Garrison. A remarkable scatterplot. The American Statistician, 47(1):12–19, February 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684775.

Shettle:1998:LMS

[SG98] Carolyn Shettle and Catherine Gaddy. The labor market for statisticians and other scientists. The American Statistician, 52(4):295–302, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/tas/abstracts_98/shettle.html;
 http://www.jstor.org/stable/2685430.

Schwertman:1990:SNP

[SGC90] Neil C. Schwertman, A. J. Gilks, and J. Cameron. A simple noncalculus proof that the median minimizes the sum of the absolute deviations. The American Statistician, 44(1):38–39, February 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684955.

Shuster:1992:LE

[SGN^{+}92] Jonathan J. Shuster, Sander Greenland, David P. Nichols, Ayala Cohen, and James R. Lackritz. Letters to the editor. The American Statistician, 46(2):163–165, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684190.

Schulman:1996:LE

[SGO^{+}96] Andrew Schulman, Stephen J. Gange, Cornel G. Ormsby, Roger G. Johnston, Clarence W. Lienhard, and Ian T. Jolliffe. Letters to the editor. The American Statistician, 50(3):280–282, August 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684679.

Shaffer:1991:GMT

[Sha91a] Juliet Popper Shaffer. The Gauss–Markov theorem and random regressors. The American Statistician, 45(4):269–273, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684451.

Shah:1991:RBC

[Sha91b] Arvind K. Shah. Relationship between the coefficients of determination of algebraically related models. The American Statistician, 45(4):300–301, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684459.

Sharma:1994:BRB

[Sha94] Subhash C. Sharma. Book review: [untitled]. The American Statistician, 48(2):172–173, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684280.

Sherr:1993:BRB

[She93] Lawrence A. Sherr. Book review: [untitled]. The American Statistician, 47(2):120–121, May 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685192.

Sherman:1997:CSM

[She97] Michael Sherman. Comparing the sample mean and the sample median: An exploration in the exponential power family. The American Statistician, 51(1):52–54, February 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts/
sherman.html;   http://www.jstor.org/stable/2684695.

Sherman:1998:BRB

[She98] Malcolm J. Sherman. Book review: A Probabilistic Analysis of the Sacco and Vanzetti Evidence. The American Statistician, 52(1):83, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685573.

Shoemaker:1995:TDD

[Sho95] Lewis H. Shoemaker. Tests for differences in dispersion based on quantiles. The American Statistician, 49(2):179–182, May 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684634.

Short:1996:BRBa

[Sho96a] Thomas H. Short. Book review: [untitled]. The American Statistician, 50(2):194, May 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684441.

Short:1996:BRBb

[Sho96b] Thomas H. Short. Book review: [untitled]. The American Statistician, 50(3):278, August 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684677.

Short:1997:DBB

[Sho97] Thomas H. Short. Discussion — Bayes for beginners? some reasons to hesitate. The American Statistician, 51(3):263–264, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684899.

Short:1998:BRE

[Sho98] Thomas H. Short. Book review: Elementary Bayesian Statistics. The American Statistician, 52(3):282–283, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685945.

Shuster:1991:SRC

[Shu91] Jonathan J. Shuster. The statistician in a reverse cocaine sting. The American Statistician, 45(2):123–124, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684373.

Searls:1990:NEV

[SI90] Donald T. Searls and Pichai Intarapanich. A note on an estimator for the variance that utilizes the kurtosis. The American Statistician, 44(4):295–296, November 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684353.

Silvapulle:1997:CEI

[Sil97] Mervyn J. Silvapulle. A curious example involving the likelihood ratio test against one-sided hypotheses. The American Statistician, 51(2):178–180, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/tas/abstracts/silvapulle.html;
 http://www.jstor.org/stable/2685415.

Simon:1994:WSP

[Sim94] Julian L. Simon. What some puzzling problems teach about the theory of simulation and the use of resampling. The American Statistician, 48(4):290–293, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684836.

Simon:1996:BRB

[Sim96] Gary Simon. Book review: [untitled]. The American Statistician, 50(3):277, August 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684675.

Simonoff:1998:LRC

[Sim98] Jeffrey S. Simonoff. Logistic regression, categorical predictors, and goodness-of-fit: It depends on who you ask. The American Statistician, 52(1):10–14, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.amstat.org/publications/tas/abstracts_98/simonoff.
html;   http://www.jstor.org/stable/2685558.

Shah:1993:TMU

[SK93] Arvind K. Shah and K. Krishnamoorthy. Testing means using hypothesis-dependent variance estimates. The American Statistician, 47(2):115–117, May 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685190.

Schwertman:1995:RGP

[SK95] Neil C. Schwertman and Terry L. Kiser. The role of geometry in pairwise and mutual independence. The American Statistician, 49(1):56–59, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684813.

Samuels:1992:SSR

[SL92] Myra L. Samuels and Tai-Fang C. Lu. Sample size requirements for the back-of-the-envelope binomial confidence interval. The American Statistician, 46(3):228–231, August 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685222.

Stroup:1991:NNA

[SM91] W. W. Stroup and D. K. Mulitze. Nearest neighbor adjusted best linear unbiased prediction. The American Statistician, 45 (3):194–200, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684289.

Schau:1997:UMT

[SM97] Candace Schau and Nancy Mattern. Use of map techniques in teaching applied statistics courses. The American Statistician, 51(2):171–175, May 1997. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:/

/www.amstat.org/publications/tas/abstracts/schau.html;
 http://www.jstor.org/stable/2685413.

Schwertman:1991:PMN

[SMH91] Neil C. Schwertman, Thomas A. McCready, and Lesley Howard. Probability models for the NCAA regional basketball tournaments. The American Statistician, 45(1):35–38, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685236.

Smith:1995:RFO

[Smi95] Peter J. Smith. A recursive formulation of the old problem of obtaining moments from cumulants and vice versa. The American Statistician, 49(2):217–218, May 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684642.

Smith:1997:BRB

[Smi97] Paul J. Smith. Book review: [untitled]. The American Statistician, 51(3):293–295, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684909.

Smith:1998:BRB

[Smi98] Marlene Smith. Book reviews: Basic Business Statistics: A Casebook and Business Analysis Using Regression: a Casebook. The American Statistician, 52(3):281, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:/

/www.amstat.org/publications/tas/ti-83.html;   http://
www.amstat.org/publications/tas/ti-83.pdf;   http://www.
jstor.org/stable/2685943.

Stein:1997:RSS

[SMP97] Philip G. Stein, James R. Matey, and Karen Pitts. A review of statistical software for the Apple Macintosh. The American Statistician, 51(1):67–82, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:/

/www.amstat.org/publications/tas/abstracts/stein.html;
 http://www.jstor.org/stable/2684699.

SNTEI:1991:GVP

[S.N91] S.N.Tibshirani Enterprises Inc. GAIM version 1.0: a package for generalized additive models. The American Statistician, 45(4): 339–340, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684468.

Snee:1990:STC

[Sne90] Ronald D. Snee. Statistical thinking and its contribution to total quality. The American Statistician, 44(2):116–121, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684144.

Snee:1993:WMS

[Sne93] Ronald D. Snee. What’s missing in statistical education? The American Statistician, 47(2):149–154, May 1993. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685201.

Snell:1995:JSE

[Sne95] J. Laurie Snell. The journal of statistics education information service and other Internet resources for statistics teachers. The American Statistician, 49(4):372–375, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684578.

Snell:1998:BRB

[Sne98] J. Laurie Snell. Book review: An Electronic Companion to Statistics (CD-ROM). The American Statistician, 52(1):88, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685581.

Shannon:1993:LE

[SNK^{+}93] Harry S. Shannon, Geoffrey R. Norman, Harry J. Khamis, Robert L. Andrews, Donald Watts, Katherine L. Monti, Gerard E. Dallal, Ingo Klein, David Branson, John W. Seaman, Niels Keiding, Gavin G. Gregory, Bernard Flury, A. Narayanan, Ravindra Khattree, Francisco J. Samaniego, Ralph W. Bailey, James M. Lucas, Nathan Mantel, Tosiya Sato, Hirofumi Takagi, Cyrus R. Mehta, Steve Walsh, Arnold J. Stromberg, Geoff McLachlan, Sunil K. Sapra, and R. W. Farebrother. Letters to the editor. The American Statistician, 47(1):79–88, February 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684794.

Samsa:1994:ISW

[SO94] Gregory Samsa and Eugene Z. Oddone. Integrating scientific writing into a statistics curriculum: a course in statistically based scientific writing. The American Statistician, 48(2):117–119, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684260.

Solomon:1992:SCS

[Sol92] Daniel L. Solomon. [Statistical consulting is scholarship]: Invited comment. The American Statistician, 46(4):299–300, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685318.

Solow:1993:ERI

[Sol93] Andrew R. Solow. Estimating record inclusion probability. The American Statistician, 47(3):206–208, August 1993. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684979.

Solow:1995:BRB

[Sol95] Andrew R. Solow. Book review: [untitled]. The American Statistician, 49(4):399, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684588.

Sonquist:1990:BRB

[Son90] John A. Sonquist. Book review: [untitled]. The American Statistician, 44(3):236–239, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685344.

Smith:1994:MSM

[SP94] Woollcott Smith and Milton Parnes. Mean streets: The median of a size-biased sample and the population mean. The American Statistician, 48(2):106–110, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684257.

Samaranayake:1998:BRB

[SP98] V. A. Samaranayake and Jagdish K. Patel. Book review: Modern Engineering Statistics and Elements of Engineering Probability and Statistics. The American Statistician, 52(2):199–201, May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685483.

Spirer:1990:BRB

[Spi90] Herbert F. Spirer. Book review: Statistical Software Teaching Package: Statistix 3.1. The American Statistician, 44(4):303–304, November 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684356.

Spurrier:1998:TCD

[Spu98] John D. Spurrier. Teacher’s corner — discussion — A onesemester, laboratory-based, quality-oriented statistics curriculum for engineering students. The American Statistician, 52(3):241, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685933.

Stromberg:1996:EIW

[SR96] Arnold J. Stromberg and Subathra Ramanathan. Easy implementation of writing in introductory statistics courses. The American Statistician, 50(2):159–163, May 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684429.

Sharpe:1997:RAS

[SR97] N. Radke Sharpe and R. A. Roberts. The relationship among sums of squares, correlation coefficients, and suppression. The American Statistician, 51(1):46–48, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.amstat.org/publications/tas/abstracts/sharpe.html;
 http://www.jstor.org/stable/2684693.

Sterling:1995:PDR

[SRW95] T. D. Sterling, W. L. Rosenbaum, and J. J. Weinkam. Publication decisions revisited: The effect of the outcome of statistical tests on the decision to publish and vice versa. The American Statistician, 49(1):108–112, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684823.

Schwarz:1991:IPS

[SS91] Carl J. Schwarz and M. Samanta. An inductive proof of the sampling distributions for the MLE’s of the parameters in an inverse Gaussian distribution. The American Statistician, 45(3):223–225, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684295.

Sellke:1997:CIU

[SS97] Thomas M. Sellke and Sarah H. Sellke. Chebyshev inequalities for unimodal distributions. The American Statistician, 51(1):34–40, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.amstat.org/publications/tas/abstracts/sellke.html;
 http://www.jstor.org/stable/2684690.

Silva:1998:NTG

[SS98] A. Pedro Duarte Silva and Antonie Stam. Nonparametric twogroup classification: Concepts and a SAS-based software package. The American Statistician, 52(2):185–197, May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685479.

Smith:1999:CNB

[SS99] Tyler Smith and Neil C. Schwertman. Can the NCAA Basketball Tournament seeding be used to predict margin of victory? The American Statistician, 53(2):94–98, May 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685724.

Sackrowitz:1999:VRV

[SSC99] Harold Sackrowitz and Ester Samuel-Cahn. P values as random variables—expected P values. The American Statistician, 53(4): 326–331, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2686051.

Schwertman:1996:MPM

[SSH96] Neil C. Schwertman, Kathryn L. Schenk, and Brett C. Holbrook. More probability models for the NCAA regional basketball tournaments. The American Statistician, 50(1):34–38, February 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685041.

StLaurent:1990:EMG

[St.90] Roy T. St. Laurent. The equivalence of the Milliken–Graybill procedure and the score test. The American Statistician, 44(1):36–37, February 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684954.

Spencer:1995:SSI

[ST95] William J. Spencer and Paul A. Tobias. Statistics in the semiconductor industry: A competitive necessity. The American Statistician, 49(3):245–249, August 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684194.

Stephenson:1990:SSR

[Ste90] W. Robert Stephenson. A study of student reaction to the use of Minitab in an introductory statistics course. The American Statistician, 44(3):231–235, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685343.

Stern:1991:PWF

[Ste91] Hal Stern. On the probability of winning a football game. The American Statistician, 45(3):179–183, August 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684286.

Stefanski:1992:MLR

[Ste92] Leonard A. Stefanski. Monotone likelihood ratio of a “Faulty-Inspection” distribution. The American Statistician, 46(2):110– 114, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684176.

Stefanski:1996:NAG

[Ste96] L. A. Stefanski. A note on the arithmetic-geometric-harmonic mean inequalities. The American Statistician, 50(3):246–247, August 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684665.

Stigler:1994:SVQ

[Sti94] Stephen M. Stigler. [Statistical values, quality, and certification]: Invited comment. The American Statistician, 48(2):71, May 1994.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684242.

Stine:1995:GIV

[Sti95] Robert A. Stine. Graphical interpretation of variance inflation factors. The American Statistician, 49(1):53–56, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684812.

Stigler:1999:FSS

[Sti99] Stephen M. Stigler. The foundations of statistics at Stanford. The American Statistician, 53(3):263–266, August 1999. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2686107.

Stone:1998:BPC

[Sto98] Richard A. Stone. The blind paper cutter: Teaching about variation, bias, stability, and process control. The American Statistician, 52(3):244–247, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.amstat.

org/publications/tas/abstracts_98/stone.html;   http:/
/www.jstor.org/stable/2685936.

Strayhorn:1990:EER

[Str90] Joseph M. Strayhorn. Estimating the errors remaining in a data set: Techniques for quality control. The American Statistician, 44 (1):14–18, February 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684949.

Strauss:1992:MFL

[Str92] David Strauss. The many faces of logistic regression. The American Statistician, 46(4):321–327, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685327.

Stukel:1990:GME

[Stu90] Therese A. Stukel. A general model for estimating ED_{100p} for binary response dose- response data. The American Statistician, 44 (1):19–22, February 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684950.

Sukhatme:1992:PTS

[Suk92] Shashikala Sukhatme. Powers of two-sample rank tests under the Lehmann alternatives. The American Statistician, 46(3):212–214, August 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685217.

Sullivan:1996:UDF

[Sul96] Jerry Sullivan. Understanding the degrees of freedom concept by computer experiments. The American Statistician, 50(3):234–237, August 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684661.

Singer:1990:ITA

[SW90] Judith D. Singer and John B. Willett. Improving the teaching of applied statistics: Putting the data back into data analysis. The American Statistician, 44(3):223–230, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685342.

Scott:1991:TR

[SW91] Alastair Scott and Chris Wild. Transformations and R^{2}. The American Statistician, 45(2):127–129, May 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684375.

Schilling:1994:SSP

[SW94] Mark F. Schilling and Ann E. Watkins. A suggestion for sunflower plots. The American Statistician, 48(4):303–305, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684839.

Tanis:1992:BRB

[Tan92] Elliot A. Tanis. Book review: [untitled]. The American Statistician, 46(4):304–306, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685322.

Tanis:1995:BRB

[Tan95] Elliot A. Tanis. Book review: [untitled]. The American Statistician, 49(2):219–220, May 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684643.

Tanis:1999:BRM

[Tan99] Elliot A. Tanis. Book review: Maple V Student Version: Release 5 . The American Statistician, 53(4):389–392, November 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2686064.

Therber:1992:LE

[TDF^{+}92] Andrew Therber, Kevin Delucchi, Martin Feuerman, David R. Fox, Kenneth Kaminsky, and Robert F. Ling. Letters to the editor. The American Statistician, 46(4):334–337, November 1992. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685329.

Therneau:1990:BRB

[The90] Terry M. Therneau. Book review: [untitled]. The American Statistician, 44(3):239–241, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685345.

Thompson:1995:BRB

[Tho95] Paul A. Thompson. Book review: Correspondence Analysis in Statistical Package Programs. The American Statistician, 49(3):310– 316, August 1995. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2684206.

Tibshirani:1997:WFM

[Tib97] Robert Tibshirani. Who is the fastest man in the world? The American Statistician, 51(2):106–111, May 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts/
tibshirani.html;   http://www.jstor.org/stable/2685397.

Tietjen:1994:RSC

[Tie94] Gary Tietjen. Recursive schemes for calculating cumulative binomial and Poisson probabilities. The American Statistician, 48(2): 136–137, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2684267.

Tietjen:1998:RSC

[Tie98] Gary Tietjen. Recursive schemes for calculating common cumulative distributions. The American Statistician, 52(3):263–265, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.amstat.org/publications/
tas/abstracts_98/tietjen.html;   http://www.jstor.org/
stable/2685940.

Tjur:1998:NRQ

[Tju98] Tue Tjur. Nonlinear regression, quasi likelihood, and overdispersion in generalized linear models. The American Statistician, 52 (3):222–227, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.amstat.org/

publications/tas/abstracts_98/tjur.html;   http://www.
jstor.org/stable/2685928.

Tingley:1993:NOC

[TL93] Maureen Tingley and Chunhui Li. A note on obtaining confidence intervals for discrete parameters. The American Statistician, 47(1):20–23, February 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684776.

Taylor:1995:CCB

[TM95] Douglas J. Taylor and Keith E. Muller. Computing confidence bounds for power and sample size of the general linear univariate model. The American Statistician, 49(1):43–47, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684810.

Tweedie:1999:CAR

[TM99] Richard Tweedie and Kerrie Mengersen. Calculating accuracy rates from multiple assessors with limited information. The American Statistician, 53(3):233–238, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2686102.

Thomas:1993:GIP

[TO93] Guy Thomas and John O’Quigley. A geometric interpretation of partial correlation using spherical triangles. The American Statistician, 47(1):30–32, February 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684779.

Tompkins:1993:BRB

[Tom93] Diane L. Tompkins. Book review: [untitled]. The American Statistician, 47(3):235, August 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684986.

Tucker:1995:F

[Tuc95] John R. Tucker. Foreword. The American Statistician, 49(1):1, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684798.

Turner:1997:SEI

[Tur97] Danny W. Turner. A simple example illustrating a wellknown property of the correlation coefficient. The American Statistician, 51(2):170, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.amstat.org/publications/tas/abstracts/turner.html;
 http://www.jstor.org/stable/2685412.

Tepping:1991:MHH

[TW91] Benjamin J. Tepping and Joseph Waksberg. Morris H. Hansen, 1910–1990. The American Statistician, 45(1):2–3, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685230.

Uyar:1990:RPS

[UE90] Bulent Uyar and Orhan Erdem. Regression procedures in SAS: Problems? The American Statistician, 44(4):296–301, November 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684354.

Unwin:1991:DFI

[Unw91] Antony Unwin. Diamond fast: Interactive graphics for time series. The American Statistician, 45(4):340, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684469.

Verrill:1990:WGP

[VAD90] Steve Verrill, Michael Axelrod, and Mark Durst. We’ve got the positive correlation BLUEs. The American Statistician, 44(2):171– 173, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684162.

Vangel:1996:CIN

[Van96a] Mark G. Vangel. Confidence intervals for a normal coefficient of variation. The American Statistician, 50(1):21–26, February 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685039.

Vanhonacker:1996:MAR

[Van96b] Wilfried R. Vanhonacker. Meta-analysis and response surface extrapolation: A least squares approach. The American Statistician, 50(4):294–299, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684923.

VanLeeuwen:1997:NCS

[Van97] Dawn M. VanLeeuwen. A note on the covariance structure in a linear model. The American Statistician, 51(2):140–144, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts/
vanleeuwen.html;   http://www.jstor.org/stable/2685406.

Vardeman:1992:WAO

[Var92] Stephen B. Vardeman. What about the other intervals? The American Statistician, 46(3):193–197, August 1992. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685212.

Vardeman:1998:TCD

[Var98] Stephen Vardeman. Teacher’s corner — discussion — A onesemester, laboratory-based, quality-oriented statistics curriculum for engineering students. The American Statistician, 52(3):242, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685934.

Viveros:1994:BNB

[VBB94] Roman Viveros, K. Balasubramanian, and N. Balakrishnan. Binomial and negative binomial analogues under correlated Bernoulli trials. The American Statistician, 48(3):243–247, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684728.

Verrill:1994:SVQ

[Ver94] Steve Verrill. [Statistical values, quality, and certification]: Invited comment. The American Statistician, 48(2):78–79, May 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684250.

Verrill:1999:WGC

[Ver99] Steve Verrill. When good confidence intervals go bad: Predictor sort experiments and ANOVA. The American Statistician, 53 (1):38–42, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685650.

Vidakovic:1994:ARL

[Vid94] Brani Vidakovic. All roads lead to Rome — even in the honeycomb world. The American Statistician, 48(3):234–236, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684723.

Visart:1991:IID

[Vis91] N. Visart. IDAMS: Internationally developed data analysis and management software. The American Statistician, 45(4):340, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684470.

Valz:1990:SDV

[VM90] Paul D. Valz and A. Ian McLeod. A simplified derivation of the variance of Kendall’s rank correlation coefficient. The American Statistician, 44(1):39–40, February 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684956.

vanBelle:1993:SSF

[vM93] Gerald van Belle and Donald C. Martin. Sample size as a function of coefficient of variation and ratio of means. The American Statistician, 47(3):165–167, August 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684968.

Velleman:1996:MTS

[VM96] Paul F. Velleman and David S. Moore. Multimedia for teaching statistics: Promises and pitfalls. The American Statistician, 50 (3):217–225, August 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684658.

Voss:1999:RMM

[Vos99] Daniel T. Voss. Resolving the mixed models controversy. The American Statistician, 53(4):352–356, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2686056.

vonEye:1996:SNL

[vS96] Alexander von Eye and Christiane Spiel. Standard and nonstandard log-linear symmetry models for measuring change in categorical variables. The American Statistician, 50(4):300–305, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684924.

vosSavant:1991:LE

[vSMC^{+}91] Marilyn vos Savant, John P. Morgan, Narasina R. Chaganty, Ram C. Dahiga, Michael J. Doviak, Nicholas R. Farnum, and Duncan K. H. Fong. Letters to the editor. The American Statistician, 45(4):347–348, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684475.

VanEck:1990:MP

[VV90] Neal A. Van Eck and Susan Van Eck. MicrOsiris for the PC. The American Statistician, 44(1):54, February 1990. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684961.

Velleman:1993:NOI

[VW93] Paul F. Velleman and Leland Wilkinson. Nominal, ordinal, interval, and ratio typologies are misleading. The American Statistician, 47(1):65–72, February 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684788.

Wainer:1996:DE

[Wai96] Howard Wainer. Depicting error. The American Statistician, 50(2):101–111, May 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684419.

Waller:1995:DCF

[Wal95a] Lance A. Waller. Does the characteristic function numerically distinguish distributions? The American Statistician, 49(2):150–152, May 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684629.

Walsh:1995:EPN

[Wal95b] Dennis P. Walsh. Equating Poisson and normal probability functions to derive Stirling’s formula. The American Statistician, 49 (3):270–271, August 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684198.

Wallman:1998:DRE

[Wal98] Katherine K. Wallman. Data on race and ethnicity: Revising the Federal standard. The American Statistician, 52(1):31–33, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.amstat.org/publications/
tas/abstracts_98/wallman.html;   http://www.jstor.org/
stable/2685562.

Wann:1991:RE

[Wan91] Marie D. Wann. A. Ross Eckler, 1901–1991. The American Statistician, 45(3):173, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684284.

Wand:1997:DBC

[Wan97] M. P. Wand. Data-based choice of histogram bin width. The American Statistician, 51(1):59–64, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.amstat.org/publications/tas/abstracts/wand.html;
http://www.jstor.org/stable/2684697.

Wand:1999:CDB

[Wan99] M. P. Wand. Correction: Data-based choice of histogram bin width. The American Statistician, 53(2):174, May 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685743.

Wardrop:1993:BRB

[War93] Daniel M. Wardrop. Book review: [untitled]. The American Statistician, 47(3):234–235, August 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684985.

Warde:1995:BRB

[War95a] William D. Warde. Book review: [untitled]. The American Statistician, 49(1):113–114, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684825.

Wardrop:1995:SPH

[War95b] Robert L. Wardrop. Simpson’s paradox and the hot hand in basketball. The American Statistician, 49(1):24–28, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684806.

Wardell:1997:SSI

[War97] Don G. Wardell. Small-sample interval estimation of Bernoulli and Poisson parameters. The American Statistician, 51(4): 321–325, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.amstat.org/

publications/tas/abstracts/wardell.html;   http://www.
jstor.org/stable/2685899.

Wasserstein:1993:BRB

[Was93] Ronald L. Wasserstein. Book review: [untitled]. The American Statistician, 47(2):155, May 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685202.

Watts:1991:WIS

[Wat91] Donald G. Watts. Why is introductory statistics difficult to learn? and what can we do to make it easier? The American Statistician, 45(4):290–291, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684456.

Watkins:1995:BRB

[Wat95] Ann E. Watkins. Book review: Review of the Proceedings of a Conference on Data Analysis in the Schools. The American Statistician, 49(1):62–63, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684815.

Waterman:1996:BRB

[Wat96] Richard Waterman. Book review: [untitled]. The American Statistician, 50(2):194–195, May 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684442.

Wasserstein:1991:BPL

[WB91] Ronald L. Wasserstein and John E. Boyer, Jr. Bounds on the power of linear rank tests for scale parameters. The American Statistician, 45(1):10–13, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685232.

Wolfe:1991:SCA

[WC91] Nicholas A. Wolfe and Alexander J. Chester. SIMCLUST: a cluster analysis program for ecological data. The American Statistician, 45(2):158, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684391.

Wiseman:1997:MLB

[WC97] Frederick Wiseman and Sangit Chatterjee. Major league baseball player salaries: Bringing realism into introductory statistics courses. The American Statistician, 51(4):350–352, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.amstat.org/publications/tas/
abstracts/wiseman.html;   http://www.jstor.org/stable/
2685904.

Weaver:1997:BRB

[Wea97] Sheila O’Leary Weaver. Book review: [untitled]. The American Statistician, 51(1):93, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684701.

Welsch:1992:SCS

[Wel92] Roy E. Welsch. [Statistical consulting is scholarship]: Invited comment. The American Statistician, 46(4):302–304, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685321.

Weldon:1998:BRB

[Wel98] Larry Weldon. Book review: Basic Statistics: Tales of Distributions. The American Statistician, 52(1):86–87, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685579.

Westfall:1990:GPB

[Wes90] Peter H. Westfall. Graphical presentation of a basketball game. The American Statistician, 44(4):305–307, November 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684357.

Wiens:1999:WLN

[Wie99] Brian L. Wiens. When log-normal and gamma models give different results: a case study. The American Statistician, 53(2):89–93, May 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685723.

Williams:1991:MCA

[Wil91a] Richard H. Williams. Multiple comparisons among correlation coefficients. The American Statistician, 45(4):341, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684471.

Wilson:1991:ETE

[Wil91b] Hoyt G. Wilson. Electronic tables: Exact probabilities and percentage points. The American Statistician, 45(4):341, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684472.

Wilson:1992:SCS

[Wil92] William J. Wilson. Statistical consulting is scholarship. The American Statistician, 46(4):295–298, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685317.

Wild:1994:EWV

[Wil94a] C. J. Wild. Embracing the “Wider view” of statistics. The American Statistician, 48(2):163–171, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684277.

Willemain:1994:BSR

[Wil94b] Thomas R. Willemain. Bootstrap on a shoestring: Resampling using spreadsheets. The American Statistician, 48(1):40–42, February 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685085.

Wilkinson:1999:DP

[Wil99] Leland Wilkinson. Dot plots. The American Statistician, 53 (3):276–281, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2686111.

Winner:1995:BRB

[Win95] Larry Winner. Book review: [untitled]. The American Statistician, 49(1):114, February 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684826.

Wiorkowski:1990:LAB

[Wio90] John J. Wiorkowski. A lightly annotated bibliography of the publications of the American Statistical Association. The American Statistician, 44(2):106–113, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684142.

Witmer:1993:BRB

[Wit93] Jeffrey A. Witmer. Book review: [untitled]. The American Statistician, 47(1):41–42, February 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684783.

Witmer:1994:BRB

[Wit94] Jeffrey A. Witmer. Book review: [untitled]. The American Statistician, 48(3):247–248, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684729.

Witmer:1997:BBS

[Wit97a] Jeffrey A. Witmer. [Bayes for beginners? some reasons to hesitate]: Discussion. The American Statistician, 51(3):262–263, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684898.

Witmer:1997:BRB

[Wit97b] Jeffrey A. Witmer. Book review: [untitled]. The American Statistician, 51(1):97–98, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684709.

Witmer:1997:DBB

[Wit97c] Jeffrey A. Witmer. Discussion — Bayes for beginners? some reasons to hesitate. The American Statistician, 51(3):262–??, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Williford:1995:MCT

[WKB^{+}95] William O. Williford, William F. Krol, Stephen F. Bingham, Joseph F. Collins, and David G. Weiss. The multicenter clinical trials coordination center statistician: “More Than a Consultant”. The American Statistician, 49(2):221–225, May 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684644.

Wilson:1999:USC

[WL99] P. David Wilson and Patricia Langenberg. Usual and shortest confidence intervals on odds ratios from logistic regression. The American Statistician, 53(4):332–335, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2686052.

Warner:1996:UNN

[WM96] Brad Warner and Manavendra Misra. Understanding neural networks as statistical tools. The American Statistician, 50(4):284– 293, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2684922.

Wonnacott:1998:BRU

[Won98] Thomas Wonnacott. Book review: The Ultimate Resource 2 . The American Statistician, 52(2):202, May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685485.

Woodall:1993:BRB

[Woo93] William H. Woodall. Book review: [untitled]. The American Statistician, 47(3):235–236, August 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684987.

West:1998:STW

[WOR98] R. Webster West, R. Todd Ogden, and A. J. Rossini. Statistical tools on the World Wide Web. The American Statistician, 52 (3):257–262, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.amstat.org/

publications/tas/west.html;   http://www.jstor.org/stable/
2685939.

Wright:1990:WZD

[Wri90] Tommy Wright. When zero defectives appear in a sample: Upper bounds on confidence coefficients of upper bounds. The American Statistician, 44(1):40–41, February 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684957.

Wright:1992:LIR

[Wri92] Tommy Wright. Lagrange’s identity reveals correlation coefficient and straight-line connection. The American Statistician, 46(2): 106–107, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2684174.

Wild:1993:CTP

[WS93] C. J. Wild and G. A. F. Seber. Comparing two proportions from the same survey. The American Statistician, 47(3):178–181, August 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684972. See

correction [WS94].

Wild:1994:CCT

[WS94] C. J. Wild and G. A. F. Seber. Correction: “Comparing two proportions from the same survey” [Amer. Statist. 47 (1993), no. 3, 178–181]. The American Statistician, 48(3):269, ???? 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [WS93].

Wang:1993:IDP

[WSS93] Y. H. Wang, Jordan Stoyanov, and Qi Man Shao. On independence and dependence properties of a set of random events. The American Statistician, 47(2):112–115, May 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685189.

Waller:1992:PPC

[WT92] Lance A. Waller and Bruce W. Turnbull. Probability plotting with censored data. The American Statistician, 46(1):5–12, February 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684401.

Wells:1990:LE

[WTA^{+}90] Martin T. Wells, Ram C. Tiwari, I. Arizono, H. Ohta, James W. Mergerson, Gabriella M. Belli, Christopher Cox, Murray A. Jorgensen, Jacques Benichou, Mitchell H. Gail, Warren F. Kuhfeld, Brian Dawkins, Walter B. Studdiford, Colin Goodall, W. D. Kaigh, Stephen W. Looney, Robert Kinnison, James A. Gibbons, Joel R. Levin, Ronald C. Serlin, K. J. A. Revfeim, Charles R. McConnell, Robert M. Norton, R. W. Farebrother, I. J. Good, Stanley Lebergott, and Vedula N. Murty. Letters to the editor. The American Statistician, 44(1):56–65, February 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684963.

Waller:1995:ODF

[WTH95] Lance A. Waller, Bruce W. Turnbull, and J. Michael Hardin. Obtaining distribution functions by numerical inversion of characteristic functions with applications. The American Statistician, 49(4):346–350, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684571.

Wuensch:1998:BRB

[Wue98] Karl L. Wuensch. Book review: Understanding Statistics: An Introduction for the Social Sciences. The American Statistician, 52 (1):87–88, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685580.

Westfall:1997:MTD

[WW97] Peter H. Westfall and Russell D. Wolfinger. Multiple tests with discrete distributions. The American Statistician, 51(1): 3–8, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.amstat.

org/publications/tas/abstracts/westfall.html;   http:/
/www.jstor.org/stable/2684683.

Yu:1994:CSS

[Yu94a] Philip L. H. Yu. Correction: a simple statistical project: Image reconstruction. The American Statistician, 48(4):354, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684857. See

[Yu94b].

Yu:1994:SSP

[Yu94b] Philip L. H. Yu. A simple statistical project: Image reconstruction. The American Statistician, 48(1):30–34, February 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685082. See correction [Yu94a].

Zelterman:1995:ETS

[ZCM95] Daniel Zelterman, Ivan Siu-Fung Chan, and Paul W. Mielke, Jr. Exact tests of significance in higher dimensional tables. The American Statistician, 49(4):357–361, November 1995. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684573.

Zehna:1991:PI

[Zeh91] Peter W. Zehna. On proving that X̄ and S^{2} are independent. The American Statistician, 45(2):121–122, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684372.

Ziegel:1990:ASA

[ZH90] Eric R. Ziegel and J. Stuart Hunter. American Statistical Association’s section on physical and engineering sciences: Past, present, and future. The American Statistician, 44(4):267–269, November 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684344.

Zheng:1994:CCR

[ZM94] Qi Zheng and James H. Matis. Correlation coefficient revisited. The American Statistician, 48(3):240–241, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684726.

Zhou:1999:CSP

[ZPH99] Xiao-Hua Zhou, Anthony J. Perkins, and Siu L. Hui. Comparisons of software packages for generalized linear multilevel models. The American Statistician, 53(3):282–290, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2686112.

Zurawski:1991:GRD

[Zur91] John Zurawski. GENSTAT 5: Recent developments. The American Statistician, 45(2):160, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684394.